### v3

Abstract.
We study three levels in a hierarchy of nondeterminism: A nondeterministic automaton ${\mathcal{A}}$ is determinizable by pruning (DBP) if we can obtain a deterministic automaton equivalent to ${\mathcal{A}}$ by removing some of its transitions. Then, ${\mathcal{A}}$ is history deterministic (HD) if its nondeterministic choices can be resolved in a way that only depends on the past. Finally, ${\mathcal{A}}$ is semantically deterministic (SD) if different nondeterministic choices in ${\mathcal{A}}$ lead to equivalent states. Some applications of automata in formal methods require deterministic automata, yet in fact can use automata with some level of nondeterminism. For example, DBP automata are useful in the analysis of online algorithms, and HD automata are useful in synthesis and control. For automata on finite words, the three levels in the hierarchy coincide. We study the hierarchy for Büchi, co-Büchi, and weak automata on infinite words. We show that the hierarchy is strict, study the expressive power of the different levels in it, as well as the complexity of deciding the membership of a language in a given level. Finally, we describe a probability-based analysis of the hierarchy, which relates the level of nondeterminism with the probability that a random run on a word in the language is accepting. We relate the latter to nondeterministic automata that can be used when reasoning about probabilistic systems.
Key words and phrases:
Automata, Nondeterminism
1. Introduction
Nondeterminism is a fundamental notion in theoretical computer science. It allows a computing machine to examine several possible actions simultaneously. For automata on finite words, nondeterminism does not increase the expressive power, yet it leads to an exponential succinctness [28].
[…]
For automata on infinite words, nondeterminism may increase the expressive power and also leads to an exponential succinctness. For example, NBWs are strictly more expressive than DBWs [22], whereas NCWs are as expressive as DCWs [24]. In some applications of the automata-theoretic approach, such as model checking, algorithms can be based on nondeterministic automata, whereas in other applications, such as synthesis and control, they cannot. There, the advantages of nondeterminism are lost, and algorithms involve a complicated determinization construction [29] or acrobatics for circumventing determinization [21]. Essentially, the inherent difficulty of using nondeterminism in synthesis and control lies in the fact that each guess of the nondeterministic automaton should accommodate all possible futures.
A study of nondeterministic automata that can resolve their nondeterministic choices in a way that only depends on the past started in [19], where the setting is modeled by means of tree automata for derived languages. It then continued by means of history deterministic (HD) automata [15]. [^1] The notion used in [15] is good for games (GFG) automata, as they address the difficulty of playing games on top of a nondeterministic automaton. As it turns out, the property of being good for games varies in different settings and HD is good for applications beyond games. Therefore, we use the term history determinism, introduced by Colcombet in the setting of quantitative automata with cost functions [12]. A nondeterministic automaton ${\mathcal{A}}$ over an alphabet $\Sigma$ is HD if there is a strategy $g$ that maps each finite word $u\in\Sigma^{*}$ to the transition to be taken after $u$ is read; and following $g$ results in accepting all the words in the language of ${\mathcal{A}}$. Note that a state $q$ of ${\mathcal{A}}$ may be reachable via different words, and $g$ may suggest different transitions from $q$ after different words are read. Still, $g$ depends only on the past, namely on the word read so far. Obviously, there exist HD automata: deterministic ones, or nondeterministic ones that are determinizable by pruning (DBP); that is, ones that embody an equivalent deterministic automaton. In fact, the HD automata constructed in [15] are DBP. [^2] As explained in [15], the fact that the HD automata constructed there are DBP does not contradict their usefulness in practice, as their transition relation is simpler than the one of the embodied deterministic automaton and it can be defined symbolically. Beyond the theoretical interest in DBP automata, they are used for modelling online algorithms: by relating the “unbounded look ahead” of optimal offline algorithms with nondeterminism, and relating the “no look ahead” of online algorithms with determinism, it is possible to reduce questions about the competitive ratio of online algorithms and the memory they require to questions about DBPness [2, 3].
For automata on finite words, HD-NFWs are always DBP [19, 25]. For automata on infinite words, HD-NBWs and HD-NCWs are as expressive as DBWs and DCWs, respectively [19, 27], but they need not be DBP [5]. Moreover, the best known determinization construction for HD-NBWs is quadratic, and determinization of HD-NCWs has a tight exponential blow-up [17]. Thus, HD automata on infinite words are (possibly even exponentially) more succinct than deterministic ones.
Further research studies characterization, typeness, complementation, and further constructions and decision procedures for HD automata [17, 7, 4], as well as an extension of the HD setting to pushdown $\omega$-automata [23] and to alternating automata [8, 6].
A nondeterministic automaton is semantically deterministic (SD, for short) if its nondeterministic choices lead to states with the same language. Thus, for every state $q$ of the automaton and letter $\sigma\in\Sigma$, all the $\sigma$-successors of $q$ have the same language. Beyond the fact that semantically determinism is a natural relaxation of determinism, and thus deserves consideration, SD automata naturally arise in the setting of HD automata. Indeed, if different $\sigma$-successors of a state in an HD automaton have different languages, then we can prune the transitions to states whose language is contained in the language of another $\sigma$-successor. Moreover, since containment for HD automata can be checked in polynomial time, such a pruning can be done in polynomial time [14, 17]. Accordingly (see more in Section 2.3), we can assume that all HD automata are SD [17].
Hence, we obtain the following hierarchy, from deterministic to nondeterministic automata, where each level is a special case of the levels to its right.
For automata on finite words, as NFWs are as expressive as DFWs, all levels of the hierarchy coincide in their expressive power. In fact, the three internal levels coincide already in the syntactic sense: every SD-NFW is DBP. Also, given an NFW, deciding whether it is SD, HD or DBP, can each be done in polynomial time [2].
For Büchi and co-Büchi automata, the picture is less clear, and is the subject of our research. Before we describe our results, let us mention that an orthogonal level of nondeterminism is that of unambiguous automata, namely automata that have a single accepting run on each word in their languages. An unambiguous NFW for a non-empty language is SD iff it is deterministic, and a DBP-NFW need not be unambiguous. It is known, however, that an HD unambiguous NCW, or NBW, is DBP [7].
We study the following aspects and questions about the hierarchy.
Strictness
For each nondeterminism level we study for which acceptance condition it is a strict super class of its preceding level. Recall that not all HD-NBWs and HD-NCWs are DBP [5], and examples for this include also SD automata. On the other hand, all HD-NWWs (in fact, all HD-NXWs whose language can be recognized by a DWW) are DBP [7]. We show that SD-NXWs need not be HD for all ${\rm X}\in\{{\rm B,C,W}\}$. Of special interest is our result on weak automata, whose properties typically agree with these of automata on finite words. Here, while all SD-NFWs are HD, this is not the case for SD-NWWs.
Expressive power
It is known that for all ${\rm X}\in\{{\rm B,C,W}\}$, HD-NXWs are as expressive as DXWs. We extend this result to semantic determinism and show that while SD-NXWs need not be HD, they are not more expressive, thus SD-NXWs are as expressive as DXWs. Since an SD-NXW need not be HD, this extends the known frontier of nondeterministic Büchi and weak automata that are not more expressive than their deterministic counterpart.
Deciding the nondeterminism level of an automaton
It is already known that deciding the HDness of a given NXW, for ${\rm X}\in\{{\rm B,C,W}\}$, can be done in polynomial time [2, 17, 4]. On the other hand, deciding whether a given NCW is DBP is NP-complete [16]. We complete the picture in three directions. First, we show that NP-completeness of deciding DBPness applies also to NBWs. Second, we show that in both cases, hardness applies even when the given automaton is HD. Thus, while it took the community some time to get convinced that not all HD automata are DBP, in fact it is NP-complete to decide whether a given HD-NBW or HD-NCW is DBP. Third, we study also the problem of deciding whether a given NXW is SD, and show that it is PSPACE-complete. Note that our results imply that the nondeterminism hierarchy is not monotone with respect to complexity: deciding DBPness, which is closest to determinism, is NP-complete, then HDness can be checked in polynomial time, and finally SDness is PSPACE-complete. Also, as PSPACE-hardness of checking SDness applies already to NWWs, we get another, even more surprising, difference between weak automata and automata on finite words. Indeed, for NFWs, all the three levels of nondeterminism coincide and SDness can be checked in polynomial time.
Approximated determinization by pruning
We say that a nondeterministic automaton ${\mathcal{A}}$ is almost-DBP if ${\mathcal{A}}$ embodies a deterministic automaton ${\mathcal{A}}^{\prime}$ such that the probability of a random word to be in $L({\mathcal{A}})\setminus L({\mathcal{A}}^{\prime})$ is 00. Clearly, if ${\mathcal{A}}$ is DBP, then it is almost-DBP. Recall that DBPness is associated with the ability of an online algorithm to perform as good as an offline one. A typical analysis of the performance of an on-line algorithm compares its performance with that of an off-line algorithm. The notion of almost-DBPness captures cases where the on-line algorithm performs, with probability $1$, as good as the offline algorithm. From the point of view of formal-language theory, almost-DBPness captures the ability to approximate from below the language of automata that need not be DBP. We study the almost-DBPness of HD and SD automata. Our results imply that the nondeterminism hierarchy “almost collapses”: We show that for Büchi (and hence also weak) automata, semantic determinism implies almost-DBPness, thus every SD-NBW is almost-DBP. Then, for co-Büchi automata, semantic determinism is not enough, and we need HDness. Thus, there is an SD-NCW that is not almost-DBP, yet all HD-NCWs are almost-DBP.
Reasoning about Markov decision processes
Another application in which nondeterminism is problematic is reasoning about probabilistic systems. Technically, such a reasoning involves the product of a Markov decision processes that models the system with an automaton for the desired behavior. When the automaton is nondeterministic, the product need not reflect the probability in which the system satisfies the behavior. A nondeterministic automaton is good-for-MDPs (GFM) if its product with Markov decision processes maintains the probability of acceptance. Thus, GFM automata can replace deterministic automata when reasoning about probabilistic systems [13, 30]. We study the relation between semantic determinism and GFMness. We show that all GFM automata are almost-DBP. On the other hand, semantic determinism implies GFMness only in Büchi automata (that is, all SD-NBWs are GFM), and a GFM automaton need not be SD, even for weak automata.
2. Preliminaries
2.1. Automata
For a finite nonempty alphabet $\Sigma$, an infinite word $w=\sigma_{1}\cdot\sigma_{2}\cdots\in\Sigma^{\omega}$ is an infinite sequence of letters from $\Sigma$. A language $L\subseteq\Sigma^{\omega}$ is a set of infinite words. For $i,j\geq 0$, we use $w[1,i]$ to denote the (possibly empty) prefix $\sigma_{1}\cdot\sigma_{2}\cdots\sigma_{i}$ of $w$, use $w[i+1,j]$ to denote the (possibly empty) infix $\sigma_{i+1}\cdot\sigma_{i+2}\cdots\sigma_{j}$ of $w$, and use $w[i+1,\infty]$ to denote its suffix $\sigma_{i+1}\cdot\sigma_{i+2}\cdots$. We sometimes refer also to languages of finite words, namely subsets of $\Sigma^{*}$. We denote the empty word by $\epsilon$.
A nondeterministic automaton over infinite words is ${\mathcal{A}}=\langle\Sigma,Q,q_{0},\delta,\alpha\rangle$, where $\Sigma$ is an alphabet, $Q$ is a finite set of states, $q_{0}\in Q$ is an initial state, $\delta:Q\times\Sigma\to 2^{Q}\setminus\emptyset$ is a transition function, and $\alpha$ is an acceptance condition, to be defined below. For states $q$ and $s$ and a letter $\sigma\in\Sigma$, we say that $s$ is a $\sigma$-successor of $q$ if $s\in\delta(q,\sigma)$.
Note that ${\mathcal{A}}$ is total, in the sense that it has at least one successor for each state and letter. If $|\delta(q,\sigma)|=1$ for every state $q\in Q$ and letter $\sigma\in\Sigma$, then ${\mathcal{A}}$ is deterministic.
A run of ${\mathcal{A}}$ on $w=\sigma_{1}\cdot\sigma_{2}\cdots\in\Sigma^{\omega}$ is an infinite sequence of states $r=r_{0},r_{1},r_{2},\ldots\in Q^{\omega}$, such that $r_{0}=q_{0}$, and for all $i\geq 0$, we have that $r_{i+1}\in\delta(r_{i},\sigma_{i+1})$. We extend $\delta$ to sets of states and finite words in the expected way. Thus, $\delta(S,u)$ is the set of states that ${\mathcal{A}}$ may reach when it reads the word $u\in\Sigma^{*}$ from some state in $S\in 2^{Q}$. Formally, $\delta:2^{Q}\times\Sigma^{*}\to 2^{Q}$ is such that for every $S\in 2^{Q}$, finite word $u\in\Sigma^{*}$, and letter $\sigma\in\Sigma$, we have that $\delta(S,\epsilon)=S$, $\delta(S,\sigma)=\bigcup_{s\in S}\delta(s,\sigma)$, and $\delta(S,u\cdot\sigma)=\delta(\delta(S,u),\sigma)$. The transition function $\delta$ induces a transition relation $\Delta\subseteq Q\times\Sigma\times Q$, where for every two states $q,s\in Q$ and letter $\sigma\in\Sigma$, we have that $\langle q,\sigma,s\rangle\in\Delta$ iff $s\in\delta(q,\sigma)$. For a state $q\in Q$ of ${\mathcal{A}}$, we define ${\mathcal{A}}^{q}$ to be the automaton obtained from ${\mathcal{A}}$ by setting the initial state to be $q$. Thus, ${\mathcal{A}}^{q}=\langle\Sigma,Q,q,\delta,\alpha\rangle$. Two states $q,s\in Q$ are equivalent, denoted $q\sim_{{\mathcal{A}}}s$, if $L({\mathcal{A}}^{q})=L({\mathcal{A}}^{s})$.
[…]
Consider a directed graph $G=\langle V,E\rangle$. A strongly connected set in $G$ (SCS, for short) is a set $C\subseteq V$ such that for every two vertices $v,v^{\prime}\in C$, there is a path from $v$ to $v^{\prime}$. A SCS is maximal if it is maximal w.r.t containment, that is, for every non-empty set $C^{\prime}\subseteq V\setminus C$, it holds that $C\cup C^{\prime}$ is not a SCS. The maximal strongly connected sets are also termed strongly connected components (SCCs, for short). The SCC graph of $G$ is the graph defined over the SCCs of $G$, where there is an edge from an SCC $C$ to another SCC $C^{\prime}$ iff there are two vertices $v\in C$ and $v^{\prime}\in C^{\prime}$ with $\langle v,v^{\prime}\rangle\in E$. A SCC is ergodic iff it has no outgoing edges in the SCC graph.
An automaton ${\mathcal{A}}=\langle\Sigma,Q,q_{0},\delta,\alpha\rangle$ induces a directed graph $G_{{\mathcal{A}}}=\langle Q,E\rangle$, where $\langle q,q^{\prime}\rangle\in E$ iff there is a letter $\sigma\in\Sigma$ such that $\langle q,\sigma,q^{\prime}\rangle\in\Delta$. The SCSs and SCCs of ${\mathcal{A}}$ are those of $G_{{\mathcal{A}}}$.
The $\alpha$-free SCCs of ${\mathcal{A}}$ are the SCCs of ${\mathcal{A}}$ that do not contain states from $\alpha$.
A Büchi automaton ${\mathcal{A}}$ is weak [26] if for each SCC $C$ in $G_{\mathcal{A}}$, either $C\subseteq\alpha$ (in which case we say that $C$ is an accepting SCC) or $C\cap\alpha=\emptyset$ (in which case we say that $C$ is a rejecting SCC). Note that a weak automaton can be viewed as both a Büchi and a co-Büchi automaton, as a run of ${\mathcal{A}}$ visits $\alpha$ infinitely often, iff it gets trapped in an accepting SCC, iff it visits states in $Q\setminus\alpha$ only finitely often.
[…]
2.2. Probability
Consider the probability space $(\Sigma^{\omega},\mathbb{P})$ where each word $w=\sigma_{1}\cdot\sigma_{2}\cdot\sigma_{3}\cdots\in\Sigma^{\omega}$ is drawn by taking the $\sigma_{i}$’s to be independent and identically distributed $\mathrm{Unif}(\Sigma)$. Thus. for all positions $i\geq 1$ and letters $\sigma\in\Sigma$, the probability that $\sigma_{i}$ is $\sigma$ is $\frac{1}{|\Sigma|}$. Let ${\mathcal{A}}=\langle\Sigma,Q,q_{0},\delta,\alpha\rangle$ be a determnistic automaton, and let $G_{\mathcal{A}}=\langle Q,E\rangle$ be its induced directed graph. A random walk on ${\mathcal{A}}$, is a random walk on $G_{\mathcal{A}}$ with the probability matrix $P(q,p)=\frac{|\{\sigma\in\Sigma:\langle q,\sigma,p\rangle\in\Delta\}|}{|\Sigma|}$. It is not hard to see that $\mathbb{P}(L({\mathcal{A}}))$ is precisely the probability that a random walk on ${\mathcal{A}}$ is an accepting run. Note that with probability $1$, a random walk on ${\mathcal{A}}$ reaches an ergodic SCC $C\subseteq Q$, where it visits all states infinitely often. It follows that $\mathbb{P}(L({\mathcal{A}}))$ equals the probability that a random walk on ${\mathcal{A}}$ reaches an ergodic accepting SCC.
2.3. Automata with Some Nondeterminism
Consider a nondeterministic automaton ${\mathcal{A}}=\langle\Sigma,Q,q_{0},\delta,\alpha\rangle$. The automaton ${\mathcal{A}}$ is determinizable by prunning (DBP) if we can remove some of the transitions of ${\mathcal{A}}$ and get a deterministic automaton ${\mathcal{A}}^{\prime}$ that recognizes $L({\mathcal{A}})$. We then say that ${\mathcal{A}}^{\prime}$ is a deterministic pruning of ${\mathcal{A}}$.
The automaton ${\mathcal{A}}$ is history deterministic (HD, for short) if its nondeterminism can be resolved based on the past, thus on the prefix of the input word read so far. Formally, ${\mathcal{A}}$ is HD if there exists a strategy $f:\Sigma^{*}\to Q$ such that the following hold:
  1. (1)
    The strategy $f$ is consistent with the transition function. That is, $f(\epsilon)=q_{0}$, and for every finite word $u\in\Sigma^{*}$ and letter $\sigma\in\Sigma$, we have that $\langle f(u),\sigma,f(u\cdot\sigma)\rangle\in\Delta$.
  2. (2)
    Following $f$ causes ${\mathcal{A}}$ to accept all the words in its language. That is, for every infinite word $w=\sigma_{1}\cdot\sigma_{2}\cdots\in\Sigma^{\omega}$, if $w\in L({\mathcal{A}})$, then the run $f(w[1,0]),f(w[1,1]),f(w[1,2]),\ldots$, which we denote by $f(w)$, is an accepting run of ${\mathcal{A}}$ on $w$.
We say that the strategy $f$ witnesses ${\mathcal{A}}$’s HDness. Note that a strategy $f$ need not use all the states or transitions of ${\mathcal{A}}$. In particular, a state $q$ of ${\mathcal{A}}$ may be such that ${\mathcal{A}}^{q}$ is not HD. We say that a state $q$ of ${\mathcal{A}}$ is HD if ${\mathcal{A}}^{q}$ is HD.
The automaton ${\mathcal{A}}$ is semantically deterministic (SD, for short) if different nondeterministic choices in ${\mathcal{A}}$ lead to equivalent states. Thus, for every state $q\in Q$ and letter $\sigma\in\Sigma$, all the $\sigma$-successors of $q$ are equivalent: for every two states $s,s^{\prime}\in\delta(q,\sigma)$, we have that $s\sim_{{\mathcal{A}}}s^{\prime}$.
Note that every deterministic automaton ${\mathcal{A}}$ is DBP (with ${\mathcal{A}}$ being the deterministic pruning of itself). Also, every nondeterministic DBP automaton is HD. Indeed, the deterministic pruning of ${\mathcal{A}}$ induces a witness strategy. Finally, if a state $q$ of an HD automaton has a $\sigma$-successor $s$ that does not accept all the words $w$ such that $\sigma\cdot w$ is in $L({\mathcal{A}}^{q})$, then the $\sigma$-transition from $q$ to $s$ can be removed. Moreover, by  [14, 17, 4], the detection of such transitions can be done in polynomial time. Hence, we can assume that every HD automaton, and hence also every DBP automaton, is SD. [^3] Using the terminology of [17], we can assume that the residual languages of all the $\sigma$-successors of each state in an HD automaton are equivalent.
For ${\rm X}\in\{{\rm B,C,W}\}$, we use SD-NXW, HD-NXW, and DBP-NXW to refer to the different classes of NXWs. For example, SD-NBWs are nondeterministic Büchi automata that are semantically-deterministic.
3. The Syntactic and Semantic Hierarchies
In this section we study syntactic and semantic hierarchies induced by the different levels of nondeterminism. We start with the syntactic hierarchy. For two classes ${\mathcal{C}}_{1}$ and ${\mathcal{C}}_{2}$ of automata, we use ${\mathcal{C}}_{1}\preceq{\mathcal{C}}_{2}$ to indicate that every automaton in ${\mathcal{C}}_{1}$ is also in ${\mathcal{C}}_{2}$. Accordingly, ${\mathcal{C}}_{1}\prec{\mathcal{C}}_{2}$ if ${\mathcal{C}}_{1}\preceq{\mathcal{C}}_{2}$ yet there are automata in ${\mathcal{C}}_{2}$ that are not in ${\mathcal{C}}_{1}$. For example, clearly, DFW $\prec$ NFW. We first show that the nondeterminism hierarchy is strict, except for all HD-NWWs being DBP. The latter is not surprising, as all HD-NFWs are DBP, and weak automata share many properties with automata on finite words. On the other hand, unlike the case of finite words, we show that not all SD-NWWs are HD. In fact the result holds already for NWWs that accept co-safety languages, namely all whose states except for an accepting sink are rejecting.
Theorem 1.
[Syntactic Hierarchy] For ${\rm X}\in\{{\rm B,C,W}\}$, we have that DXW $\prec$ DBP-NXW $\preceq$ HD-NXW $\prec$ SD-NXW $\prec$ NXW. For ${\rm X}\in\{{\rm B,C}\}$, the second inequality is strict.
Proof 3.1.
By definition, each class is a special case of the one to its right. We prove strictness. It is easy to see that the first and last strict inequalities hold. For the first strict inequality, for all ${\rm X}\in\{{\rm B,C,W}\}$, consider a nonempty DXW ${\mathcal{A}}$, and obtain an NXW ${\mathcal{B}}$ from ${\mathcal{A}}$ by duplicating some state and the transitions to it. Then, ${\mathcal{B}}$ is a DBP-NXW that is not a DXW. For the last inequality, consider an SD-NXW ${\mathcal{A}}$, and let $\sigma$ be a letter such that ${\mathcal{A}}$ accepts some word that starts with $\sigma$. Consider the NXW ${\mathcal{C}}$ obtained from ${\mathcal{A}}$ by adding a $\sigma$-transition from ${\mathcal{A}}$’s initial state to a new rejecting sink. Then, as at least one $\sigma$-successor of the initial state of ${\mathcal{A}}$ is not empty, we have that ${\mathcal{C}}$ is an NXW that is not an SD-NXW. The relation between DBPness and HDness has already been studied. It is shown in [5] that HD-NXW need not be DBP for ${\rm X}\in\{{\rm B,C}\}$, and shown in [7] that HD-NWW are DBP.
It is left to relate HDness and SDness. Consider the NWW ${\mathcal{W}}$ in Figure 1. It is not hard to check that ${\mathcal{W}}$ is weak, and note that it is SD, as all its states recognize the language $\{a,b\}^{\omega}$. Yet $\mathcal{W}$ is not HD, as every strategy has a word with which it does not reach $q_{acc}$ – a word that forces each visit in $q_{a}$ and $q_{b}$ to be followed by a visit in $q_{0}$.
Figure 1. An SD-NWW that is not HD.
Hence, HD-NWW $\prec$ SD-NWW. As weak automata are a special case of Büchi and co-Büchi, strictness for them follows.
We continue to the semantic hierarchy, where we study the expressive power of the different classes. Now, for two classes ${\mathcal{C}}_{1}$ and ${\mathcal{C}}_{2}$ of automata, we say that ${\mathcal{C}}_{1}$ is less expressive than ${\mathcal{C}}_{2}$, denoted ${\mathcal{C}}_{1}\leq{\mathcal{C}}_{2}$, if every automaton in ${\mathcal{C}}_{1}$ has an equivalent automaton in ${\mathcal{C}}_{2}$. Accordingly, ${\mathcal{C}}_{1}={\mathcal{C}}_{2}$ if ${\mathcal{C}}_{1}\leq{\mathcal{C}}_{2}$ and ${\mathcal{C}}_{2}\leq{\mathcal{C}}_{1}$, and ${\mathcal{C}}_{1}<{\mathcal{C}}_{2}$ if ${\mathcal{C}}_{1}\leq{\mathcal{C}}_{2}$ yet ${\mathcal{C}}_{2}\neq{\mathcal{C}}_{1}$. Since NCW=DCW, we expect the hierarchy to be strict only in the cases of Büchi and weak automata. As we now show, however, semantically deterministic automata are not more expressive than deterministic ones also in the case of Büchi and weak automata.
Theorem 2.
[Semantic Hierarchy] For $X\in\{B,W\}$, we have that DXW $=$ DBP-NXW $=$ HD-NXW $=$ SD-NXW $<$ NXW.
Proof 3.2.
In [19, 17], the authors suggest variants of the subset construction that determinize HD-NBWs. As we argue below, the construction in [17] is correct also when applied to SD-NBWs. Moreover, it preserves weakness. Thus, DBW$=$SD-NBW and DWW$=$SD-NWW. Also, the last inequality follows from the fact DBW$<$NBW and DWW$<$NWW [22]. Given an NBW ${\mathcal{A}}=\langle\Sigma,Q,q_{0},\delta,\alpha\rangle$, the DBW generated in [17] [^4] The construction in [17] assumes automata with transition-based acceptance, and (regardless of this) is slightly different: when $\alpha$ is visited, ${\mathcal{A}}^{\prime}$ continues with a single state from the set of successors. However, the key point remains the same: ${\mathcal{A}}$ being SD enables ${\mathcal{A}}^{\prime}$ to maintain only subsets of states, rather than Safra trees, which makes determinization much easier. is ${\mathcal{A}}^{\prime}=\langle\Sigma,Q^{\prime},q^{\prime}_{0},\delta^{\prime},\alpha^{\prime}\rangle$, where $Q^{\prime}=2^{Q}$, $q^{\prime}_{0}=\{q_{0}\}$, $\alpha^{\prime}=\{S\in 2^{Q}:S\subseteq\alpha\}$, and the transition function $\delta^{\prime}$ is defined for every subset $S\in 2^{Q}$ and letter $\sigma\in\Sigma$ as follows. If $\delta(S,\sigma)\cap\alpha=\emptyset$, then $\delta^{\prime}(S,\sigma)=\delta(S,\sigma)$. Otherwise, namely if $\delta(S,\sigma)\cap\alpha\neq\emptyset$, then $\delta^{\prime}(S,\sigma)=\delta(S,\sigma)\cap\alpha$.
Thus, we proceed as the standard subset construction, except that whenever a constructed set contains a state in $\alpha$, we leave in the set only states in $\alpha$. Accordingly, every reachable state $S\in Q^{\prime}$ contains only $\alpha$-states of ${\mathcal{A}}$ or only $\bar{\alpha}$-states of ${\mathcal{A}}$. Note that as ${\mathcal{A}}$ is SD, then for every two states $q,q^{\prime}\in Q$, letter $\sigma\in\Sigma$, and transitions $\langle q,\sigma,s\rangle,\langle q^{\prime},\sigma,s^{\prime}\rangle\in\Delta$, if $q\sim_{{\mathcal{A}}}q^{\prime}$, then $s\sim_{{\mathcal{A}}}s^{\prime}$. Consequently, every reachable state $S$ of ${\mathcal{A}}^{\prime}$ consists of $\sim_{\mathcal{A}}$-equivalent states. As we formally prove in Appendix A, these properties guarantee that indeed $L({\mathcal{A}}^{\prime})=L({\mathcal{A}})$ and that weakness of ${\mathcal{A}}$ is maintained in ${\mathcal{A}}^{\prime}$.
4. Deciding the Nondeterminism Level of an Automaton
In this section we study the complexity of the problem of deciding the nondeterminism level of a given automaton. Note that we refer here to the syntactic class (e.g., deciding whether a given NBW is HD) and not to the semantic one (e.g., deciding whether a given NBW has an equivalent HD-NBW). Indeed, by Theorem 2, the latter boils down to deciding whether the language of a given NXW, for ${\rm X}\in\{{\rm B,C,W}\}$, can be recognized by a DXW, which is well known: the answer is always “yes” for an NCW, and the problem is PSPACE-complete for NBWs and NWWs [20]. [^5] The proof in [20] is for NBWs, yet the arguments there apply also for weak automata.
Our results are summarized in Table 1. The entries there describe the complexity of deciding whether a given NXW belongs to a certain nondeterministic level (for example, the upper left cell describes the complexity of deciding whether an NBW is DBP). In fact, the results described are valid already when the given NXW is known to be one level above the questioned one with only two exceptions – deciding the DBPness of an HD-NCW, whose NP-hardness is proved in Th. 8, and deciding whether a given NWW is DBP, which is PTIME in general, and is $O(1)$ when the given NWW is HD, in which case the answer is always “yes”.
DBP HD SD
NBW NP-complete PTIME PSPACE-complete
4 [4] 3
NCW NP-complete PTIME PSPACE-complete
[16] [17] 3
NWW PTIME PTIME PSPACE-complete
[17, 4] [17, 4] 3
Table 1. The complexity of deciding the nondeterminizm level of an NXW, for ${\rm X}\in\{{\rm B,C,W}\}$. The results are valid also in the case the given NXW is one level to the right in the nondeterminism hierarchy, with the exception of deciding the DBPness of an HD-NCW, whose NP-hardness is proved in Th. 8, and deciding the DBPness of an HD-NWW, where the answer is always positive [7].
We start with the complexity of deciding semantic determinism.
Theorem 3.
Deciding whether an NXW is semantically deterministic is PSPACE-complete, for ${\rm X}\in\{{\rm B,C,W}\}$.
Proof 4.1.
Membership in PSPACE is easy, as we check SDness by polynomially many checks of language equivalence. Formally, given an NXW ${\mathcal{A}}=\langle\Sigma,Q,q_{0},\delta,\alpha\rangle$, a PSPACE algorithm goes over all states $q\in Q$, letters $\sigma$, and $\sigma$-successors $s$ and $s^{\prime}$ of $q$, and checks that $s\sim_{\mathcal{A}}s^{\prime}$. Since language equivalence can be checked in PSPACE [32] and there are polynomially many checks to perform, we are done.
Proving PSPACE-hardness, we do a reduction from polynomial-space Turing machines. Given a Turing machine $T$ with space complexity $s:\mathbb{N}\rightarrow\mathbb{N}$, we construct in time polynomial in $|T|$ and $s(0)$, an NWW ${\mathcal{A}}$ of size linear in $T$ and $s(0)$, such that ${\mathcal{A}}$ is SD iff $T$ accepts the empty tape [^6] This is sufficient, as one can define a generic reduction from every language $L$ in PSPACE as follows. Let $T_{L}$ be a Turing machine that decides $L$ in polynomial space $f(n)$. On input $w$ for the reduction, the reduction considers the machine $T_{w}$ that on every input, first erases the tape, writes $w$ on its tape, and then runs as $T_{L}$ on $w$. Then, the reduction outputs an automaton ${\mathcal{A}}$, such that $T_{w}$ accepts the empty tape iff ${\mathcal{A}}$ is SD. Note that the space complexity of $T_{w}$ is $s(n)=\max(n,f(|w|))$, and that $w$ is in $L$ iff $T_{w}$ accepts the empty tape. Since ${\mathcal{A}}$ is constructed in time polynomial in $s(0)=f(|w|)$ and $|T_{w}|=\mathrm{poly}(|w|)$, it follows that the reduction is polynomial in $|w|$.. Clearly, this implies a lower bound also for NBWs and NCWs. Let $n_{0}=s(0)$. Thus, each configuration in the computation of $T$ on the empty tape uses at most $n_{0}$ cells.
We assume that $T$ halts from all configurations (that is, not just from these reachable from an initial configuration of $T$); indeed, by adding to $T$ a step-counter that uses polynomial-space, one can transform a polynomial-space Turing machine that need not halt from all configurations to one that does halt.
We also assume, without loss of generality, that once $T$ reaches a final (accepting or rejecting) state, it erases the tape, moves with its reading head to the leftmost cell, and moves to the initial state. Thus, all computations of $T$ are infinite and after visiting a final configuration for the first time, they consists of repeating the same finite computation on the empty tape that uses $n_{0}$ tape cells.
We define ${\mathcal{A}}$ so that it accepts a word $w$ iff
  • (C1) $w$ is not a suffix of an encoding of a legal computation of $T$ that uses at most $n_{0}$ cells, or
  • (C2) $w$ includes an encoding of an accepting configuration of $T$ that uses at most $n_{0}$ cells.
It is not hard to see that if $T$ accepts the empty tape, then ${\mathcal{A}}$ is universal (that is, accepts all words). Indeed, each word $w$ is either not a suffix of an encoding of a legal computation of $T$ that uses at most $n_{0}$ cells, in which case $w$ is accepted thanks to C1, or
$w$ is such a suffix, in which case, since we assume that $T$ halts from any configuration, the encoded computation eventually reaches a final configuration and, by the definition of $T$, continues by an encoding of the computation of $T$ on the empty tape. Thus, $w$ eventually includes the encoding of the computation of $T$ on the empty tape. In particular, since $T$ accepts the empty tape, $w$ includes an encoding of an accepting configuration that uses at most $n_{0}$ tape cells, and so $w$ is accepted thanks to C2.
Also, if $T$ rejects the empty tape, then ${\mathcal{A}}$ rejects the word $w_{\varepsilon}$ that encodes the computation of $T$ on the empty tape. Indeed, C1 is not satisfied, and since by definition of $T$, the infinite computation of $T$ on the empty tape is just the same finite rejecting computation of $T$ on the empty tape repeated, then $w_{\varepsilon}$ clearly does not include an encoding of an accepting configuration, and so C2 is not satisfied too.
In order to define ${\mathcal{A}}$ so that it is SD iff $T$ accepts the empty tape, we define all its states to be universal iff $T$ accepts the empty tape. Intuitively, we do it by letting ${\mathcal{A}}$ guess and check the existence of an infix that witnesses satisfaction of C1 or C2, and also let it, at each point of its operation, go back to the initial state, where it can guess again. Recall that ${\mathcal{A}}$ is universal when $T$ accepts the empty tape. That is, for every infinite word $w$, all the suffixes of $w$ satisfy C1 or C2. Thus, ${\mathcal{A}}$ making a bad guess on $w$ does not prevent it from later branching into an accepting run.
We now describe the operation of ${\mathcal{A}}$ in more detail (see Figure 2).
Figure 2. The structure of the NWW constructed in 3.
In its initial state, ${\mathcal{A}}$ guesses which of C1 and C2 is satisfied. In case ${\mathcal{A}}$ guesses that C1 is satisfied, it guesses the place in which $w$ includes a violation of the encoding. As we detail in Appendix B, this amounts to guessing whether the violation is in the encoding of a single configuration encoded in $w$, or whether the violation is of the transition function of $T$. For the second type of violations, ${\mathcal{A}}$ may guess, in each step, that the next three letters encode a position in a configuration and the letter to come $n_{0}$ letters later, namely at the same position in the successive configuration, is different from the one that should appear in a legal encoding of two successive configurations. If a violation is detected, ${\mathcal{A}}$ moves to an accepting sink. Otherwise, ${\mathcal{A}}$ returns to the initial state and $w$ gets another chance to be accepted.
In case ${\mathcal{A}}$ guesses that C2 is satisfied, it essentially waits to see a letter that encodes the accepting state of $T$, and when such a letter arrives it moves to the accepting sink.
Also, whenever ${\mathcal{A}}$ waits to witness some behavior, namely, waits for a position where there is a violation of the encoding, or waits for an encoding of an accepting configuration of $T$ that uses $n_{0}$ cells, it may nondeterministically, upon reading the next letter, return to the initial state. It is not hard to see that ${\mathcal{A}}$ can be defined in size linear in $T$ and $n_{0}$. As the only accepting states of ${\mathcal{A}}$ is the accepting sink, it is clearly weak, and in fact describes a co-safety language.
We prove that $T$ accepts the empty tape iff ${\mathcal{A}}$ is SD. First, if $T$ rejects the empty tape, then ${\mathcal{A}}$ is not SD. To see this, consider the word $w_{\varepsilon}$ that encodes the computation of $T$ on the empty tape, and let $w_{\varepsilon}^{\prime}$ be a word that is obtained from $w_{\varepsilon}$ by making a single violation in the first letter. Note that $w_{\varepsilon}^{\prime}\in L({\mathcal{A}})$ since it has a violation. Note also that any proper suffix of $w_{\varepsilon}^{\prime}$ is also a suffix of $w_{\varepsilon}$, and hence is not in $L({\mathcal{A}})$ as it encodes a suffix of a computation of $T$ that uses at most $n_{0}$ tape cells and does not include an encoding of an accepting configuration of $T$. Consequently, the word $w_{\varepsilon}^{\prime}$ can be accepted by ${\mathcal{A}}$ only by guessing a violation that is caused by the first letter. In particular, if we guess incorrectly that there is a violation on the second letter of $w^{\prime}_{\varepsilon}$, then that guess fails, and after it we cannot branch to an accepting run. This shows that ${\mathcal{A}}$ is not SD.
For the other direction, we show that if $T$ accepts the empty tape, then all the states of ${\mathcal{A}}$ are universal. In particular, all the states of ${\mathcal{A}}$ are equivalent, which clearly implies that ${\mathcal{A}}$ is SD. First, as argued above, if $T$ accepts the empty tape, then ${\mathcal{A}}$ is universal. In addition, by the definition of ${\mathcal{A}}$, for every infinite word $w$ and for all states $q$ of ${\mathcal{A}}$ that are not the accepting sink, there is a path from $q$ to the initial state that is labeled by a prefix of $w$. Thus, the language of all states is universal, and we are done.
Thus, we conclude that $T$ accepts the empty tape iff ${\mathcal{A}}$ is SD. In Appendix B, we give the full technical details of the construction of ${\mathcal{A}}$.
Theorem 4.
The problem of deciding whether a given NBW or HD-NBW is DBP is NP-complete.
Proof 4.2.
For membership in NP, observe we can check that a witness deterministic pruning ${\mathcal{A}}^{\prime}$ is equivalent to ${\mathcal{A}}$ by checking whether $L({\mathcal{A}})\subseteq L({\mathcal{A}}^{\prime})$. Since ${\mathcal{A}}^{\prime}$ is deterministic, the latter can be checked in polynomial time. For NP-hardness, we describe a parsimonious polynomial time reduction from SAT. That is, given a CNF formula $\varphi$, we construct an HD-NBW ${\mathcal{A}}_{\varphi}$ such that there is a bijection between assignments to the variables of $\varphi$ and DBWs embodied in ${\mathcal{A}}_{\varphi}$, and an assignment satisfies $\varphi$ iff its corresponding embodied DBW is equivalent to ${\mathcal{A}}_{\varphi}$. In particular, $\varphi$ is satisfiable iff ${\mathcal{A}}_{\varphi}$ is DBP.
Consider a SAT instance $\varphi$ over the variable set $X=\{x_{1},\ldots,x_{n}\}$ and with $m\geq 1$ clauses $C=\{c_{1},\ldots,c_{m}\}$. For $n\geq 1$, let $[n]=\{1,2,\ldots,n\}$. For a variable $x_{k}\in X$, let $C^{0}_{k}\subseteq C$ be the set of clauses in which $x_{k}$ appears negatively, and let $C^{1}_{k}\subseteq C$ be the set of clauses in which $x_{k}$ appears positively. For example, if $c_{1}=x_{1}\lor\lnot x_{2}\lor x_{3}$, then $c_{1}$ is in $C^{1}_{1}$, $C_{2}^{0}$, and $C^{1}_{3}$. Assume that all clauses depend on at least two different variables (that is, no clause is a tautology or forces an assignment to a single variable). Let $\Sigma_{n,m}=X\cup C$, and let
$R_{n,m}=(X\cdot C)^{*}\cdot\{x_{1}\cdot c_{j}\cdot x_{2}\cdot c_{j}\cdots x_{n}\cdot c_{j}:j\in[m]\}\subseteq\Sigma^{*}_{n,m}.$
We construct an HD-NBW ${\mathcal{A}}_{\varphi}$ that recognizes $L_{n,m}=(R_{n,m})^{\omega}$, and is DBP iff $\varphi$ is satisfiable.
Let $D_{n,m}$ be a DFW that recognizes $R_{n,m}$ with $O(n\cdot m)$ states, a single accepting state $p$, and an initial state $q_{0}$ that is visited only once in all runs. For example, we can define $D_{n,m}=\langle\Sigma_{n,m},Q_{n,m},q_{0},\delta_{n,m},\{p\}\rangle$ as follows: from $q_{0}$, the DFW expects to read only words in $(X\cdot C)^{*}$ – upon a violation of this pattern, it goes to a rejecting sink. Now, if the pattern is respected, then with $X\setminus\{x_{1}\}$, the DFW goes to two states where it loops with $C\cdot(X\setminus\{x_{1}\})$ and, upon reading $x_{1}$ from all states that expect to see letters in $X$, it branches with each $c_{j}$, for all $j\in[m]$, to a path where it hopes to detect an $x_{2}\cdot c_{j}\cdots x_{n}\cdot c_{j}$ suffix. If the detection is completed successfully, it goes to the accepting state $p$. Otherwise, it returns to the two-state loop. Now, we define ${\mathcal{A}}_{\varphi}=\langle\Sigma_{n,m},Q_{\varphi},p,\delta_{\varphi},\{q_{0}\}\rangle$, where $Q_{\varphi}=Q_{n,m}\cup\{q^{i}_{k}:(i,k)\in\{0,1\}\times[n]\}$. The idea behind ${\mathcal{A}}_{\varphi}$ is as follows. From state $p$ (that is, the accepting state of $D_{n,m}$, which is now the initial state of ${\mathcal{A}}_{\varphi}$), the NBW ${\mathcal{A}}_{\varphi}$ expects to read a letter in $X$. When it reads $x_{k}$, for $1\leq k\leq n$, it nondeterministically branches to the states $q_{k}^{0}$ and $q_{k}^{1}$. Intuitively, when it branches to $q_{k}^{0}$, it guesses that the clause that comes next is one that is satisfied when $x_{k}=0$, namely a clause in $C^{0}_{k}$. Likewise, when it branches to $q_{k}^{1}$, it guesses that the clause that comes next is one that is satisfied when $x_{k}=1$, namely a clause in $C^{1}_{k}$. When the guess is successful, ${\mathcal{A}}_{\varphi}$ moves to the $\alpha$-state $q_{0}$. When the guess is not successful, it returns to $p$. Implementing the above intuition, transitions from the states $Q_{n,m}\setminus\{p\}$ are inherited from $D_{n,m}$, and transitions from the states in $\{q^{i}_{k}:(i,k)\in\{0,1\}\times[n]\}\cup\{p\}$ are defined as follows (see also Figure 3).
  • For all $k\in[n]$, we have that $\delta_{\varphi}(p,x_{k})=\{q^{0}_{k},q^{1}_{k}\}$.
  • For all $k\in[n]$, $i\in{\{0,1\}}$, and $j\in[m]$, if $c_{j}\in C_{k}^{i}$, then $\delta_{\varphi}(q^{i}_{k},c_{j})=\{q_{0}\}$. Otherwise, $\delta_{\varphi}(q^{i}_{k},c_{j})=\{p\}$. For example, if $c_{1}=x_{1}\lor\lnot x_{2}\lor x_{3}$, then $\delta_{\varphi}(q^{0}_{2},c_{1})=\{q_{0}\}$ and $\delta_{\varphi}(q^{1}_{2},c_{1})=\{p\}$.
Figure 3. The transitions to and from the states $q_{k}^{0}$ and $q_{k}^{1}$ in ${\mathcal{A}}_{\varphi}$.
Note that $p$ is the only nondeterministic state of ${\mathcal{A}}_{\varphi}$ and that for every deterministic pruning of ${\mathcal{A}}_{\varphi}$, all the words in $(X\cdot C)^{\omega}$ have an infinite run in the pruned automaton. This run, however, may eventually loop in $\{p\}\cup\{q^{0}_{k},q^{1}_{k}:k\in[n]\}$. Note also, that for readability purposes, the automaton ${\mathcal{A}}_{\varphi}$ is not total. Specifically, the states of ${\mathcal{A}}_{\varphi}$ are partitioned into states that expect to see letters in $X$ and states that expect to see letters in $C$. In particular, all infinite paths in ${\mathcal{A}}_{\varphi}$ are labeled by words in $(X\cdot C)^{\omega}$. Thus, when defining an HD strategy $g$ for ${\mathcal{A}}_{\varphi}$, we only need to define $g$ on prefixes in $(X\cdot C)^{*}\cup(X\cdot C)^{*}\cdot X$.
In the following propositions, we prove that ${\mathcal{A}}_{\varphi}$ is an HD NBW recognizing $L_{n,m}$, and that ${\mathcal{A}}_{\varphi}$ is DBP iff $\varphi$ is satisfiable.
Proposition 5.
$L({\mathcal{A}}_{\varphi})\subseteq L_{n,m}$.
Proof 4.3.
As already mentioned, all infinite paths of ${\mathcal{A}}_{\varphi}$, accepting or rejecting, are labeled by words in $(X\cdot C)^{\omega}$. Further, any accepting run of ${\mathcal{A}}_{\varphi}$ has infinitely many sub-runs that are accepting finite runs of $D_{n,m}$. Let $\infty R_{n,m}$ be the language of infinite words that include infinitely many disjoint finite subwords in $R_{n,m}$. Since $L_{n,m}=(R_{n,m})^{\omega}=(X\cdot C)^{\omega}\cap(\infty R_{n,m})$, it follows that $L({\mathcal{A}}_{\varphi})\subseteq L_{n,m}$.
Proposition 6.
There exists a strategy $g:\Sigma^{*}\rightarrow Q_{\varphi}$ for ${\mathcal{A}}_{\varphi}$ that accepts all words in $L_{n,m}$. Formally, for all $w\in L_{n,m}$, the run $g(w)=g(w[1,0]),g(w[1,1]),g(w[1,2]),\ldots$, is an accepting run of ${\mathcal{A}}_{\varphi}$ on $w$.
Proof 4.4.
The definition of $L_{n,m}$ is such that when reading a prefix that ends with a subword of the form $x_{1}\cdot c_{j}$, for some $j\in[m]$, then we can guess that the word continues with $x_{2}\cdot c_{j}\cdot x_{3}\cdot c_{j}\cdots x_{n}\cdot c_{j}$; thus that $c_{j}$ is the clause that is going to repeat. Therefore, when we are at state $p$ after reading a word that ended with $x_{1}\cdot c_{j}$, and we read $x_{2}$, it is a good HD strategy to move to a state $q^{i}_{2}$ such that the assignment $x_{2}=i$ satisfies $c_{j}$ (if such $i\in\{0,1\}$ exists; otherwise the strategy can choose arbitrary between $q^{0}_{2}$ and $q^{1}_{2}$), and if the run gets back to $p$, the strategy continues with assignments that hope to satisfy $c_{j}$, until the run gets to $q_{0}$ or another occurrence of $x_{1}$ is detected. Note that while it is not guaranteed that for all $k\in[n]$ there is $i\in\{0,1\}$ such that the assignment $x_{k}=i$ satisfies $c_{j}$, it is guaranteed that such an $i$ exists for at least two different $k$’s (we assume that all clauses depend on at least two variables). Thus, even though we a priori miss an opportunity to satisfy $c_{j}$ with an assignment to $x_{1}$, it is guaranteed that there is another $2\leq k\leq n$ such that $c_{j}$ can be satisfied by $x_{k}$. We define $g$ inductively as follows. Recall that ${\mathcal{A}}_{\varphi}$ is nondeterministic only in the state $p$, and so in all other states, the strategy $g$ follows the only possible transition. First, for all $k\in[n]$, we define $g(x_{k})=q^{0}_{k}$. Let $v\in(X\cdot C)^{*}\cdot X$, be such that $g$ has already been defined on $v$ and let $j\in[m]$. Since $v\notin(X\cdot C)^{*}$, we have that $g(v)\neq p$ and so $g(v\cdot c_{j})$ is uniquely defined. We continue and define $g$ on $u=v\cdot c_{j}\cdot x_{k}$, for all $k\in[n]$. If $g(v\cdot c_{j})\neq p$, then $g(u)$ is uniquely defined. Otherwise, $g(v\cdot c_{j})=p$ and we define $g(u)$ as follows,
  • If $k=1$, then we define $g(u)=q^{0}_{1}$.
  • If $k>1$ and $x_{k}$ participates in $c_{j}$, then we define $g(u)=q^{i}_{k}$, where $i\in\{0,1\}$ is such that $c_{j}\in C^{i}_{k}$. Note that since we assume that $c_{j}$ is not a tautology, then $i$ is uniquely defined.
  • If $k>1$ and $x_{k}$ does not participate in $c_{j}$, then the value of $c_{j}$ is not affected by the assignment to $x_{k}$, and in that case we define $g(u)=q^{0}_{k}$.
The reason for the distinction between the cases $k=1$ and $k>1$ is that when we see a finite word that ended with $c_{j}\cdot x_{1}$, then there is no special reason to hope that the next letter is going to be $c_{j}$. This is in contrast, for example, to the case we have seen a word that ends with $c_{j^{\prime}}\cdot x_{1}\cdot c_{j}\cdot x_{2}$, where it is worthwhile to guess we are about to see $c_{j}$ as the next letter. By the definition of $g$, it is consistent with $\Delta_{\varphi}$. In Appendix C we formally prove that $g$ is a winning HD strategy for ${\mathcal{A}}_{\varphi}$. Namely, that for all $w\in L_{n,m}$, the run $g(w)$ on $w$, generated by $g$ is accepting.
We now examine the relation between prunings of ${\mathcal{A}}_{\varphi}$ and assignments to $\varphi$, and conclude the proof by proving the following:
Proposition 7.
The formula $\varphi$ is satisfiable iff the HD-NBW ${\mathcal{A}}_{\varphi}$ is DBP.
Proof 4.5.
Consider an assignment $i_{1},\ldots,i_{n}\in\{0,1\}$, for $X$. I.e., $x_{k}=i_{k}$ for all $k\in[n]$. A possible memoryless HD strategy is to always move from $p$ to $q^{i_{k}}_{k}$ when reading $x_{k}$. This describes a one to one correspondence between assignments for $X$ and prunings of ${\mathcal{A}}_{\varphi}$. Assume that the assignment $i_{k}\in\{0,1\}$, for $k\in[n]$, satisfies $\varphi$. Then, the corresponding pruning recognizes $L_{n,m}$. Indeed, instead of trying to satisfy the last read clause $c_{j}$, we may ignore this extra information, and rely on the fact that one of the assignments $x_{k}=i_{k}$ is going to satisfy $c_{j}$. In other words, the satisfiability of $\varphi$ allows us to ignore the history and still accept all words in $L_{n,m}$, which makes ${\mathcal{A}}_{\varphi}$ DBP. On the other hand, if an assignment does not satisfy some clause $c_{j}$, then the corresponding pruning fails to accept the word $(x_{1}\cdot c_{j}\cdots x_{n}\cdot c_{j})^{\omega}$, which shows that if $\varphi$ is not satisfiable then ${\mathcal{A}}_{\varphi}$ is not DBP. In Appendix C we formally prove that there is a one to one correspondence between prunings of ${\mathcal{A}}_{\varphi}$ and assignments to $\varphi$, and that an assignment satisfies $\varphi$ iff the corresponding pruning recognizes $L_{n,m}$, implying Proposition 7.
We continue to co-Büchi automata. In [16], the authors prove that deciding the DBPness of a given NCW is NP-complete. Below we extend their proof to HD-NCWs.
Theorem 8.
The problem of deciding whether a given HD-NCW is DBP is NP-complete.
Proof 4.6.
The upper bound is as in [16]. For the lower bound, we analyze the reduction in [16] and prove that the NCW constructed there is HD. The reduction is from the Hamiltonian-cycle problem: given a connected graph $G=\langle[n],E\rangle$, the reduction outputs an NCW ${\mathcal{A}}_{G}$ over the alphabet $[n]$ that is obtained from $G$ by adding self loops to all vertices, labelling the loop at a vertex $i$ by the letter $i$, and labelling the edges from vertex $i$ to all its neighbours in $G$ by every letter $j\neq i$. Then, the co-Büchi condition requires a run to eventually get stuck at a self-loop [^7] The exact reduction is more complicated and involves an additional letter #normal-#\## that forces each deterministic pruning of ${\mathcal{A}}_{G}$ to proceed to the same neighbour of $i$ upon reading a letter $j\neq i$ from the vertex $i$.. Accordingly, $L({\mathcal{A}}_{G})=[n]^{*}\cdot\bigcup_{i\in[n]}i^{\omega}$.
It is not hard to see that ${\mathcal{A}}_{G}$ is HD. Indeed, an HD strategy can decide to which neighbour of $i$ to proceed with a letter $j\neq i$ by following a cycle $c$ that traverses all the vertices of the graph $G$. Since when we read $j\neq i$ at vertex $i$ we move to a neighbour state, then by following the cycle $c$ upon reading $i^{\omega}$, we eventually reach the vertex $i$ and get stuck at the $i$-labeled loop. Thus, the Hamiltonian-cycle problem is reduced to DBPness of an HD-NCW, and we are done.
5. Approximated Determinization by Pruning
Consider a nondeterministic automaton ${\mathcal{A}}$. We say that ${\mathcal{A}}$ is almost-DBP if it can be pruned to a deterministic automaton that retains almost all the words in $L({\mathcal{A}})$. [^8] Not to be confused with the definition of almost-DBPness in [2], where it is used to describe automata that are deterministic in their transition function yet may have a set of initial states. Formally, an NXW ${\mathcal{A}}$ is almost-DBP if there is a DXW ${\mathcal{A}}^{\prime}$ embodied in ${\mathcal{A}}$ such that $\mathbb{P}(L({\mathcal{A}})\setminus L({\mathcal{A}}^{\prime}))=0$. Thus, while ${\mathcal{A}}^{\prime}$ need not accept all the words accepted by ${\mathcal{A}}$, it rejects only a negligible set of words in $L({\mathcal{A}})$. Note that if ${\mathcal{A}}$ is DBP or $\mathbb{P}(L({\mathcal{A}}))=0$, then ${\mathcal{A}}$ is almost-DBP. Indeed, if ${\mathcal{A}}$ is DBP, then the DXW ${\mathcal{A}}^{\prime}$ embodies in ${\mathcal{A}}$ with $L({\mathcal{A}}^{\prime})=L({\mathcal{A}})$ is such that $L({\mathcal{A}})\setminus L({\mathcal{A}}^{\prime})=\emptyset$, and if $\mathbb{P}(L({\mathcal{A}}))=0$, then for every DBW ${\mathcal{A}}^{\prime}$ embodied in ${\mathcal{A}}$, we have that $\mathbb{P}(L({\mathcal{A}})\setminus L({\mathcal{A}}^{\prime}))=0$.
In this section we study approximated determinization by pruning. We first show that, unsurprisingly, almost-DBPness is not a trivial property:
Theorem 9.
There is an NWW (and hence, also an NBW and an NCW) that is not almost-DBP.
Proof 5.1.
Consider the NWW ${\mathcal{A}}_{1}$ in Figure 4. It is not hard to see that $L({\mathcal{A}}_{1})=\{a,b\}^{\omega}$, and so $\mathbb{P}(L({\mathcal{A}}_{1}))=1$. Moreover, every deterministic pruning of ${\mathcal{A}}_{1}$ is such that $q_{rej}$ is reachable from all states, which implies that $\{q_{rej}\}$ is the only ergodic SCC of any pruning. Since $\{q_{rej}\}$ is $\alpha$-free, it follows that every deterministic pruning of ${\mathcal{A}}_{1}$ recognizes a language of measure zero, and hence ${\mathcal{A}}_{1}$ is not almost-DBP. As an example, consider the deterministic pruning ${\mathcal{A}}^{\prime}_{1}$ described on the right hand side of Figure 4. The only ergodic SCC of ${\mathcal{A}}^{\prime}_{1}$ is $\alpha$-free, and as such $\mathbb{P}(L({\mathcal{A}}^{\prime}_{1}))=0$.
Figure 4. An NWW that is not almost-DBP.
We continue and study the almost-DBPness of HD and SD automata. We show that while for Büchi (and hence also weak) automata, semantic determinism implies almost-DBPness, thus every SD-NBW is almost-DBP, for co-Büchi automata semantic determinism is not enough, and we need HDness. Thus, there is an SD-NCW that is not almost-DBP, yet all HD-NCWs are almost-DBP.
[…]
  1. (1)
    Random picks a letter $\sigma\in\Sigma$ uniformly, and the game moves to position $(q,\sigma)$.
  2. (2)
    Eve picks a transition $(q,\sigma,p)\in\Delta$, and the game moves to position $p$.
A probabilistic strategy for Eve is $f:(Q\times\Sigma)^{+}\rightarrow[0,1]^{Q}$, where for all histories $x\in(Q\times\Sigma)^{*}$ and positions of Eve $(q,\sigma)\in Q\times\Sigma$, the function $d=f(x\cdot(q,\sigma)):Q\rightarrow[0,1]$, is a distribution on $Q$ such that $d(p)\neq 0$ implies that $p\in\delta(q,\sigma)$. As usual, we say that a strategy $f$ is memoryless, if it depends only on the current position, thus for all histories $x,y\in(Q\times\Sigma)^{*}$ and positions of Eve $(q,\sigma)\in Q\times\Sigma$, it holds that $f(x\cdot(q,\sigma))=f(y\cdot(q,\sigma))$. A strategy for Eve is pure if for all histories $x\in(Q\times\Sigma)^{*}$ and positions of Eve $(q,\sigma)\in Q\times\Sigma$, there is a position $p\in\delta(q,\sigma)$ such that $f(x\cdot(q,\sigma))(p)=1$. When Eve plays according to a strategy $f$, the outcome of the game can be viewed as a run $r_{f}=q_{f}^{0},q_{f}^{1},q_{f}^{2},\ldots$ in ${\mathcal{A}}$, over a random word $w_{f}\in\Sigma^{\omega}$. (The word that is generated in a play is independent of the strategy of Eve, but we use the notion $w_{f}$ to emphasize that we are considering the word that is generated in a play where Eve plays according to $f$).
Let $Q_{rej}=\{q\in Q:\mathbb{P}(L({\mathcal{A}}^{q}))=0\}$. The outcome $r_{f}$ of the game is winning for Eve iff $r_{f}$ is accepting, or $r_{f}$ visits $Q_{rej}$. Note also that for all positions $q\in Q_{rej}$ and $p\in Q$, if $p$ is reachable from $q$, then $p\in Q_{rej}$. Hence, the winning condition can be defined by the Büchi objective $\alpha\cup Q_{rej}$. Note that $r_{f}$ is winning for Eve iff $\inf(r_{f})\subseteq Q_{rej}$ or $\inf(r_{f})\cap\alpha\neq\emptyset$. We say that $f$ is an almost-sure winning strategy, if $r_{f}$ is winning for Eve with probability $1$, and Eve almost-sure wins in ${\mathcal{G}}_{\mathcal{A}}$ if she has an almost-sure winning strategy.
Theorem 10.
All SD-NBWs are almost-DBP.
Proof 5.2.
Let ${\mathcal{A}}$ be an SD-NBW, and consider the simple stochastic Büchi game $\mathcal{G}_{\mathcal{A}}$. We first show that Eve has a probabilistic strategy to win ${\mathcal{G}}_{\mathcal{A}}$ with probability $1$, even without assuming that ${\mathcal{A}}$ is semantically deterministic. Consider the probabilistic strategy $g$ where from $(q,\sigma)$, Eve picks one of the $\sigma$-successors of $q$ uniformly by random. Note that this strategy is memoryless, and hence the outcome of the game can be thought as a random walk in ${\mathcal{A}}$ that starts at $q_{0}$ and gives positive probabilities to all transitions. Thus, with probability $1$, the run $r_{g}$ is going to reach an ergodic SCC of ${\mathcal{A}}$ and visit all its states. If the run $r_{g}$ reaches an $\alpha$-free ergodic SCC, then $\inf(r_{g})\subseteq Q_{rej}$, and hence $r_{g}$ is then winning for Eve. Otherwise, $r_{g}$ reaches a non $\alpha$-free ergodic SCC, and with probability $1$, it visits all the states in that SCC. Thus, with probability $1$, we have $\inf(r_{g})\cap\alpha\neq 0$, and $r_{g}$ is winning for Eve. Overall, Eve wins ${\mathcal{G}}_{\mathcal{A}}$ with probability $1$ when playing according to $g$.
For a strategy $f$ for Eve, we say that $r_{f}$ is correct if $w_{f}\in L({\mathcal{A}})$ implies that $r_{f}$ is accepting. Note that $w_{f}\notin L({\mathcal{A}})$ always implies that $r_{f}$ is rejecting. We show that if ${\mathcal{A}}$ is SD, then for every winning strategy $f$ for Eve, we have that $r_{f}$ is correct with probability $1$. Consider a winning strategy $f$ for Eve. Since $f$ is winning for Eve, we have that with probability 1, either $r_{f}$ is an accepting run or $\inf(r_{f})\subseteq Q_{rej}$. Thus, it is sufficient to prove that if $\inf(r_{f})\subseteq Q_{rej}$, then $w_{f}\in L({\mathcal{A}})$ with probability 00, since otherwise, almost surly $r_{f}$ is accepting, and in particular is correct.
By semantic determinism, for all $i\geq 0$, it holds that $w_{f}\in L({\mathcal{A}})$ iff $w_{f}[i+1,\infty]\in L({\mathcal{A}}^{q_{f}^{i}})$. Moreover, the infix $w_{f}[i+1,\infty]$ is independent of $q^{i}_{f}$, and hence for all $q\in Q$ and $i\geq 0$, the event $w_{f}[i+1,\infty]\in L({\mathcal{A}}^{q})$ is independent of $q^{i}_{f}$. Thus, for all $q\in Q$ and $i\geq 0$, it holds that $\mathbb{P}(w_{f}\in L({\mathcal{A}})\>|\>q_{f}^{i}=q)=\mathbb{P}(w_{f}[i+1,\infty]\in L({\mathcal{A}}^{q}))=\mathbb{P}(L({\mathcal{A}}^{q}))$. Hence, by the definition of $Q_{rej}$, and by the fact that $Q_{rej}$ is finite, we have that $\mathbb{P}(w_{f}\in L({\mathcal{A}})\>|\>q_{f}^{i}\in Q_{rej})=0$ for all $i\geq 0$, and so $\mathbb{P}(w_{f}\in L({\mathcal{A}})\>|\>r_{f}\text{ visits }Q_{rej})=0$. Overall, we showed that if $f$ is winning for Eve and ${\mathcal{A}}$ is SD, then $\mathbb{P}(r_{f}\text{ is correct})=1$.
Hence, by pure memoryless determinacy of simple stochastic parity games [10], we may consider a pure memoryless winning strategy $f$ for Eve in ${\mathcal{G}}_{\mathcal{A}}$, and by the above we know that a random walk in the corresponding pruned DBW is correct with probability 1. Formally, since $f$ is pure memoryless, it induces a pruning of ${\mathcal{A}}$. We denote this pruning by ${\mathcal{A}}^{f}$, and think of $r_{f}$ as a random walk in ${\mathcal{A}}^{f}$. As we saw, since $f$ is a winning strategy and ${\mathcal{A}}$ is SD, we have that $r_{f}$ is correct with probability 1. Note that $\mathbb{P}(L({\mathcal{A}})\setminus L({\mathcal{A}}^{f}))$, is precisely the probability that a random word $w_{f}$ is in $L({\mathcal{A}})$ but not accepted by ${\mathcal{A}}^{f}$. Namely, the probability that $r_{f}$ is not correct. Hence, $\mathbb{P}(L({\mathcal{A}})\setminus L({\mathcal{A}}^{f}))=0$, and ${\mathcal{A}}$ is almost-DBP.
We continue to co-Büchi automata and show that unlike the case of Büchi, here semantic determinism does not imply almost-DBPness. Note that the NWW in Figure 4 is not SD.
Theorem 11.
There is an SD-NCW that is not almost-DBP.
Proof 5.3.
Consider the NCW ${\mathcal{A}}_{2}$ in Figure 5. It is not hard to see that $L({\mathcal{A}}_{2})=\{a,b\}^{\omega}$, and hence $\mathbb{P}(L({\mathcal{A}}_{2}))=1$. In fact all the states $q$ of ${\mathcal{A}}_{2}$ have $L({\mathcal{A}}_{2}^{q})=\{a,b\}^{\omega}$, and so it is semantically deterministic.
Figure 5. An SD-NCW that is not almost-DBP.
Moreover, every deterministic pruning of ${\mathcal{A}}_{2}$ is strongly connected and not $\alpha$-free. It follows that any deterministic pruning of ${\mathcal{A}}_{2}$ recognizes a language of measure zero, and hence ${\mathcal{A}}_{2}$ is not almost-DBP. As an example, consider the deterministic pruning ${\mathcal{A}}^{\prime}_{2}$ described on the right hand side of Figure 5. It is easy to see that ${\mathcal{A}}^{\prime}_{2}$ is strongly connected and not $\alpha$-free, and as such, $\mathbb{P}(L({\mathcal{A}}^{\prime}_{2}))=0$. Thus, ${\mathcal{A}}_{2}$ is not almost-DBP.
[…]
Lemma 12.
If $L$ is NCW-recognizable, then $\mathbb{P}(L\setminus\mbox{\it co-safe}(L))=0$.
Proof 5.4.
Consider an NCW-recognizable language $L$. Since NCW=DCW, there is a DCW ${\mathcal{D}}$ that recognizes $L$. Assume without loss of generality that ${\mathcal{D}}$ has a single state $q$ with $L({\mathcal{D}}^{q})=\Sigma^{\omega}$, in particular, $C=\{q\}$ is the only ergodic $\alpha$-free SCC of ${\mathcal{D}}$. Then, for every word $w\in\Sigma^{\omega}$, we have that $w\in\mbox{\it co-safe}(L)$ iff the run of ${\mathcal{D}}$ on $w$ reaches $C$. Hence, the probability that $w\in L\setminus\mbox{\it co-safe}(L)$ equals the probability that $inf(r)$ is $\alpha$-free but is not an ergodic SCC of ${\mathcal{D}}$. Since the later happens w.p 0, we have that $\mathbb{P}(L\setminus\mbox{\it co-safe}(L))=0$.
[…]
Theorem 13.
All HD-NCWs are almost-DBP.
Proof 5.5.
Let ${\mathcal{A}}=\langle\Sigma,Q,q_{0},\delta,\alpha\rangle$ be an HD-NCW. Consider the NCW ${\mathcal{A}}^{\prime}=\langle\Sigma,Q,q_{0},\delta,\alpha^{\prime}\rangle$, where $\alpha^{\prime}=\alpha\cup\{q\in Q:L({\mathcal{A}}^{q})\neq\Sigma^{\omega}\}$. We prove that ${\mathcal{A}}^{\prime}$ is an HD-NCW with $L({\mathcal{A}}^{\prime})=\mbox{\it co-safe}(L({\mathcal{A}}))$. Consider a word $w\in L({\mathcal{A}}^{\prime})$, and let $r=q_{0},q_{1},q_{2},\ldots$ be an accepting run of ${\mathcal{A}}^{\prime}$ on $w$. There exists a prefix $x\in\Sigma^{*}$ of $w$ such that $r$ reaches some $q\notin\alpha^{\prime}$ when reading $x$. Hence $L({\mathcal{A}}^{q})=\Sigma^{\omega}$, and so $x\cdot\Sigma^{\omega}\subseteq L({\mathcal{A}})$. That is, $x$ is a good prefix and $w\in\mbox{\it co-safe}(L({\mathcal{A}}))$. Thus, $L({\mathcal{A}}^{\prime})\subseteq\mbox{\it co-safe}(L({\mathcal{A}}))$.
[…]
Since $w\in\mbox{\it co-safe}(L({\mathcal{A}}))$, it has a good prefix $x\in\Sigma^{*}$. Since $f$ is an HD strategy and $x$ is a good prefix, it must be that $L({\mathcal{A}}^{f(x)})=\Sigma^{\omega}$. Also, since all the extensions of $x$ are also good prefixes, the above holds also for all states that $f$ visits after $f(x)$. Hence, the run $r$ visits states $q$ with $L({\mathcal{A}}^{q})\neq\Sigma^{\omega}$ only during the finite prefix $x$. Thus, $\inf(r)\cap\alpha^{\prime}=\emptyset$, and we are done.
[…]
6. Nondeterminism and Reasoning about MDPs
A probabilistic open system can be modeled by a Markov decision process (MDP, for short). As we define formally below, an MDP describes the behavior of the system when it interacts with its environment. The behavior is probabilistic: each input provided by the environment induces a probability on the reaction of the system. Reasoning about the behavior of an MDP can proceed by taking its product with a deterministic automaton that describes a desired behavior for the MDP. Indeed, this product contains information on the probability that the interaction results in a behavior accepted by the automaton. When the automaton is nondeterministic, this information is lost, and thus reasoning about MDPs involves determinization. An automaton is said to be good for MDPs (GFM), if its product with MDPs maintains the probability of acceptance, and can therefore replace deterministic automata when reasoning about stochastic behaviors [13, 30]. In Section 5, we related semantic determinism with almost-DBPness. In this section we relate semantic determinism with GFMness.
6.1. Good-for-MDP automata
We first define MDPs and good-for-MDP automata. A $\Sigma$-labeled Markov decision process (MDP) is ${\mathcal{M}}=\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{{\mathcal{M}}},\Sigma,\tau\rangle$, where $S$ is a finite set of states, $s_{0}\in S$ is an initial state, and for every $s\in S$, the set $A_{s}$ is a finite set of actions that are available in $s$. Let $A=\bigcup_{s\in S}A_{s}$ be the set of all actions. Then, $\mathrm{Pr}_{\mathcal{M}}:S\times A\times S\nrightarrow[0,1]$ is a (partial) stochastic transition function: for every two states $s,s^{\prime}\in S$ and action $a\in A_{s}$, we have that $\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})$ is the probability of moving from $s$ to $s^{\prime}$ when action $a$ is taken. Accordingly, for every $s\in S$ and $a\in A_{s}$, we have $\Sigma_{s^{\prime}\in S}\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})=1$. Finally, $\Sigma$ is a finite alphabet, and $\tau:S\rightarrow\Sigma$ is a labeling function on the states.
Consider a nondeterministic automaton ${\mathcal{A}}$ with alphabet $\Sigma$, and a $\Sigma$-labeled MDP ${\mathcal{M}}$, and consider the following game between a player and nature. The positions of the game are pairs $\langle s,q\rangle$, where $s\in S$ is a state of ${\mathcal{M}}$, and $q\in Q$ is a state of ${\mathcal{A}}$. The game starts by the player picking an action $a\in A_{s_{0}}$ available from the initial state of ${\mathcal{M}}$, and the game moves to $\langle s,q_{0}\rangle$ with probability $\mathrm{Pr}_{\mathcal{M}}(s_{0},a,s)$. Then, at each round from position $\langle s,q\rangle$, the player chooses an action $a\in A_{s}$ and a state $q^{\prime}\in\delta(q,\tau(s))$, and then nature chooses a state $s^{\prime}\in S$ with probability $\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})$, and the game moves to $\langle s^{\prime},q^{\prime}\rangle$. The objective of the player is to maximize the probability of generating an accepting run of ${\mathcal{A}}$. When the automaton ${\mathcal{A}}$ is deterministic, the player has to only pick actions, as the state $q^{\prime}\in\delta(q,\tau(s))$ is unique. When the automaton is nondeterministic, the player may fail to generate an accepting run of ${\mathcal{A}}$ even when the actions he provides induce a word in $L({\mathcal{A}})$. Indeed, each nondeterministic choice may lead to accepting only a subset of the words in $L({\mathcal{A}})$. Intuitively, ${\mathcal{A}}$ is Good-For-MDP, if for every MDP ${\mathcal{M}}$, the nondeterminism in ${\mathcal{A}}$ does not prevent the player from maximizing the probability of generating an accepting run. We now formalize this intuition.
A run of ${\mathcal{M}}$ is an infinite sequence of states $r=\langle s_{0},s_{1},s_{2},\ldots\rangle\in S^{\omega}$, such that for all $i\geq 0$ there exists $a\in A_{s_{i}}$ with $\mathrm{Pr}_{\mathcal{M}}(s_{i},a,s_{i+1})>0$. A finite run is a prefix of such a sequence. Let $\Omega({\mathcal{M}})$ denote the set of runs of ${\mathcal{M}}$ from the initial state $s_{0}$, and let $\Pi({\mathcal{M}})$ denote the set of finite such runs $\langle s_{0},s_{1},\ldots,s_{k}\rangle\in S^{k+1}$, for some $k\geq 0$. For a run $r\in\Omega({\mathcal{M}})$, we define the corresponding labeled run as $\tau(r)=\tau(s_{1}),\tau(s_{2}),\tau(s_{3}),\ldots\in\Sigma^{\omega}$. Note that the label of the first state $s_{0}$ of a run is not part of $\tau(r)$.
Consider the set $A$ of actions. Let $\mathrm{dist}(A)$ be the set of distributions over $A$. A distribution $d\in\mathrm{dist}(A)$ is a point distribution if there exists $a\in A$ such that $d(a)=1$. A strategy for an MDP ${\mathcal{M}}$ is a function $\mu:\Pi({\mathcal{M}})\rightarrow\mathrm{dist}(A)$, which suggests to the player a distribution on the available actions given the history of the game so far. The strategy should suggest an available action, thus $\mu(s_{0},s_{1},s_{2},\ldots,s_{k})(a)>0$ implies that $a\in A_{s_{k}}$. Let $\Pi({\mathcal{M}},\mu)$ denote the subset of $\Pi({\mathcal{M}})$ that includes exactly all finite runs that agree with the strategy $\mu$. Formally, we have that $\langle s_{0}\rangle\in\Pi({\mathcal{M}},\mu)$, and for all $p^{\prime}=\langle s_{0},s_{1},\ldots,s_{k}\rangle\in\Pi({\mathcal{M}},\mu)$, and $s_{k+1}\in S$, we have that $p=p^{\prime}\cdot s_{k+1}\in\Pi({\mathcal{M}},\mu)$ iff there exists $a_{k+1}\in A_{s_{k}}$ such that $\mu(p^{\prime})(a_{k+1})>0$ and $\mathrm{Pr}_{{\mathcal{M}}}(s_{k},a_{k+1},s_{k+1})>0$. Thus, a run in $\Pi({\mathcal{M}},\mu)$ can only be extended using actions that have a positive probability according to $\mu$. Let $\Omega({\mathcal{M}},\mu)$ denote the subset of $\Omega({\mathcal{M}})$ that includes exactly all runs $r$ such that for every $p\in\Pi({\mathcal{M}})$ that is a prefix of $r$, it holds that $p\in\Pi({\mathcal{M}},\mu)$. Let ${\mathcal{F}}({\mathcal{M}})$ be the set of all strategies.
We say that a strategy $\mu$ is pure if for all $p\in\Pi({\mathcal{M}})$, the distribution $\mu(p)$ is a point distribution, and is mixed otherwise. We say that $\mu$ has finite-memory, if there exists a finite set $M$ of memories, with a distinguished initial memory $m_{0}\in M$, and a memory update function $\eta:M\times S\to M$ such that $\mu$ only depends on the current memory and MDP state. Formally, $\eta$ is extended to memories and finite sequences in $S^{*}$ as follows: $\eta:M\times S^{*}\to M$ is such that for every memory $m\in M$, finite sequence $p\in S^{*}$, and MDP state $s\in S$, we have that $\eta(m,\epsilon)=m$ and $\eta(m,p\cdot s)=\eta(\eta(m,p),s)$.
Then, for all $p,p^{\prime}\in\Pi({\mathcal{M}},\mu)$ and $s\in S$ such that $p\cdot s,p^{\prime}\cdot s\in\Pi({\mathcal{M}},\mu)$, if $\eta(m_{0},p)=\eta(m_{0},p^{\prime})$ then $\mu(p\cdot s)=\mu(p^{\prime}\cdot s)$. Equivalently, $\mu$ can be thought as a function $\mu:M\times S\rightarrow\mathrm{dist}(A)\times M$, that given the current memory $m$ and the MDP state $s$, outputs a distribution on the available actions from $s$ and the next memory state to move to.
Indeed, we say that $\langle m,s\rangle\in M\times S$ are feasible by $\mu$ if there exists $p\in\Pi({\mathcal{M}},\mu)$ such that $p\cdot s\in\Pi({\mathcal{M}},\mu)$ and $\eta(m_{0},p)=m$. Then for every feasible $\langle m,s\rangle\in M\times S$, let $p_{m,s}\in\Pi({\mathcal{M}},\mu)$ be a witness for feasibility. We can now define $\mu^{\prime}:M\times S\nrightarrow\mathrm{dist}(A)\times M$ as a partial function on all feasible $\langle m,s\rangle\in M\times S$ by $\mu^{\prime}(m,s)=\langle\mu(p_{m,s}\cdot s),\eta(m,s)\rangle$. Observe that $\mu(p\cdot s)$ only depends on $s$ and the memory $m=\eta(m_{0},p)$. That is, for all $p\cdot s\in\Pi({\mathcal{M}},\mu)$ we have that $\mu(p\cdot s)=\mu(p_{m,s}\cdot s)$ for $m=\eta(m_{0},p)$. Thus, if $\mu^{\prime}(m,s)=\langle d,m^{\prime}\rangle$, then $\mu(p\cdot s)=d$ for all $p\cdot s\in\Pi({\mathcal{M}},\mu)$ such that $\eta(m_{0},p)=m$.
We say that $\mu$ is memoryless if it has a finite memory $M$ with $|M|=1$. Equivalently, if for all $p,p^{\prime}\in\Pi({\mathcal{M}},\mu)$ and $s\in S$ such that $p\cdot s,p^{\prime}\cdot s\in\Pi({\mathcal{M}},\mu)$, it holds that $\mu(p\cdot s)=\mu(p^{\prime}\cdot s)$. Namely, if $\mu$ only depends on the current state, and not on the history beforehand.
A Markov Chain (MC, for short) is an MDP with a dummy set of actions, $A_{s}=\{\ast\}$ for all $s\in S$. Formally, an MC is ${\mathcal{C}}=\langle S,s_{0},\mathrm{Pr}_{\mathcal{C}},\Sigma,\tau\rangle$, where $S$ is the set of states, $s_{0}\in S$ is the initial state, $\mathrm{Pr}_{\mathcal{C}}:S\times S\rightarrow[0,1]$ is a stochastic transition function: for every $s\in S$ it holds that $\sum_{s^{\prime}\in S}\mathrm{Pr}_{\mathcal{C}}(s,s^{\prime})=1$. A random process of ${\mathcal{C}}$, is an infinite sequence of random variables $S_{0},S_{1},S_{2},\ldots$ with values in $S$, such that $S_{0}=s_{0}$ with probability 1, and for all $s_{1},s_{2},\ldots,s_{i+1}\in S$, $i\geq 0$, it holds that $\mathbb{P}_{\mathcal{C}}(S_{i+1}=s_{i+1}|S_{0}=s_{0},\ldots,S_{i}=s_{i})=\mathrm{Pr}_{\mathcal{C}}(s_{i},s_{i+1})$.
Given a strategy $\mu$, we can obtain from ${\mathcal{M}}$ and $\mu$ the MC ${\mathcal{C}}({\mathcal{M}},\mu)=\langle\Pi({\mathcal{M}},\mu),s_{0},\mathrm{Pr}_{\mathcal{M}}^{\mu}\rangle$ in which the choice of actions is resolved according to $\mu$. Formally, if $p,p^{\prime}\in\Pi({\mathcal{M}},\mu)$ are such that there are $s,s^{\prime}\in S$, such that $p^{\prime}$ ends with $s^{\prime}$ and $p=p^{\prime}\cdot s$, then $\mathrm{Pr}_{\mathcal{M}}^{\mu}(p^{\prime},p)=\sum_{a\in A_{s^{\prime}}}\mu(p^{\prime})(a)\cdot\mathrm{Pr}_{\mathcal{M}}(s^{\prime},a,s)$. It is clear that when $\mu$ is memoryless, that is, $\mu$ only depends on the last position in $p^{\prime}$, then ${\mathrm{Pr}}_{\mathcal{M}}^{\mu}$ only depends on $s^{\prime}$ and $s$, and hence the projection of the chain ${\mathcal{C}}({\mathcal{M}},\mu)$ onto the last position in a path is a finite MC. Formally, let $\iota:\Pi({\mathcal{M}},\mu)\rightarrow S$ be the function that maps a finite path to its last state. Let $S^{\prime}=\iota(\Pi({\mathcal{M}},\mu))$ be the set of all states for which there is a positive probability to reach when playing with $\mu$, and for every $s\in S^{\prime}$ let $p_{s}\in\Pi({\mathcal{M}},\mu)$ be some path with $\iota(p_{s})=s$. Let ${\mathrm{Pr}^{\prime}}_{{\mathcal{M}}}^{\mu}:S^{\prime}\times S^{\prime}\rightarrow[0,1]$ be the stochastic transition function that is defined by ${\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}(s,s^{\prime})=\sum_{a\in A_{s}}\mu(p_{s})(a)\cdot\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})$. Then, if $\langle p_{0},p_{1},p_{2},\ldots\rangle\in(\Pi({\mathcal{M}},\mu))^{\omega}$ is an infinite sequence of increasing paths that are sampled according to ${\mathcal{C}}({\mathcal{M}},\mu)$, then the sequence $\langle\iota(p_{0}),\iota(p_{1}),\iota(p_{2}),\ldots\rangle\in S^{\omega}$ satisfies $\mathbb{P}(\iota(p_{i+1})=s^{\prime}|\iota(p_{i})=s)={\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}(s,s^{\prime})$ for all $s^{\prime},s\in S^{\prime}$.
Let ${\mathcal{M}}=\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{\mathcal{M}},\Sigma,\tau\rangle$ be an MDP and ${\mathcal{A}}=\langle\Sigma,Q,q_{0},\delta,\alpha\rangle$ be a nondeterministic automaton. When the player takes actions in ${\mathcal{M}}$ the infinite word $\tau(r)\in\Sigma^{\omega}$ is generated. Assume that, simultaneously, the player also generates a run of ${\mathcal{A}}$. This process can be modeled by an MDP that is a product of ${\mathcal{M}}$ with ${\mathcal{A}}$, and in each step the player takes an action of ${\mathcal{M}}$ and chooses a successor state in ${\mathcal{A}}$ that extends the current run on the new letter drawn by ${\mathcal{M}}$. Formally, we define the MDP ×${\mathcal{M}}\times{\mathcal{A}}=\langle S^{\prime},\langle s_{0},\bot\rangle,(A_{\langle s,q\rangle})_{\langle s,q\rangle\in S^{\prime}},\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{A}}},\Sigma,\gamma\rangle$ that is defined as follows. The set of states of ×${\mathcal{M}}\times{\mathcal{A}}$ is $S^{\prime}=(S\times Q)\cup\{\langle s_{0},\bot\rangle\}$, for $\bot\notin Q$. For all $\langle s,q\rangle\in S\times Q$, the set of actions $A_{\langle s,q\rangle}$ available from $\langle s,q\rangle$ consists of all pairs $\langle a,q^{\prime}\rangle\in A_{s}\times\delta(q,\tau(s))$. In addition, the actions available from $\langle s_{0},\bot\rangle$ are $A_{\langle s_{0},\bot\rangle}=A_{s_{0}}\times\{q_{0}\}$. Then, for all $\langle s,\ast\rangle\in S^{\prime}$ and $\langle a,q^{\prime}\rangle\in A_{\langle s,\ast\rangle}$, we have $\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{A}}}(\langle s,\ast\rangle,\langle a,q^{\prime}\rangle,\langle s^{\prime},q^{\prime}\rangle)=\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$ and $\gamma(s,\ast)=\tau(s)$. Given a finite or infinite run $r\in\Pi({\mathcal{M}}\times{\mathcal{A}})\cup\Omega({\mathcal{M}}\times{\mathcal{A}})$, $r=\langle\langle s_{0},\bot\rangle,\langle s_{1},q_{0}\rangle,\langle s_{2},q_{1}\rangle,\ldots\rangle$, we define $\pi_{\mathcal{A}}(r)=\langle q_{0},q_{1},q_{2},\ldots\rangle$, and $\pi_{\mathcal{M}}(r)=\langle s_{0},s_{1},s_{2},\ldots\rangle$ to be the ${\mathcal{A}}$-run and ${\mathcal{M}}$-run that correspond to $r$ respectively. Conversely, for $r_{\mathcal{M}}=\langle s_{0},s_{1},s_{2},\ldots\rangle\in S^{*}\cup S^{\omega}$ and $r_{\mathcal{A}}=\langle q_{0},q_{1},q_{2},\ldots\rangle\in(Q_{\mathcal{A}})^{*}\cup(Q_{\mathcal{A}})^{\omega}$ of the same length, let $r_{\mathcal{M}}\oplus r_{\mathcal{A}}=\langle\langle s_{0},\bot\rangle,\langle s_{1},q_{0}\rangle,\langle s_{2},q_{1}\rangle,\ldots\rangle$. Notice that for all $r_{\mathcal{M}}\oplus r_{\mathcal{A}}\in\Pi({\mathcal{M}}\times{\mathcal{A}})\cup\Omega({\mathcal{M}}\times{\mathcal{A}})$, we have $\pi_{\mathcal{A}}(r_{\mathcal{M}}\oplus r_{\mathcal{A}})=r_{\mathcal{A}}$ and $\pi_{\mathcal{M}}(r_{\mathcal{M}}\oplus r_{\mathcal{A}})=r_{\mathcal{M}}$.
The semantic satisfaction probability of a strategy $\mu\in{\mathcal{F}}({\mathcal{M}})$ with respect to ${\mathcal{A}}$, denoted $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)$, is defined as:
$\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)={\mathrm{Pr}}_{\mathcal{M}}^{\mu}(r\in\Omega({\mathcal{M}},\mu)|\tau(r)\in L({\mathcal{A}})).$
That is, $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)$ is the probability that when the player plays with $\mu$, the word $\tau(r)$ generated by ${\mathcal{M}}$ is in $L({\mathcal{A}})$. Then, the semantic satisfaction probability of an MDP ${\mathcal{M}}$ with respect to ${\mathcal{A}}$, denoted $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$, is defined by
$\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}=\sup_{\mu\in{\mathcal{F}}({\mathcal{M}})}\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu).$
That is, $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$ is the least upper bound on how good the player can do when playing with ${\mathcal{M}}$ and ${\mathcal{A}}$. It is well known that for every nondeterministic automaton ${\mathcal{A}}$ with an $\omega$-regular acceptance condition, the suprimum is attained by a pure finite-memory strategy[11].
Similarly, the syntactic satisfaction probability of $\mu\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})$ and ${\mathcal{M}}$, denoted $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu)$ and $\mathrm{PSyn}^{\mathcal{A}}_{{\mathcal{M}}}$ respectively, are defined by
$\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu)={\mathrm{Pr}}_{{\mathcal{M}}\times{\mathcal{A}}}^{\mu}(r\in\Omega({\mathcal{M}}\times{\mathcal{A}},\mu)|\pi_{\mathcal{A}}(r)\mbox{ is accepting}),$
and
$\mathrm{PSyn}^{\mathcal{A}}_{{\mathcal{M}}}=\sup_{\mu\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})}\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu).$
As with the semantic case, the value $\mathrm{PSyn}^{\mathcal{A}}_{{\mathcal{M}}}$ is attained. Moreover, it is guaranteed to be attained by a pure memoryless strategy [11].
In general, $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}\leq\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$, but equality is not guaranteed, as the optimal resolution of nondeterministic choices may require access to future events.
An automaton ${\mathcal{A}}$ is Good-For-MDPs (GFM, for short) if for every MDP ${\mathcal{M}}$, it holds that $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}=\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}$.
6.2. The nondeterminism hierarchy and GFMness
In this section we locate GFMness in the nondeterminism hierarchy. For the semantic hierarchy, it is proved in [13] that GFM-NBWs are as expressive as NBWs. More specifically, every NBW can be translated into an equivalent GFM-NBW. Unlike the case of co-Büchi automata, where nondeterministic and deterministic automata have the same expressive power, making the expressiveness equivalence of GFM-NCWs and NCW straightforward, NBWs are more expressive than DBWs, making the result about GFM-NBWs interesting. Another aspect that has been studied is deciding GFMness [30]. In  [30], the authors prove an exponential-time upper bound as well as a polynomial-space lower bound for the problem of deciding whether an automaton is GFM-NBW.
We study the syntactic hierarchy, we first point out that every HD automaton is GFM. Then, by relating GFMness and almost-DBPness, we show that SD-NBWs are GFM, yet SD-NCWs need not be GFM. Thus, with respect to the syntactic hierarchy, GFM-NBWs are between SD and nondeterministic Büchi automata, while GFM-NCWs are incomparable to SD-NCWs.
Theorem 14.
Every HD automaton is GFM.
Proof 6.1.
Consider an HD automaton ${\mathcal{A}}$. Let $f$ be an HD strategy for ${\mathcal{A}}$. Then, every strategy $\mu\in{\mathcal{F}}({\mathcal{M}})$ can be augmented with $f$ to obtain a strategy $\mu\times f\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})$ such that $\mathrm{PSem}^{\mathcal{A}}_{{\mathcal{M}}}(\mu)=\mathrm{PSem}^{\mathcal{A}}_{{\mathcal{M}}\times{\mathcal{A}}}(\mu\times f)=\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu\times f)$. Thus, $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}\leq\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}$, and so ${\mathcal{A}}$ is GFM.
Theorem 15.
Every GFM automaton is almost-DBP.
Proof 6.2.
Consider an $\omega$-regular automaton ${\mathcal{A}}$ over an alphabet $\Sigma$, and consider the MDP ${\mathcal{M}}=\langle\Sigma,\sigma_{0},(A_{\sigma})_{\sigma\in\Sigma},\mathrm{Pr}_{{\mathcal{M}}},\mathsf{id}\rangle$, where $\sigma_{0}\in\Sigma$ is an arbitrary initial state, $A_{\sigma}=\{\ast\}$ is a dummy set of actions for each letter $\sigma\in\Sigma$, and $\mathrm{Pr}_{{\mathcal{M}}}(\sigma,\ast,\sigma^{\prime})=|\Sigma|^{-1}$, for all $\sigma,\sigma^{\prime}\in\Sigma$. By [11], there is a pure memoryless strategy $\mu^{*}:S\times Q_{\mathcal{A}}\cup\{\langle s_{0},\bot\rangle\}\rightarrow A\times Q_{\mathcal{A}}$ that attains $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}$. Observe that since $\mathrm{Pr}_{{\mathcal{M}}}(\sigma,\ast,\cdot)=(\sigma^{\prime}\mapsto|\Sigma|^{-1})$ is the uniform distribution on $\Sigma$ for all $\sigma\in\Sigma$, then ${\mathcal{M}}$ is essentially an MC that generates a random word $w=\sigma_{1},\sigma_{2},\sigma_{3},\ldots\in\Sigma^{\omega}$, such that all its letters are independent and distributed uniformly. Thus, the membership $w\in L({\mathcal{A}})$, where $w$ is the word generated by ${\mathcal{C}}({\mathcal{M}},\mu^{\prime})$ (or by ${\mathcal{C}}({\mathcal{M}}\times{\mathcal{A}},\mu^{\prime})$) for $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}})$ (respectively $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})$), is independent of the strategy $\mu^{\prime}$. In particular $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$ equals the probability that a random word $w$ is in $L({\mathcal{A}})$. Note that it must be that $\mu^{*}(\sigma_{0},\bot)=\langle\ast,q_{0}\rangle$; i.e., there is no choice at the initial state of ×${\mathcal{M}}\times{\mathcal{A}}$. Also, as $A_{\sigma}=\{\ast\}$, for all $\sigma\in\Sigma$, we can think of $\mu^{*}$ as a function $\mu^{*}:\Sigma\times Q_{\mathcal{A}}\rightarrow Q_{\mathcal{A}}$, with $\mu^{*}(\sigma,q)\in\delta_{\mathcal{A}}(q,\sigma)$ for all $\langle\sigma,q\rangle\in\Sigma\times Q_{\mathcal{A}}$. Thus, $\mu^{*}$ induces a pruning ${\mathcal{A}}_{\mu^{*}}$ of ${\mathcal{A}}$, and $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu^{*})$ equals to the probability that a random word is accepted by ${\mathcal{A}}_{\mu^{*}}$. Now, if ${\mathcal{A}}$ is GFM, then $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu^{*})=\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}=\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$. Hence, the probability that a random word $w\in\Sigma^{\omega}$ is accepted by ${\mathcal{A}}_{\mu^{*}}$ equals the probability that $w\in L({\mathcal{A}})$. That is, ${\mathcal{A}}_{\mu^{*}}$ is a deterministic pruning of ${\mathcal{A}}$ that loses only a set of words of measure zero. Thus, ${\mathcal{A}}$ is almost-DBP, and we are done.
Recall that in Theorem 11, we argued that there is an SD-NCW that is not almost-DBP. By Theorem 15, we can conclude that the SD-NCW described there, is also not GFM. Thus, while HDness implies GFMness, for co-Büchi automata, SDness need not imply GFMness. In the rest of this section we prove that for Büchi automata, SDness does imply GFMness, thus all SD-NBWs are GFM. In particular, Theorem 15 suggests an alternative proof of Theorem 10, which states that all SD-NBWs are almost-DBP.
Proposition 16.
If ${\mathcal{A}}$ is an SD-NBW and $\mu\in{\mathcal{F}}({\mathcal{M}})$ is a finite memory strategy, then there exists $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})$ such that $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)=\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu^{\prime})$.
Proof 6.3.
Let $\mu:M\times S\rightarrow\mathrm{dist}(A)\times M$, $\mu\in{\mathcal{F}}({\mathcal{M}})$, be a finite memory strategy, and let ${\mathcal{C}}={\mathcal{C}}({\mathcal{M}},\mu)$ be the Markov chain where actions of ${\mathcal{M}}$ are resolved according to $\mu$. Notice that since $\mu$ uses $M$ as a memory structure, we may think of ${\mathcal{C}}$ as a Markov chain with positions $M\times S$, and a stochastic transition function $\mathrm{Pr}_{\mathcal{C}}:(M\times S)^{2}\rightarrow[0,1]$ that is defined $\mathrm{Pr}_{\mathcal{C}}(\langle m,s\rangle,\langle m^{\prime},s^{\prime}\rangle)=\sum_{a\in A_{s}}d(a)\cdot\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$ when $\mu(m,s)=\langle d,m^{\prime}\rangle$, and otherwise $\mathrm{Pr}_{\mathcal{C}}(\langle m,s\rangle,\langle m^{\prime},s^{\prime}\rangle)=0$. Let $r=\langle\langle m_{0},s_{0}\rangle,\langle m_{1},s_{1}\rangle,\langle m_{2},s_{2}\rangle,\ldots\rangle\in(M\times S)^{\omega}$ be a random process sampled according to ${\mathcal{C}}$. Let $r_{\mathcal{A}}=\langle q_{0},q_{1},q_{2},\ldots\rangle\in(Q_{\mathcal{A}})^{\omega}$ be a random run of ${\mathcal{A}}$ over the word $\tau(r)$ that is sampled by playing according to $\mu$ in ${\mathcal{M}}$. I.e., given $\langle m_{0},s_{0}\rangle,\langle m_{1},s_{1}\rangle,\ldots,\langle m_{k},s_{k}\rangle$, for $k\geq 1$, and $q_{0},q_{1},\ldots,q_{k-1}$, we set $q_{k}$ to be one of the $\tau(s_{k})$-successors of $q_{k-1}$ at random. Notice that the composition $r\oplus r_{\mathcal{A}}=\langle\langle m_{0},s_{0},q_{0}\rangle,\langle m_{1},s_{1},q_{1}\rangle,\ldots\rangle$ is a process of a finite Markov chain ${\mathcal{C}}_{\mathcal{A}}$ with a set of positions $M\times S\times Q_{\mathcal{A}}$. Indeed, the stochastic transition function of ${\mathcal{C}}_{\mathcal{A}}$ can be expressed as follows:
[…]
Since ${\mathcal{C}}_{\mathcal{A}}$ is a finite Markov chain, it holds that with probability 1 the random process $r\oplus r_{\mathcal{A}}$ of ${\mathcal{C}}_{\mathcal{A}}$ eventually enters an ergodic component and traverses all its states. We say that an ergodic component $C\subseteq M\times S\times Q_{\mathcal{A}}$ of ${\mathcal{C}}_{\mathcal{A}}$ is $\alpha_{\mathcal{A}}$-accepting if the projection of $C$ onto $Q_{\mathcal{A}}$ intersects $\alpha_{\mathcal{A}}$. I.e., $C$ is $\alpha_{\mathcal{A}}$-accepting if there exist $s\in S$, $m\in M$ and $q\in\alpha_{\mathcal{A}}$ such that $\langle m,s,q\rangle\in C$. Thus, if $r\oplus r_{\mathcal{A}}$ eventually enters an $\alpha_{\mathcal{A}}$-accepting ergodic component $C$, then with probability 1 the process traverses all the positions in $C$ infinitely often, and in particular with probability 1 $r_{\mathcal{A}}$ is an accepting run of ${\mathcal{A}}$ over the generated word $\tau(r)$. Thus, since with probability 1 the process of ${\mathcal{C}}_{\mathcal{A}}$ eventually enters an ergodic component, it follows that the probability that $\tau(r)\in L({\mathcal{A}})$ but $r_{\mathcal{A}}$ is rejecting equals the probability that $r\oplus r_{\mathcal{A}}$ enters an ergodic component that is not $\alpha_{\mathcal{A}}$-accepting and $\tau(r)\in L({\mathcal{A}})$. We prove that if $r\oplus r_{\mathcal{A}}$ enters an ergodic component $C$ that is not $\alpha_{\mathcal{A}}$-accepting then $\tau(r)\notin L({\mathcal{A}})$. Note that we claim that the implication $C$ not $\alpha_{\mathcal{A}}$-accepting implies that $\tau(r)\notin L({\mathcal{A}})$, is surely, and not only almost-surely. Consider an ergodic component $C$ that is not $\alpha_{\mathcal{A}}$-accepting, and assume that $r\oplus r_{\mathcal{A}}$ is such that $\langle m_{k},s_{k},q_{k}\rangle\in C$ for some $k\geq 0$. Assume by way of contradiction that $\tau(r)\in L({\mathcal{A}})$. Since ${\mathcal{A}}$ is SD, it is possible to extend the finite run $q_{0},q_{1},\ldots q_{k}$ of ${\mathcal{A}}$ over $\tau(s_{1}),\tau(s_{2}),\ldots,\tau(s_{k})$ into an accepting run of ${\mathcal{A}}$ over $\tau(r)$. In particular there exist $j>k$ and $q^{\prime}_{k+1},q^{\prime}_{k+2},\ldots,q^{\prime}_{j}\in(Q_{\mathcal{A}})^{*}$, such that $q^{\prime}_{j}\in\alpha_{\mathcal{A}}$ and $q_{0},\ldots,q_{k},q^{\prime}_{k+1},\ldots,q^{\prime}_{j}$ is a finite run of ${\mathcal{A}}$ over $\tau(s_{1}),\ldots,\tau(s_{j})$. By definition of the stochastic transition function of ${\mathcal{C}}_{\mathcal{A}}$, it follows that there is a positive probability to move to $\langle m_{k+1},s_{k+1},q^{\prime}_{k+1}\rangle$ from $\langle m_{k},s_{k},q_{k}\rangle$. Similarly, it follows by induction that for all $k+1<i\leq j$ there is a positive probability to move to $\langle m_{i},s_{i},q^{\prime}_{i}\rangle$ from $\langle m_{i-1},s_{i-1},q^{\prime}_{i-1}\rangle$. In particular $\langle m_{j},s_{j},q^{\prime}_{j}\rangle\in C$, in contradiction to the assumption that $C$ is not $\alpha_{\mathcal{A}}$-accepting. Let $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})$ be the strategy for ×${\mathcal{M}}\times{\mathcal{A}}$ that resolves the actions of ${\mathcal{M}}$ according to $\mu$ and in addition generates a random run of ${\mathcal{A}}$ over $\tau(r)$. It is not hard to see that ${\mathcal{C}}_{\mathcal{A}}$ induces the same distribution over infinite paths in $(S\times Q_{\mathcal{A}})^{\omega}$ as the Markov chain ${\mathcal{C}}({\mathcal{M}}\times{\mathcal{A}},\mu^{\prime})$. We proved that there is a zero probability that $\tau(r)\in L({\mathcal{A}})$ and $r_{\mathcal{A}}$ is rejecting. Thus, $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu^{\prime})=\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)$.
We can now conclude with the desired connection between SDness and GFMness.
We first examine whether SDness implies GFMness, and show that the answer depends on the acceptance condition.
Theorem 17.
All SD-NBWs are GFM, yet there is an SD-NCW that is not GFM.
Proof 6.4.
The co-Büchi case follows from Theorems 11 and 15. For Büchi automata, consider an SD-NBW ${\mathcal{A}}=\langle\Sigma,Q_{\mathcal{A}},q_{0},\delta_{\mathcal{A}},\alpha_{\mathcal{A}}\rangle$, and a $\Sigma$-labeled MDP ${\mathcal{M}}$. Let $\mu$ be a finite memory strategy such that $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)=\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$ given by [11]. Then by Proposition 16 it holds that there exists $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})$ such that $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu^{\prime})=\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)$, and we get $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}=\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu^{\prime})\leq\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}$. Since $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}\leq\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$ holds for all ${\mathcal{A}}$ and ${\mathcal{M}}$, we have the equality $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}=\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$.
We continue and check whether GFMness implies SDness, and show that the answer is negative even for weak automata.
Theorem 18.
There is a GFM-NWW that is not SD.
Proof 6.5.
[…]
Figure 6. A GFM-NWW that is not SD. Missing transitions lead to a rejecting sink.
Appendix A Determinization of SD-NBWs
In this section we prove formally that the DBW ${\mathcal{A}}^{\prime}$ is equivalent to the SD-NBW ${\mathcal{A}}$, and that if ${\mathcal{A}}$ is weak, then so is ${\mathcal{A}}^{\prime}$. The following two propositions follow immediately from the definitions:
Proposition 19.
Consider states $q\in Q$ and $S\in Q^{\prime}$, a letter $\sigma\in\Sigma$, and transitions $\langle q,\sigma,q^{\prime}\rangle$ and $\langle S,\sigma,S^{\prime}\rangle$ of ${\mathcal{A}}$ and ${\mathcal{A}}^{\prime}$, respectively. If $q$ is ${\mathcal{A}}$-equivalent to the states in $S$, then $q^{\prime}$ is ${\mathcal{A}}$-equivalent to the states in $S^{\prime}$.
Proposition 20.
Consider a state $S$ of ${\mathcal{A}}^{\prime}$ and a letter $\sigma\in\Sigma$. If $\langle S,\sigma,S^{\prime}\rangle\in\Delta^{\prime}$ and $S^{\prime}\notin\alpha^{\prime}$, then all the $\sigma$-successors of a state $s\in S$ are in $S^{\prime}\setminus\alpha$.
We can now prove the correctness of the construction:
Proposition 21.
The automata ${\mathcal{A}}$ and ${\mathcal{A}}^{\prime}$ are equivalent.
Proof A.1.
We first prove that $L({\mathcal{A}}^{\prime})\subseteq L({\mathcal{A}})$. Let $r_{{\mathcal{A}}^{\prime}}=S_{0},S_{1},S_{2},\ldots$ be an accepting run of ${\mathcal{A}}^{\prime}$ on a word $w=\sigma_{1}\cdot\sigma_{2}\cdots$. We construct an accepting run of ${\mathcal{A}}$ on $w$. Since $r_{{\mathcal{A}}^{\prime}}$ is accepting, there are infinitely many positions $j_{1},j_{2},\ldots$ with $S_{j_{i}}\in\alpha^{\prime}$.
We also define $j_{0}=0$.
Consider the DAG $G=\langle V,E\rangle$, where
  • $V\subseteq Q\times\mathbb{N}$ is the union $\bigcup_{i\geq 0}(S_{j_{i}}\times\{i\})$.
  • $E\subseteq\bigcup_{i\geq 0}(S_{j_{i}}\times\{i\})\times(S_{j_{i+1}}\times\{i+1\})$ is such that for all $i\geq 0$, it holds that $E(\langle s^{\prime},i\rangle,\langle s,i+1\rangle)$ iff there is a finite run from $s^{\prime}$ to $s$ over $w[j_{i}+1,j_{i+1}]$. Then, we label this edge by the run from $s^{\prime}$ to $s$.
By the definition of ${\mathcal{A}}^{\prime}$, for every $j\geq 0$ and state $s_{j+1}\in S_{j+1}$, there is a state $s_{j}\in S_{j}$ such that $\langle s_{j},\sigma_{j},s_{j+1}\rangle\in\Delta$. Thus, it follows by induction that for every $i\geq 0$ and state $s_{i+1}\in S_{j_{i+1}}$, there is a state $s_{i}\in S_{j_{i}}$ such that there is a finite run from $s_{i}$ to $s_{i+1}$ on $w[j_{i}+1,j_{i+1}]$. Thus, the DAG $G$ has infinitely many reachable vertices from the vertex $\langle q_{0},0\rangle$. Also, as the nondeterminism degree of ${\mathcal{A}}$ is finite, so is the branching degree of $G$. Thus, by König’s Lemma, $G$ includes an infinite path, and the labels along the edges of this path define a run of ${\mathcal{A}}$ on $w$. Since for all $i\geq 1$, the state $S_{j_{i}}$ is in $\alpha^{\prime}$, and so all the states in $S_{j_{i}}$ are in $\alpha$, this run is accepting, and we are done.
For the other direction, assume that $w=\sigma_{1}\cdot\sigma_{2}\cdots\in L({\mathcal{A}})$, and let $r=r_{0},r_{1},\ldots$ be an accepting run of ${\mathcal{A}}$ on $w$. Let $S_{0},S_{1},S_{2}\ldots$ be the run of ${\mathcal{A}}^{\prime}$ on $w$, and assume, by way of contradiction, that there is a position $j\geq 0$ such that $S_{j},S_{j+1},\ldots$ is an $\alpha$-free run on the suffix $w[j+1,\infty]$. Then, an iterative application of Proposition 20 implies that all the runs of a state $s_{j}\in S_{j}$ on $w[j+1,\infty]$ are $\alpha$-free in ${\mathcal{A}}$. Also, an iterative application of Proposition 19 implies that $r_{j}\sim_{\mathcal{A}}s_{j}$, and since $r$ is an accepting run of ${\mathcal{A}}$, it holds that ${\mathcal{A}}^{s_{j}}$ has an accepting run on $w[j+1,\infty]$, and we have reached a contradiction.
It is left to prove that weakness of ${\mathcal{A}}$ is preserved in ${\mathcal{A}}^{\prime}$.
Proposition 22.
If ${\mathcal{A}}$ is an NWW, then ${\mathcal{A}}^{\prime}$ is a DWW.
Proof A.2.
Assume by way of contradiction that there are reachable states $S\in\alpha^{\prime}$ and $S^{\prime}\notin\alpha^{\prime}$, and an infinite run $r_{{\mathcal{A}}^{\prime}}=S_{0},S_{1},S_{2},\ldots$ that visits both $S$ and $S^{\prime}$ infinitely often. Recall that a reachable state in $Q^{\prime}$ contains only $\alpha$-states of ${\mathcal{A}}$ or only $\bar{\alpha}$-states of ${\mathcal{A}}$. Hence, $S^{\prime}$ contains only $\bar{\alpha}$-states of ${\mathcal{A}}$. As in the proof of Proposition 21, the run $r_{{\mathcal{A}}^{\prime}}$ induces an infinite run $r_{\mathcal{A}}=s_{0},s_{1},s_{2},\ldots$, where for all positions $j\geq 0$, it holds that $s_{j}\in S_{j}$. Since the run $r_{{\mathcal{A}}^{\prime}}$ visits $S$ infinitely often, then $r_{\mathcal{A}}$ visits infinitely many $\alpha$-states. Likewise, since $r_{{\mathcal{A}}^{\prime}}$ visits $S^{\prime}$ infinitely often, then $r_{\mathcal{A}}$ also visits infinitely many $\bar{\alpha}$-states. This contradicts the weakness of ${\mathcal{A}}$, and we are done.
Appendix B Details of the Reduction in Theorem 3
We describe the technical details of the construction of ${\mathcal{A}}$. Let $T=\langle\Gamma,Q,\rightarrow,q_{0},q_{acc},q_{rej}\rangle$, where $\Gamma$ is the working alphabet, $Q$ is the set of states, $\rightarrow\subseteq Q\times\Gamma\times Q\times\Gamma\times\{L,R\}$ is the transition relation (we use $(q,a)\rightarrow(q^{\prime},b,\Delta)$ to indicate that when $T$ is in state $q$ and it reads the input $a$ in the current tape cell, it moves to state $q^{\prime}$, writes $b$ in the current tape cell, and its reading head moves one cell to the left/right, according to $\Delta$), $q_{0}$ is the initial state, $q_{acc}$ is the accepting state, and $q_{rej}$ is the rejecting one.
The transitions function $\rightarrow$ is defined also for the final states $q_{acc}$ and $q_{rej}$: when a computation of $T$ reaches them, it erases the tape, goes to the leftmost cell in the tape, and moves to the initial state $q_{0}$. Recall that $s:\mathbb{N}\rightarrow\mathbb{N}$ is the polynomial space function of $T$. Thus, when $T$ runs on the empty tape, it uses at most $n_{0}=s(0)$ cells.
We use strings of the form #$\#\gamma_{1}\gamma_{2}\ldots(q,\gamma_{i})\ldots\gamma_{n_{0}}$ to encode a configuration of $T$ on a word of length at most $n_{0}$. That is, a configuration starts with ##\##, and all its other letters are in $\Gamma$, except for one letter in $Q\times\Gamma$. The meaning of such a configuration is that the $j$’th cell in $T$, for $1\leq j\leq n_{0}$, is labeled $\gamma_{j}$, the reading head points at cell $i$, and $T$ is in state $q$. For example, the initial configuration of $T$ is #$\#(q_{0},b)b\ldots b$ (with $n_{0}-1$ occurrences of $b$’s) where $b$ stands for an empty cell. We can now encode a computation of $T$ by a sequence of configurations.
Let $\Sigma=\{\#\}\cup\Gamma\cup(Q\times\Gamma)$ and let #$\#\sigma_{1}\ldots\sigma_{n_{0}}\#\sigma^{\prime}_{1}\ldots\sigma^{\prime}_{n_{0}}$ be two successive configurations of $T$. We also set $\sigma_{0}$, $\sigma^{\prime}_{0}$, and $\sigma_{n_{0}+1}$ to ##\##. For each triple $\langle\sigma_{i-1},\sigma_{i},\sigma_{i+1}\rangle$ with $1\leq i\leq n_{0}$, we know, by the transition relation of $T$, what $\sigma^{\prime}_{i}$ should be. In addition, the letter ##\## should repeat exactly every $n_{0}+1$ letters. Let $next(\sigma_{i-1},\sigma_{i},\sigma_{i+1})$ denote our expectation for $\sigma^{\prime}_{i}$. That is,
  • $next(\sigma_{i-1},\sigma_{i},\sigma_{i+1})=\sigma_{i}$ if $\sigma_{i}=\#$, or if non of $\sigma_{i-1},\sigma_{i}$ and $\sigma_{i+1}$ are in $Q\times\Gamma$.
  • $next(\sigma_{i-1},\sigma_{i},\sigma_{i+1})=\#$ if |$|\{\sigma_{i-1},\sigma_{i},\sigma_{i+1}\}\cap Q\times\Gamma|\geq 2$, or if |$|\{\sigma_{i-1},\sigma_{i},\sigma_{i+1}\}\cap\{\#\}|\geq 2$  [^10] These case describes an illegal encoding of a configuration, and there is no correct expectation of the letter that comes in the same position in the next configuration. Thus, the choice $next(\sigma_{i-1},\sigma_{i},\sigma_{i+1})=\#$ is arbitrary..
  • $next((q,\gamma_{i-1}),\gamma_{i},\gamma_{i+1})=next((q,\gamma_{i-1}),\gamma_{i},\#)=$
    $\left\{\begin{array}[]{ll}\gamma_{i}&\mbox{If $(q,\gamma_{i-1})\rightarrow(q^{\prime},\gamma^{\prime}_{i-1},L)$}\\ (q^{\prime},\gamma_{i})&\mbox{If $(q,\gamma_{i-1})\rightarrow(q^{\prime},\gamma^{\prime}_{i-1},R)$}\end{array}\right.$
  • $next(\gamma_{i-1},\gamma_{i},(q,\gamma_{i+1}))=next(\#,\gamma_{i},(q,\gamma_{i+1}))=$
    $\left\{\begin{array}[]{ll}\gamma_{i}&\mbox{If $(q,\gamma_{i+1})\rightarrow(q^{\prime},\gamma^{\prime}_{i+1},R)$}\\ (q^{\prime},\gamma_{i})&\mbox{If $(q,\gamma_{i+1})\rightarrow(q^{\prime},\gamma^{\prime}_{i},L)$}\end{array}\right.$
  • $next(\#,(q,\gamma_{i}),\gamma_{i+1})=(q^{\prime},\gamma^{\prime})$ where $(q,\gamma_{i})\rightarrow(q^{\prime},\gamma^{\prime}_{i},L)$ 111111 We assume that the reading head of $T$ stays in place when it is above the leftmost cell and the transition function directs it to move left..
  • $next(\gamma_{i-1},(q,\gamma_{i}),\gamma_{i+1})=next(\gamma_{i-1},(q,\gamma_{i}),\#)=\gamma^{\prime}_{i}$ where $(q,\gamma_{i})\rightarrow(q^{\prime},\gamma^{\prime}_{i},\Delta)$.
Consistency with $next$ now gives us a necessary condition for a word to encode a legal computation that uses $n_{0}$ tape cells.
In order to accept words that satisfy C1, namely detect a violation of the encoding, we distinguish between two types of violations: a violation of a single encoded configuration, and a violation of $next$. In order to detect a violation of $next$, the NWW ${\mathcal{A}}$ use its nondeterminism and guesses a triple $\langle\sigma_{i-1},\sigma_{i},\sigma_{i+1}\rangle\in\Sigma^{3}$ and guesses a position in the word, where it checks whether the three letters to be read starting this position are $\sigma_{i-1},\sigma_{i}$, and $\sigma_{i+1}$, and checks whether $next(\sigma_{i-1},\sigma_{i},\sigma_{i+1})$ is not the letter to come $n_{0}+1$ letters later. Once ${\mathcal{A}}$ sees such a violation, it goes to an accepting sink. If $next$ is respected, or if the guessed triple and position is not successful, then ${\mathcal{A}}$ returns to its initial state. Also, at any point that ${\mathcal{A}}$ still waits to guess a position of a triple, it can guess to return back to the initial state.
In order to detect a violation of a single encoded configuration, the NWW ${\mathcal{A}}$ accepts words that include in them a subword $x$ of length $n_{0}+1$, which is either ##\##-free, or is of the form #$\#\cdot y$, for a finite word $y$ that does not encode a legal configuration that uses $n_{0}$ cells. Specifically, $y$ is not of the form $\Gamma^{i}\cdot(Q\times\Gamma)\cdot\Gamma^{n_{0}-(i+1)}$, for some $10\leq i\leq n_{0}-$ 1. Indeed, such words either include a long subword with no ##\##, or have some block of size $n_{0}$ that follows a ##\##, that does not encode a legal configuration of $T$ when it uses $n_{0}$ cells. If ${\mathcal{A}}$ succeeds to detect such a violation, it goes to an accepting sink, otherwise, ${\mathcal{A}}$ returns to its initial state.
In order to accept words that satisfy C2, namely detect an encoding of an accepting configuration of $T$, the NWW ${\mathcal{A}}$ guesses a position where it checks if the next $n_{0}+1$ letters are of the form #$\#\cdot x\cdot(q_{acc},\gamma)\cdot y$, for $x,y\in\Gamma^{*}$ and $\gamma\in\Gamma$. If the guess succeeded, then ${\mathcal{A}}$ goes to the accepting sink, and otherwise it goes back to the initial state. At any point that ${\mathcal{A}}$ waits to guess a position where the accepting configuration begins, it can guess to return back to the initial state.
Appendix C Correctness and full details of the reduction in Theorem 4
We first prove that the HD strategy $g$ defined in Proposition 6 satisfies two essential properties. Then, in Lemma 25, we show that these properties imply that $g$ is a winning HD strategy for ${\mathcal{A}}_{\varphi}$.
Lemma 23.
For all $u\in(X\cdot C)^{*}$ and $v\in R_{n,m}$, if $g(u)=p$, then there is a prefix $y\in(X\cdot C)^{*}$ of $v$ such that $g(u\cdot y)=q_{0}$.
Proof C.1.
Let $j\in[m]$ and $2\leq k\leq n$ be such that $v$ ends with the word $x_{k}\cdot c_{j}\cdot x_{k+1}\cdot c_{j}\cdots x_{n}\cdot c_{j}$, and $k$ is the minimal index that is greater than $1$ for which $x_{k}$ participates in $c_{j}$. Since we assume that each of the clauses of $\varphi$ depends on at least two variables, such $k>1$ exists. Let $i\in{\{0,1\}}$ be minimal with $c_{j}\in C_{k}^{i}$, and let $z\in(X\cdot C)^{*}$ be a prefix of $v$ such that $v=z\cdot x_{k}\cdot c_{j}\cdots x_{n}\cdot c_{j}$. If there is a prefix $y\in(X\cdot C)^{*}$ of $z$ such that $g(u\cdot y)=q_{0}$, then we are done. Otherwise, $g(u\cdot z)=p$. By definition of $g$ and the choice of $k$, we know that $g(u\cdot z\cdot x_{k})=q^{i}_{k}$, where the assignment $x_{k}=i$ satisfies $c_{j}$. Thus, if we take $y=z\cdot x_{k}\cdot c_{j}$, then $g(u\cdot y)=q_{0}$, and $y$ is a prefix of $v$.
Lemma 24.
For all $u\in(X\cdot C)^{*}$ and $v\in R_{n,m}$, if $g(u)=q_{0}$, there is a prefix $z\in(X\cdot C)^{*}$ of $v$ such that $g(u\cdot z)=p$.
Proof C.2.
This follows immediately from the fact that $D_{n,m}$ is a DFW that recognizes $R_{n,m}$ and $p$ is the only accepting state of $D_{n,m}$. Thus, we may take $z$ to be the minimal prefix of $v$ that is in $R_{n,m}$.
Recall that an HD strategy $g:\Sigma^{*}\rightarrow Q$ has to agree with the the transitions of ${\mathcal{A}}_{\varphi}$. That is, for all $w\in(X\cdot C)^{*}$, $x_{k}\in X$, and $c_{j}\in C$, it holds that $(g(w),x_{k},g(w\cdot x_{k}))$ and $(g(w\cdot x_{k}),c_{j},g(x\cdot x_{k}\cdot c_{j}))$ are in $\Delta_{\varphi}$. In addition, if $g$ satisfies the conditions in Lemmas 23 and 24, we say that $g$ supports a $(p,q_{0})$-circle.
Lemma 25.
If $g:\Sigma^{*}\rightarrow Q$ is consistent with $\Delta_{\varphi}$ and supports a $(p,q_{0})$-circle, then for all words $w\in L_{n,m}$, the run $g(w)$ is accepting.
Proof C.3.
Consider a word $w\in L_{n,m}=(R_{n,m})^{\omega}$. Observe that if $w^{\prime}\in(X\cdot C)^{\omega}$ is a suffix of $w$, then $w^{\prime}\in L_{n,m}$, and hence has a prefix in $R_{n,m}$. Thus, if $g$ supports a $(p,q_{0})$-circle, there exist $y_{1},z_{1}\in(X\cdot C)^{*}$, such that $y_{1}\cdot z_{1}$ is a prefix of $w$, $g(y_{1})=q_{0}$, and $g(y_{1}\cdot z_{1})=p$. Let $w^{\prime}\in(X\cdot C)^{\omega}$ be the suffix of $w$ with $w=y_{1}\cdot z_{1}\cdot w^{\prime}$. By the above, $w^{\prime}\in L_{n,m}$, and we can now apply again the assumption on $g$ to obtain $y_{2},z_{2}\in(X\cdot C)^{*}$ such that $y_{2}\cdot z_{2}$ is a prefix of $w^{\prime}$, $g(y_{1}\cdot z_{1}\cdot y_{2})=q_{0}$, and $g(y_{1}\cdot z_{1}\cdot y_{2}\cdot z_{2})=p$. By iteratively applying this argument, we construct $\{y_{i},z_{i}:i\geq 1\}\subseteq(X\cdot C)^{*}$, such that $w^{i}=y_{1}\cdot z_{1}\cdot y_{2}\cdot z_{2}\cdots y_{i-1}\cdot z_{i-1}\cdot y_{i}$ is a prefix of $w$, and $g(w^{i})=q_{0}$, for all $i\geq 1$. We conclude that $q_{0}\in\inf(g(w))$, and hence $g(w)$ is accepting.
It is easy to see that there is a correspondence between assignments to the variables in $X$ and deterministic prunnings of ${\mathcal{A}}_{\varphi}$. Indeed, a pruning of $p$ amounts to choosing, for each $k\in[n]$, a value $i_{k}\in{\{0,1\}}$: the assignment $x_{k}=i_{k}$ corresponds to keeping the transition $\langle p,x_{k},q^{i_{k}}_{k}\rangle$ and removing the transition $\langle p,x_{k},q^{\lnot i_{k}}_{k}\rangle$. For an assignment $a:X\rightarrow{\{0,1\}}$, we denote by ${\mathcal{A}}_{\varphi}^{a}$ the deterministic pruning of ${\mathcal{A}}_{\varphi}$ that is associated with $a$. We prove that $a$ satisfies $\varphi$ iff ${\mathcal{A}}_{\varphi}^{a}$ is equivalent to ${\mathcal{A}}_{\varphi}$. Thus, the number of deterministic prunnings of ${\mathcal{A}}_{\varphi}$ that result in a DBW equivalent to ${\mathcal{A}}_{\varphi}$, equals to the number of assignments that satisfy $\varphi$. In particular, $\varphi$ is satisfiable iff ${\mathcal{A}}_{\varphi}$ is DBP.
Proposition 26.
For every assignment $a:X\rightarrow{\{0,1\}}$, we have that $L({\mathcal{A}}^{a}_{\varphi})=L({\mathcal{A}}_{\varphi})$ iff $\varphi$ is satisfied by $a$.
Proof C.4.
Assume first that $\varphi$ is not satisfied by $a$. We prove that $L_{n,m}\neq L({\mathcal{A}}^{a}_{\varphi})$. Let $j\in[m]$ be such that $c_{j}$ is not satisfied by $a$. I.e, for all $k\in[n]$ the assignment $x_{k}=i_{k}$ does not satisfy $c_{j}$. Since $q^{i}_{k}$ is reachable in ${\mathcal{A}}_{\varphi}^{a}$ iff $i=i_{k}$, and all $c_{j}$-labeled transitions from $\{q^{i_{k}}_{k}:k\in[n]\}$ are to $p$, it follows that the run of ${\mathcal{A}}_{\varphi}^{a}$ on $\{x_{1}\cdot c_{j}\cdot x_{2}\cdot c_{j}\cdots x_{n}\cdot c_{j}\}^{\omega}$ never visits $q_{0}$, and hence is rejecting. Thus, $(x_{1}\cdot c_{j}\cdot x_{2}\cdot c_{j}\cdots x_{n}\cdot c_{j})^{\omega}\in L_{n,m}\setminus L({\mathcal{A}}_{\varphi}^{a})$. For the other direction, we assume that $a$ satisfies $\varphi$ and prove that $L({\mathcal{A}}_{\varphi}^{a})=L_{n,m}$. Let $g^{a}:\Sigma^{*}\rightarrow Q$ be the memoryless strategy that correspond to the pruning ${\mathcal{A}}^{a}_{\varphi}$. By Lemma 25, it is sufficient proving that $g^{a}$ supports a $(p,q_{0})$-circle. Note that every strategy for $L_{n,m}$ satisfies Lemma 24. Indeed, the proof only uses the fact that $D_{n,m}$ is a DFW that recognizes $R_{n,m}$ with a single accepting state $p$. Thus, we only need to prove that $g^{a}$ satisfies Lemma 23. That is, for all $u\in(X\cdot C)^{*}$ and $v\in R_{n,m}$, if $g^{a}(u)=p$, then there is a prefix $y\in(X\cdot C)^{*}$ of $v$, such that $g^{a}(u\cdot y)=q_{0}$. Consider such words $u$ and $v$, and let $j\in[m]$ be such that $c_{j}$ is the last letter of $v$. Let $k\in[n]$ be the minimal index for which $c_{j}\in C^{i_{k}}_{k}$, and let $z\in(X\cdot C)^{*}$ be a prefix of $v$ such that $v=z\cdot x_{k}\cdot c_{j}\cdot x_{k+1}\cdot c_{j}\cdots x_{n}\cdot c_{j}$. If there exists a prefix $y\in(X\cdot C)^{*}$ of $z$ such that $g^{a}(u\cdot y)=q_{0}$, then we are done. Otherwise, the finite run of ${\mathcal{A}}_{\varphi}^{a}$ on $z$ from $p$, returns back to $p$, and hence $g^{a}(u\cdot z)=p$. Now $g^{a}(u\cdot z\cdot x_{k})=q^{i_{k}}_{k}$, and since $x_{k}=i_{k}$ satisfies $c_{j}$ we have $g^{a}(u\cdot z\cdot x_{k}\cdot c_{j})=q_{0}$. Thus, we may take $y=z\cdot x_{k}\cdot c_{j}$ which is a prefix of $v$, and we are done.

### v4

@@
 It allows a computing machine to examine several possible actions simultaneously.
-For automata on finite words, nondeterminism does not increase the expressive power, yet it leads to an exponential succinctness [28].
+For automata on finite words, nondeterminism does not increase the expressive power, yet it leads to an exponential succinctness [RS59].
 […]
 For automata on infinite words, nondeterminism may increase the expressive power and also leads to an exponential succinctness.
-For example, NBWs are strictly more expressive than DBWs [22], whereas NCWs are as expressive as DCWs [24].
+For example, NBWs are strictly more expressive than DBWs [Lan69], whereas NCWs are as expressive as DCWs [MH84].
 In some applications of the automata-theoretic approach, such as model checking, algorithms can be based on nondeterministic automata, whereas in other applications, such as synthesis and control, they cannot.
-There, the advantages of nondeterminism are lost, and algorithms involve a complicated determinization construction [29] or acrobatics for circumventing determinization [21].
+There, the advantages of nondeterminism are lost, and algorithms involve a complicated determinization construction [Saf88] or acrobatics for circumventing determinization [KV05b].
 Essentially, the inherent difficulty of using nondeterminism in synthesis and control lies in the fact that each guess of the nondeterministic automaton should accommodate all possible futures.
-A study of nondeterministic automata that can resolve their nondeterministic choices in a way that only depends on the past started in [19], where the setting is modeled by means of tree automata for derived languages.
-It then continued by means of history deterministic (HD) automata [15].
-[^1] The notion used in [15] is good for games (GFG) automata, as they address the difficulty of playing games on top of a nondeterministic automaton.
+A study of nondeterministic automata that can resolve their nondeterministic choices in a way that only depends on the past started in [KSV06], where the setting is modeled by means of tree automata for derived languages.
+It then continued by means of history deterministic (HD) automata [HP06].
+[^1] The notion used in [HP06] is good for games (GFG) automata, as they address the difficulty of playing games on top of a nondeterministic automaton.
 As it turns out, the property of being good for games varies in different settings and HD is good for applications beyond games.
-Therefore, we use the term history determinism, introduced by Colcombet in the setting of quantitative automata with cost functions [12].
+Therefore, we use the term history determinism, introduced by Colcombet in the setting of quantitative automata with cost functions [Col09].
 A nondeterministic automaton ${\mathcal{A}}$ over an alphabet $\Sigma$ is HD if there is a strategy $g$ that maps each finite word $u\in\Sigma^{*}$ to the transition to be taken after $u$ is read; and following $g$ results in accepting all the words in the language of ${\mathcal{A}}$.
@@
 Obviously, there exist HD automata: deterministic ones, or nondeterministic ones that are determinizable by pruning (DBP); that is, ones that embody an equivalent deterministic automaton.
-In fact, the HD automata constructed in [15] are DBP.
-[^2] As explained in [15], the fact that the HD automata constructed there are DBP does not contradict their usefulness in practice, as their transition relation is simpler than the one of the embodied deterministic automaton and it can be defined symbolically.
-Beyond the theoretical interest in DBP automata, they are used for modelling online algorithms: by relating the “unbounded look ahead” of optimal offline algorithms with nondeterminism, and relating the “no look ahead” of online algorithms with determinism, it is possible to reduce questions about the competitive ratio of online algorithms and the memory they require to questions about DBPness [2, 3].
-For automata on finite words, HD-NFWs are always DBP [19, 25].
-For automata on infinite words, HD-NBWs and HD-NCWs are as expressive as DBWs and DCWs, respectively [19, 27], but they need not be DBP [5].
-Moreover, the best known determinization construction for HD-NBWs is quadratic, and determinization of HD-NCWs has a tight exponential blow-up [17].
+In fact, the HD automata constructed in [HP06] are DBP.
+[^2] As explained in [HP06], the fact that the HD automata constructed there are DBP does not contradict their usefulness in practice, as their transition relation is simpler than the one of the embodied deterministic automaton and it can be defined symbolically.
+Beyond the theoretical interest in DBP automata, they are used for modelling online algorithms: by relating the “unbounded look ahead” of optimal offline algorithms with nondeterminism, and relating the “no look ahead” of online algorithms with determinism, it is possible to reduce questions about the competitive ratio of online algorithms and the memory they require to questions about DBPness [AKL10, AK15].
+For automata on finite words, HD-NFWs are always DBP [KSV06, Mor03].
+For automata on infinite words, HD-NBWs and HD-NCWs are as expressive as DBWs and DCWs, respectively [KSV06, NW98], but they need not be DBP [BKKS13].
+Moreover, the best known determinization construction for HD-NBWs is quadratic, and determinization of HD-NCWs has a tight exponential blow-up [KS15].
 Thus, HD automata on infinite words are (possibly even exponentially) more succinct than deterministic ones.
-Further research studies characterization, typeness, complementation, and further constructions and decision procedures for HD automata [17, 7, 4], as well as an extension of the HD setting to pushdown $\omega$-automata [23] and to alternating automata [8, 6].
+Further research studies characterization, typeness, complementation, and further constructions and decision procedures for HD automata [KS15, BKS17, BK18], as well as an extension of the HD setting to pushdown $\omega$-automata [LZ20] and to alternating automata [BL19, BKLS20].
 A nondeterministic automaton is semantically deterministic (SD, for short) if its nondeterministic choices lead to states with the same language.
@@
 Indeed, if different $\sigma$-successors of a state in an HD automaton have different languages, then we can prune the transitions to states whose language is contained in the language of another $\sigma$-successor.
-Moreover, since containment for HD automata can be checked in polynomial time, such a pruning can be done in polynomial time [14, 17].
-Accordingly (see more in Section 2.3), we can assume that all HD automata are SD [17].
+Moreover, since containment for HD automata can be checked in polynomial time, such a pruning can be done in polynomial time [HKR02, KS15].
+Accordingly (see more in Section 2.3), we can assume that all HD automata are SD [KS15].
 Hence, we obtain the following hierarchy, from deterministic to nondeterministic automata, where each level is a special case of the levels to its right.
@@
 In fact, the three internal levels coincide already in the syntactic sense: every SD-NFW is DBP.
-Also, given an NFW, deciding whether it is SD, HD or DBP, can each be done in polynomial time [2].
+Also, given an NFW, deciding whether it is SD, HD or DBP, can each be done in polynomial time [AKL10].
 For Büchi and co-Büchi automata, the picture is less clear, and is the subject of our research.
@@
 An unambiguous NFW for a non-empty language is SD iff it is deterministic, and a DBP-NFW need not be unambiguous.
-It is known, however, that an HD unambiguous NCW, or NBW, is DBP [7].
+It is known, however, that an HD unambiguous NCW, or NBW, is DBP [BKS17].
 We study the following aspects and questions about the hierarchy.
@@
 For each nondeterminism level we study for which acceptance condition it is a strict super class of its preceding level.
-Recall that not all HD-NBWs and HD-NCWs are DBP [5], and examples for this include also SD automata.
-On the other hand, all HD-NWWs (in fact, all HD-NXWs whose language can be recognized by a DWW) are DBP [7].
+Recall that not all HD-NBWs and HD-NCWs are DBP [BKKS13], and examples for this include also SD automata.
+On the other hand, all HD-NWWs (in fact, all HD-NXWs whose language can be recognized by a DWW) are DBP [BKS17].
 We show that SD-NXWs need not be HD for all ${\rm X}\in\{{\rm B,C,W}\}$.
@@
 Deciding the nondeterminism level of an automaton
-It is already known that deciding the HDness of a given NXW, for ${\rm X}\in\{{\rm B,C,W}\}$, can be done in polynomial time [2, 17, 4].
-On the other hand, deciding whether a given NCW is DBP is NP-complete [16].
+It is already known that deciding the HDness of a given NXW, for ${\rm X}\in\{{\rm B,C,W}\}$, can be done in polynomial time [AKL10, KS15, BK18].
+On the other hand, deciding whether a given NCW is DBP is NP-complete [KM18].
 We complete the picture in three directions.
@@
 Recall that DBPness is associated with the ability of an online algorithm to perform as good as an offline one.
-A typical analysis of the performance of an on-line algorithm compares its performance with that of an off-line algorithm.
-The notion of almost-DBPness captures cases where the on-line algorithm performs, with probability $1$, as good as the offline algorithm.
-From the point of view of formal-language theory, almost-DBPness captures the ability to approximate from below the language of automata that need not be DBP.
+A typical analysis of the performance of an online algorithm compares its performance with that of an offline algorithm.
+The notion of almost-DBPness captures cases where the online algorithm performs, with probability $1$, as good as the offline algorithm.
+From the point of view of formal language theory, almost-DBPness captures the ability to approximate from below the language of automata that need not be DBP.
 We study the almost-DBPness of HD and SD automata.
@@
 A nondeterministic automaton is good-for-MDPs (GFM) if its product with Markov decision processes maintains the probability of acceptance.
-Thus, GFM automata can replace deterministic automata when reasoning about probabilistic systems [13, 30].
+Thus, GFM automata can replace deterministic automata when reasoning about probabilistic systems [HPS+20, STZ22].
 We study the relation between semantic determinism and GFMness.
@@
 The $\alpha$-free SCCs of ${\mathcal{A}}$ are the SCCs of ${\mathcal{A}}$ that do not contain states from $\alpha$.
-A Büchi automaton ${\mathcal{A}}$ is weak [26] if for each SCC $C$ in $G_{\mathcal{A}}$, either $C\subseteq\alpha$ (in which case we say that $C$ is an accepting SCC) or $C\cap\alpha=\emptyset$ (in which case we say that $C$ is a rejecting SCC).
+A Büchi automaton ${\mathcal{A}}$ is weak [MSS88] if for each SCC $C$ in $G_{\mathcal{A}}$, either $C\subseteq\alpha$ (in which case we say that $C$ is an accepting SCC) or $C\cap\alpha=\emptyset$ (in which case we say that $C$ is a rejecting SCC).
 Note that a weak automaton can be viewed as both a Büchi and a co-Büchi automaton, as a run of ${\mathcal{A}}$ visits $\alpha$ infinitely often, iff it gets trapped in an accepting SCC, iff it visits states in $Q\setminus\alpha$ only finitely often.
@@
 for all positions $i\geq 1$ and letters $\sigma\in\Sigma$, the probability that $\sigma_{i}$ is $\sigma$ is $\frac{1}{|\Sigma|}$.
-Let ${\mathcal{A}}=\langle\Sigma,Q,q_{0},\delta,\alpha\rangle$ be a determnistic automaton, and let $G_{\mathcal{A}}=\langle Q,E\rangle$ be its induced directed graph.
-A random walk on ${\mathcal{A}}$, is a random walk on $G_{\mathcal{A}}$ with the probability matrix $P(q,p)=\frac{|\{\sigma\in\Sigma:\langle q,\sigma,p\rangle\in\Delta\}|}{|\Sigma|}$.
+Let ${\mathcal{A}}=\left\langle\Sigma,Q,q_{0},\delta,\alpha\right\rangle$ be a deterministic automaton, and let $G_{\mathcal{A}}=\left\langle Q,E\right\rangle$ be its induced directed graph.
+A random walk on ${\mathcal{A}}$, is a random walk on $G_{\mathcal{A}}$ with the probability matrix $P(q,p)=\frac{|\{\sigma\in\Sigma:\left\langle q,\sigma,p\right\rangle\in\Delta\}|}{|\Sigma|}$.
 It is not hard to see that $\mathbb{P}(L({\mathcal{A}}))$ is precisely the probability that a random walk on ${\mathcal{A}}$ is an accepting run.
@@
 Consider a nondeterministic automaton ${\mathcal{A}}=\langle\Sigma,Q,q_{0},\delta,\alpha\rangle$.
-The automaton ${\mathcal{A}}$ is determinizable by prunning (DBP) if we can remove some of the transitions of ${\mathcal{A}}$ and get a deterministic automaton ${\mathcal{A}}^{\prime}$ that recognizes $L({\mathcal{A}})$.
+The automaton ${\mathcal{A}}$ is determinizable by pruning (DBP) if we can remove some of the transitions of ${\mathcal{A}}$ and get a deterministic automaton ${\mathcal{A}}^{\prime}$ that recognizes $L({\mathcal{A}})$.
 We then say that ${\mathcal{A}}^{\prime}$ is a deterministic pruning of ${\mathcal{A}}$.
@@
     The strategy $f$ is consistent with the transition function.
-That is, $f(\epsilon)=q_{0}$, and for every finite word $u\in\Sigma^{*}$ and letter $\sigma\in\Sigma$, we have that $\langle f(u),\sigma,f(u\cdot\sigma)\rangle\in\Delta$.
+That is, $f(\epsilon)=q_{0}$, and for every finite word $u\in\Sigma^{*}$ and letter $\sigma\in\Sigma$, we have that $\left\langle f(u),\sigma,f(u\cdot\sigma)\right\rangle\in\Delta$.
   2. (2)
@@
 Finally, if a state $q$ of an HD automaton has a $\sigma$-successor $s$ that does not accept all the words $w$ such that $\sigma\cdot w$ is in $L({\mathcal{A}}^{q})$, then the $\sigma$-transition from $q$ to $s$ can be removed.
-Moreover, by  [14, 17, 4], the detection of such transitions can be done in polynomial time.
+Moreover, by  [HKR02, KS15, BK18], the detection of such transitions can be done in polynomial time.
 Hence, we can assume that every HD automaton, and hence also every DBP automaton, is SD.
-[^3] Using the terminology of [17], we can assume that the residual languages of all the $\sigma$-successors of each state in an HD automaton are equivalent.
+[^3] Using the terminology of [KS15], we can assume that the residual languages of all the $\sigma$-successors of each state in an HD automaton are equivalent.
 For ${\rm X}\in\{{\rm B,C,W}\}$, we use SD-NXW, HD-NXW, and DBP-NXW to refer to the different classes of NXWs.
@@
 Then, as at least one $\sigma$-successor of the initial state of ${\mathcal{A}}$ is not empty, we have that ${\mathcal{C}}$ is an NXW that is not an SD-NXW.
-The relation between DBPness and HDness has already been studied.
-It is shown in [5] that HD-NXW need not be DBP for ${\rm X}\in\{{\rm B,C}\}$, and shown in [7] that HD-NWW are DBP.
+The relation between DBPness and HDness has already been studied. It is shown in [BKKS13] that HD-NXW need not be DBP for ${\rm X}\in\{{\rm B,C}\}$, and shown in [BKS17] that HD-NWW are DBP.
 It is left to relate HDness and SDness.
@@
 Proof 3.2.
-In [19, 17], the authors suggest variants of the subset construction that determinize HD-NBWs.
-As we argue below, the construction in [17] is correct also when applied to SD-NBWs.
+In [KSV06, KS15], the authors suggest variants of the subset construction that determinize HD-NBWs.
+As we argue below, the construction in [KS15] is correct also when applied to SD-NBWs.
 Moreover, it preserves weakness.
 Thus, DBW$=$SD-NBW and DWW$=$SD-NWW.
-Also, the last inequality follows from the fact DBW$<$NBW and DWW$<$NWW [22].
-Given an NBW ${\mathcal{A}}=\langle\Sigma,Q,q_{0},\delta,\alpha\rangle$, the DBW generated in [17] [^4] The construction in [17] assumes automata with transition-based acceptance, and (regardless of this) is slightly different: when $\alpha$ is visited, ${\mathcal{A}}^{\prime}$ continues with a single state from the set of successors.
+Also, the last inequality follows from the fact DBW$<$NBW and DWW$<$NWW [Lan69].
+Given an NBW ${\mathcal{A}}=\left\langle\Sigma,Q,q_{0},\delta,\alpha\right\rangle$, the DBW generated in [KS15] [^4] The construction in [KS15] assumes automata with transition-based acceptance, and (regardless of this) is slightly different: when $\alpha$ is visited, ${\mathcal{A}}^{\prime}$ continues with a single state from the set of successors.
 However, the key point remains the same: ${\mathcal{A}}$ being SD enables ${\mathcal{A}}^{\prime}$ to maintain only subsets of states, rather than Safra trees, which makes determinization much easier.
-is ${\mathcal{A}}^{\prime}=\langle\Sigma,Q^{\prime},q^{\prime}_{0},\delta^{\prime},\alpha^{\prime}\rangle$, where $Q^{\prime}=2^{Q}$, $q^{\prime}_{0}=\{q_{0}\}$, $\alpha^{\prime}=\{S\in 2^{Q}:S\subseteq\alpha\}$, and the transition function $\delta^{\prime}$ is defined for every subset $S\in 2^{Q}$ and letter $\sigma\in\Sigma$ as follows.
+is ${\mathcal{A}}^{\prime}=\left\langle\Sigma,Q^{\prime},q^{\prime}_{0},\delta^{\prime},\alpha^{\prime}\right\rangle$, where $Q^{\prime}=2^{Q}$, $q^{\prime}_{0}=\{q_{0}\}$, $\alpha^{\prime}=\{S\in 2^{Q}:S\subseteq\alpha\}$, and the transition function $\delta^{\prime}$ is defined for every subset $S\in 2^{Q}$ and letter $\sigma\in\Sigma$ as follows.
 If $\delta(S,\sigma)\cap\alpha=\emptyset$, then $\delta^{\prime}(S,\sigma)=\delta(S,\sigma)$.
@@
 Note that we refer here to the syntactic class (e.g., deciding whether a given NBW is HD) and not to the semantic one (e.g., deciding whether a given NBW has an equivalent HD-NBW).
-Indeed, by Theorem 2, the latter boils down to deciding whether the language of a given NXW, for ${\rm X}\in\{{\rm B,C,W}\}$, can be recognized by a DXW, which is well known: the answer is always “yes” for an NCW, and the problem is PSPACE-complete for NBWs and NWWs [20].
-[^5] The proof in [20] is for NBWs, yet the arguments there apply also for weak automata.
+Indeed, by Theorem 2, the latter boils down to deciding whether the language of a given NXW, for ${\rm X}\in\{{\rm B,C,W}\}$, can be recognized by a DXW, which is well known: the answer is always “yes” for an NCW, and the problem is PSPACE-complete for NBWs and NWWs [KV05a].
+[^5] The proof in [KV05a] is for NBWs, yet the arguments there apply also for weak automata.
 Our results are summarized in Table 1.
 The entries there describe the complexity of deciding whether a given NXW belongs to a certain nondeterministic level (for example, the upper left cell describes the complexity of deciding whether an NBW is DBP).
-In fact, the results described are valid already when the given NXW is known to be one level above the questioned one with only two exceptions – deciding the DBPness of an HD-NCW, whose NP-hardness is proved in Th.
-8, and deciding whether a given NWW is DBP, which is PTIME in general, and is $O(1)$ when the given NWW is HD, in which case the answer is always “yes”.
+In fact, the results described are valid already when the given NXW is known to be one level above the questioned one with only two exceptions – deciding the DBPness of an HD-NCW, whose NP-hardness is proved in Theorem 8, and deciding whether a given NWW is DBP, which is PTIME in general, and is $O(1)$ when the given NWW is HD, in which case the answer is always “yes”.
 DBP HD SD
 NBW NP-complete PTIME PSPACE-complete
-4 [4] 3
+4 [BK18] 3
 NCW NP-complete PTIME PSPACE-complete
-[16] [17] 3
+[KM18] [KS15] 3
 NWW PTIME PTIME PSPACE-complete
-[17, 4] [17, 4] 3
+[KS15, BK18] [KS15, BK18] 3
 Table 1.
-The complexity of deciding the nondeterminizm level of an NXW, for ${\rm X}\in\{{\rm B,C,W}\}$.
-The results are valid also in the case the given NXW is one level to the right in the nondeterminism hierarchy, with the exception of deciding the DBPness of an HD-NCW, whose NP-hardness is proved in Th.
-8, and deciding the DBPness of an HD-NWW, where the answer is always positive [7].
+The complexity of deciding the nondeterminism level of an NXW, for ${\rm X}\in\{{\rm B,C,W}\}$.
+The results are valid also in the case the given NXW is one level to the right in the nondeterminism hierarchy, with the exception of deciding the DBPness of an HD-NCW, whose NP-hardness is proved in Theorem 8, and deciding the DBPness of an HD-NWW, where the answer is always positive [BKS17].
 We start with the complexity of deciding semantic determinism.
@@
 Membership in PSPACE is easy, as we check SDness by polynomially many checks of language equivalence.
-Formally, given an NXW ${\mathcal{A}}=\langle\Sigma,Q,q_{0},\delta,\alpha\rangle$, a PSPACE algorithm goes over all states $q\in Q$, letters $\sigma$, and $\sigma$-successors $s$ and $s^{\prime}$ of $q$, and checks that $s\sim_{\mathcal{A}}s^{\prime}$.
-Since language equivalence can be checked in PSPACE [32] and there are polynomially many checks to perform, we are done.
+Formally, given an NXW ${\mathcal{A}}=\left\langle\Sigma,Q,q_{0},\delta,\alpha\right\rangle$, a PSPACE algorithm goes over all states $q\in Q$, letters $\sigma$, and $\sigma$-successors $s$ and $s^{\prime}$ of $q$, and checks that $s\sim_{\mathcal{A}}s^{\prime}$.
+Since language equivalence can be checked in PSPACE [SVW87] and there are polynomially many checks to perform, we are done.
 Proving PSPACE-hardness, we do a reduction from polynomial-space Turing machines.
-Given a Turing machine $T$ with space complexity $s:\mathbb{N}\rightarrow\mathbb{N}$, we construct in time polynomial in $|T|$ and $s(0)$, an NWW ${\mathcal{A}}$ of size linear in $T$ and $s(0)$, such that ${\mathcal{A}}$ is SD iff $T$ accepts the empty tape [^6] This is sufficient, as one can define a generic reduction from every language $L$ in PSPACE as follows.
+Given a Turing machine $T$ with $|T|$ many transitions and space complexity $s:\mathbb{N}\rightarrow\mathbb{N}$, we construct in time polynomial in $|T|$ and $s(0)$, an NWW ${\mathcal{A}}$ of size linear in $T$ and $s(0)$, such that ${\mathcal{A}}$ is SD iff $T$ accepts the empty tape [^6] This is sufficient, as one can define a generic reduction from every language $L$ in PSPACE as follows.
 Let $T_{L}$ be a Turing machine that decides $L$ in polynomial space $f(n)$.
@@
 It is not hard to see that if $T$ accepts the empty tape, then ${\mathcal{A}}$ is universal (that is, accepts all words).
-Indeed, each word $w$ is either not a suffix of an encoding of a legal computation of $T$ that uses at most $n_{0}$ cells, in which case $w$ is accepted thanks to C1, or
-$w$ is such a suffix, in which case, since we assume that $T$ halts from any configuration, the encoded computation eventually reaches a final configuration and, by the definition of $T$, continues by an encoding of the computation of $T$ on the empty tape.
+Indeed, each word $w$ is either not a suffix of an encoding of a legal computation of $T$ that uses at most $n_{0}$ cells, in which case $w$ is accepted thanks to C1, or $w$ is such a suffix, in which case, since we assume that $T$ halts from any configuration, the encoded computation eventually reaches a final configuration and, by the definition of $T$, continues by an encoding of the computation of $T$ on the empty tape.
 Thus, $w$ eventually includes the encoding of the computation of $T$ on the empty tape.
@@
 Also, whenever ${\mathcal{A}}$ waits to witness some behavior, namely, waits for a position where there is a violation of the encoding, or waits for an encoding of an accepting configuration of $T$ that uses $n_{0}$ cells, it may nondeterministically, upon reading the next letter, return to the initial state.
-It is not hard to see that ${\mathcal{A}}$ can be defined in size linear in $T$ and $n_{0}$.
+It is not hard to see that ${\mathcal{A}}$ can be defined in size linear in $|T|$ and $n_{0}$.
 As the only accepting states of ${\mathcal{A}}$ is the accepting sink, it is clearly weak, and in fact describes a co-safety language.
@@
 Thus, we conclude that $T$ accepts the empty tape iff ${\mathcal{A}}$ is SD. In Appendix B, we give the full technical details of the construction of ${\mathcal{A}}$.
+Note that the considerations in the proof of Theorem 3 can be used in order to prove that the universality problem for SD-NBWs is PSPACE-hard [AK23].
 Theorem 4.
@@
 Otherwise, it returns to the two-state loop.
-Now, we define ${\mathcal{A}}_{\varphi}=\langle\Sigma_{n,m},Q_{\varphi},p,\delta_{\varphi},\{q_{0}\}\rangle$, where $Q_{\varphi}=Q_{n,m}\cup\{q^{i}_{k}:(i,k)\in\{0,1\}\times[n]\}$.
+Now, we define ${\mathcal{A}}_{\varphi}=\left\langle\Sigma_{n,m},Q_{\varphi},p,\delta_{\varphi},\{q_{0}\}\right\rangle$, where $Q_{\varphi}=Q_{n,m}\cup\{q^{i}_{k}:(i,k)\in\{0,1\}\times[n]\}$.
 The idea behind ${\mathcal{A}}_{\varphi}$ is as follows.
@@
 When the guess is not successful, it returns to $p$.
-Implementing the above intuition, transitions from the states $Q_{n,m}\setminus\{p\}$ are inherited from $D_{n,m}$, and transitions from the states in $\{q^{i}_{k}:(i,k)\in\{0,1\}\times[n]\}\cup\{p\}$ are defined as follows (see also Figure 3).
+Implementing the above intuition, transitions from the states $Q_{n,m}\setminus\{p\}$ are inherited from $D_{n,m}$, and transitions from the states in $\{q^{i}_{k}:(i,k)\in\left\{0,1\right\}\times[n]\}\cup\{p\}$ are defined as follows (see also Figure 3).
   • For all $k\in[n]$, we have that $\delta_{\varphi}(p,x_{k})=\{q^{0}_{k},q^{1}_{k}\}$.
@@
 In Appendix C we formally prove that there is a one to one correspondence between prunings of ${\mathcal{A}}_{\varphi}$ and assignments to $\varphi$, and that an assignment satisfies $\varphi$ iff the corresponding pruning recognizes $L_{n,m}$, implying Proposition 7.
-We continue to co-Büchi automata. In [16], the authors prove that deciding the DBPness of a given NCW is NP-complete. Below we extend their proof to HD-NCWs.
+We continue to co-Büchi automata. In [KM18], the authors prove that deciding the DBPness of a given NCW is NP-complete. Below we extend their proof to HD-NCWs.
 Theorem 8.
@@
 Proof 4.6.
-The upper bound is as in [16].
-For the lower bound, we analyze the reduction in [16] and prove that the NCW constructed there is HD.
-The reduction is from the Hamiltonian-cycle problem: given a connected graph $G=\langle[n],E\rangle$, the reduction outputs an NCW ${\mathcal{A}}_{G}$ over the alphabet $[n]$ that is obtained from $G$ by adding self loops to all vertices, labelling the loop at a vertex $i$ by the letter $i$, and labelling the edges from vertex $i$ to all its neighbours in $G$ by every letter $j\neq i$.
-Then, the co-Büchi condition requires a run to eventually get stuck at a self-loop [^7] The exact reduction is more complicated and involves an additional letter #normal-#\## that forces each deterministic pruning of ${\mathcal{A}}_{G}$ to proceed to the same neighbour of $i$ upon reading a letter $j\neq i$ from the vertex $i$..
+The upper bound follows by Lemma 18 in [KM18] where the authors prove that checking if an automaton is DBP is in NP already for general NCWs.
+For the lower bound, we analyze the reduction in [KM18] and prove that the NCW constructed there is HD.
+The reduction is from the Hamiltonian-cycle problem: given a connected graph $G=\left\langle[n],E\right\rangle$, the reduction outputs an NCW ${\mathcal{A}}_{G}$ over the alphabet $[n]$ that is obtained from $G$ by adding self loops to all vertices, labelling the loop at a vertex $i$ by the letter $i$, and labelling the edges from vertex $i$ to all its neighbours in $G$ by every letter $j\neq i$.
+Then, the co-Büchi condition requires a run to eventually get stuck at a self-loop [^7] The exact reduction is more complicated and involves an additional letter ##\## that forces each deterministic pruning of ${\mathcal{A}}_{G}$ to proceed to the same neighbour of $i$ upon reading a letter $j\neq i$ from the vertex $i$..
 Accordingly, $L({\mathcal{A}}_{G})=[n]^{*}\cdot\bigcup_{i\in[n]}i^{\omega}$.
@@
 We say that ${\mathcal{A}}$ is almost-DBP if it can be pruned to a deterministic automaton that retains almost all the words in $L({\mathcal{A}})$.
-[^8] Not to be confused with the definition of almost-DBPness in [2], where it is used to describe automata that are deterministic in their transition function yet may have a set of initial states.
+[^8] Not to be confused with the definition of almost-DBPness in [AKL10], where it is used to describe automata that are deterministic in their transition function yet may have a set of initial states.
 Formally, an NXW ${\mathcal{A}}$ is almost-DBP if there is a DXW ${\mathcal{A}}^{\prime}$ embodied in ${\mathcal{A}}$ such that $\mathbb{P}(L({\mathcal{A}})\setminus L({\mathcal{A}}^{\prime}))=0$.
@@
 We first show that Eve has a probabilistic strategy to win ${\mathcal{G}}_{\mathcal{A}}$ with probability $1$, even without assuming that ${\mathcal{A}}$ is semantically deterministic.
-Consider the probabilistic strategy $g$ where from $(q,\sigma)$, Eve picks one of the $\sigma$-successors of $q$ uniformly by random.
+Consider the probabilistic strategy $g$ where from $(q,\sigma)$, Eve picks one of the $\sigma$-successors of $q$ uniformly at random.
 Note that this strategy is memoryless, and hence the outcome of the game can be thought as a random walk in ${\mathcal{A}}$ that starts at $q_{0}$ and gives positive probabilities to all transitions.
@@
 Note that $w_{f}\notin L({\mathcal{A}})$ always implies that $r_{f}$ is rejecting.
-We show that if ${\mathcal{A}}$ is SD, then for every winning strategy $f$ for Eve, we have that $r_{f}$ is correct with probability $1$.
-Consider a winning strategy $f$ for Eve.
-Since $f$ is winning for Eve, we have that with probability 1, either $r_{f}$ is an accepting run or $\inf(r_{f})\subseteq Q_{rej}$.
-Thus, it is sufficient to prove that if $\inf(r_{f})\subseteq Q_{rej}$, then $w_{f}\in L({\mathcal{A}})$ with probability 00, since otherwise, almost surly $r_{f}$ is accepting, and in particular is correct.
+We show that if ${\mathcal{A}}$ is SD, then for every almost-sure winning strategy $f$ for Eve, we have that $r_{f}$ is correct with probability $1$.
+Consider an almost-sure winning strategy $f$ for Eve.
+Since $f$ is almost-sure winning for Eve, we have that with probability 1, either $r_{f}$ is an accepting run or $\inf(r_{f})\subseteq Q_{rej}$.
+Thus, it is sufficient to prove that if $\inf(r_{f})\subseteq Q_{rej}$, then $w_{f}\in L({\mathcal{A}})$ with probability 00, since otherwise, almost surely $r_{f}$ is accepting, and in particular is correct.
 By semantic determinism, for all $i\geq 0$, it holds that $w_{f}\in L({\mathcal{A}})$ iff $w_{f}[i+1,\infty]\in L({\mathcal{A}}^{q_{f}^{i}})$.
-Moreover, the infix $w_{f}[i+1,\infty]$ is independent of $q^{i}_{f}$, and hence for all $q\in Q$ and $i\geq 0$, the event $w_{f}[i+1,\infty]\in L({\mathcal{A}}^{q})$ is independent of $q^{i}_{f}$.
+Moreover, the suffix $w_{f}[i+1,\infty]$ is independent of $q^{i}_{f}$, and hence for all $q\in Q$ and $i\geq 0$, the event $w_{f}[i+1,\infty]\in L({\mathcal{A}}^{q})$ is independent of $q^{i}_{f}$.
 Thus, for all $q\in Q$ and $i\geq 0$, it holds that $\mathbb{P}(w_{f}\in L({\mathcal{A}})\>|\>q_{f}^{i}=q)=\mathbb{P}(w_{f}[i+1,\infty]\in L({\mathcal{A}}^{q}))=\mathbb{P}(L({\mathcal{A}}^{q}))$.
@@
 Overall, we showed that if $f$ is winning for Eve and ${\mathcal{A}}$ is SD, then $\mathbb{P}(r_{f}\text{ is correct})=1$.
-Hence, by pure memoryless determinacy of simple stochastic parity games [10], we may consider a pure memoryless winning strategy $f$ for Eve in ${\mathcal{G}}_{\mathcal{A}}$, and by the above we know that a random walk in the corresponding pruned DBW is correct with probability 1.
+Hence, by pure memoryless determinacy of simple stochastic parity games [CJH03], we may consider a pure memoryless winning strategy $f$ for Eve in ${\mathcal{G}}_{\mathcal{A}}$, and by the above we know that a random walk in the corresponding pruned DBW is correct with probability 1.
 Formally, since $f$ is pure memoryless, it induces a pruning of ${\mathcal{A}}$.
@@
 Assume without loss of generality that ${\mathcal{D}}$ has a single state $q$ with $L({\mathcal{D}}^{q})=\Sigma^{\omega}$, in particular, $C=\{q\}$ is the only ergodic $\alpha$-free SCC of ${\mathcal{D}}$.
-Then, for every word $w\in\Sigma^{\omega}$, we have that $w\in\mbox{\it co-safe}(L)$ iff the run of ${\mathcal{D}}$ on $w$ reaches $C$.
-Hence, the probability that $w\in L\setminus\mbox{\it co-safe}(L)$ equals the probability that $inf(r)$ is $\alpha$-free but is not an ergodic SCC of ${\mathcal{D}}$.
-Since the later happens w.p 0, we have that $\mathbb{P}(L\setminus\mbox{\it co-safe}(L))=0$.
+Then, for every word $w\in\Sigma^{\omega}$, we have that $w\in\mbox{\it co-safe}(L)$ iff the run $r$ of ${\mathcal{D}}$ on $w$ reaches $C$.
+Hence, the probability that $w\in L\setminus\mbox{\it co-safe}(L)$ equals the probability that $\inf(r)$ is $\alpha$-free but is not an ergodic SCC of ${\mathcal{D}}$.
+Since the later happens with probability 0, we have that $\mathbb{P}(L\setminus\mbox{\it co-safe}(L))=0$.
 […]
@@
 Proof 5.5.
-Let ${\mathcal{A}}=\langle\Sigma,Q,q_{0},\delta,\alpha\rangle$ be an HD-NCW.
-Consider the NCW ${\mathcal{A}}^{\prime}=\langle\Sigma,Q,q_{0},\delta,\alpha^{\prime}\rangle$, where $\alpha^{\prime}=\alpha\cup\{q\in Q:L({\mathcal{A}}^{q})\neq\Sigma^{\omega}\}$.
+Let ${\mathcal{A}}=\left\langle\Sigma,Q,q_{0},\delta,\alpha\right\rangle$ be an HD-NCW.
+Consider the NCW ${\mathcal{A}}^{\prime}=\left\langle\Sigma,Q,q_{0},\delta,\alpha^{\prime}\right\rangle$, where $\alpha^{\prime}=\alpha\cup\{q\in Q:L({\mathcal{A}}^{q})\neq\Sigma^{\omega}\}$.
 We prove that ${\mathcal{A}}^{\prime}$ is an HD-NCW with $L({\mathcal{A}}^{\prime})=\mbox{\it co-safe}(L({\mathcal{A}}))$.
@@
 When the automaton is nondeterministic, this information is lost, and thus reasoning about MDPs involves determinization.
-An automaton is said to be good for MDPs (GFM), if its product with MDPs maintains the probability of acceptance, and can therefore replace deterministic automata when reasoning about stochastic behaviors [13, 30].
+An automaton is said to be good-for-MDPs (GFM), if its product with MDPs maintains the probability of acceptance, and can therefore replace deterministic automata when reasoning about stochastic behaviors [HPS+20, STZ22].
 In Section 5, we related semantic determinism with almost-DBPness.
 In this section we relate semantic determinism with GFMness.
+The terminology used in this section is based on [HPS+20].
 6.1. Good-for-MDP automata
 We first define MDPs and good-for-MDP automata.
-A $\Sigma$-labeled Markov decision process (MDP) is ${\mathcal{M}}=\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{{\mathcal{M}}},\Sigma,\tau\rangle$, where $S$ is a finite set of states, $s_{0}\in S$ is an initial state, and for every $s\in S$, the set $A_{s}$ is a finite set of actions that are available in $s$.
+A $\Sigma$-labeled Markov decision process (MDP) is ${\mathcal{M}}=\left\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{{\mathcal{M}}},\Sigma,\tau\right\rangle$, where $S$ is a finite set of states, $s_{0}\in S$ is an initial state, and for every $s\in S$, the set $A_{s}$ is a finite set of actions that are available in $s$.
 Let $A=\bigcup_{s\in S}A_{s}$ be the set of all actions.
@@
 Finally, $\Sigma$ is a finite alphabet, and $\tau:S\rightarrow\Sigma$ is a labeling function on the states.
-Consider a nondeterministic automaton ${\mathcal{A}}$ with alphabet $\Sigma$, and a $\Sigma$-labeled MDP ${\mathcal{M}}$, and consider the following game between a player and nature.
-The positions of the game are pairs $\langle s,q\rangle$, where $s\in S$ is a state of ${\mathcal{M}}$, and $q\in Q$ is a state of ${\mathcal{A}}$.
-The game starts by the player picking an action $a\in A_{s_{0}}$ available from the initial state of ${\mathcal{M}}$, and the game moves to $\langle s,q_{0}\rangle$ with probability $\mathrm{Pr}_{\mathcal{M}}(s_{0},a,s)$.
-Then, at each round from position $\langle s,q\rangle$, the player chooses an action $a\in A_{s}$ and a state $q^{\prime}\in\delta(q,\tau(s))$, and then nature chooses a state $s^{\prime}\in S$ with probability $\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})$, and the game moves to $\langle s^{\prime},q^{\prime}\rangle$.
-The objective of the player is to maximize the probability of generating an accepting run of ${\mathcal{A}}$.
-When the automaton ${\mathcal{A}}$ is deterministic, the player has to only pick actions, as the state $q^{\prime}\in\delta(q,\tau(s))$ is unique.
-When the automaton is nondeterministic, the player may fail to generate an accepting run of ${\mathcal{A}}$ even when the actions he provides induce a word in $L({\mathcal{A}})$.
-Indeed, each nondeterministic choice may lead to accepting only a subset of the words in $L({\mathcal{A}})$.
-Intuitively, ${\mathcal{A}}$ is Good-For-MDP, if for every MDP ${\mathcal{M}}$, the nondeterminism in ${\mathcal{A}}$ does not prevent the player from maximizing the probability of generating an accepting run.
+Consider a nondeterministic automaton ${\mathcal{N}}$ with alphabet $\Sigma$, and a $\Sigma$-labeled MDP ${\mathcal{M}}$, and consider the following game between a player and nature.
+The positions of the game are pairs $\left\langle s,q\right\rangle$, where $s\in S$ is a state of ${\mathcal{M}}$, and $q\in Q$ is a state of ${\mathcal{N}}$.
+The game starts by the player picking an action $a\in A_{s_{0}}$ available from the initial state of ${\mathcal{M}}$, and the game moves to $\left\langle s,q_{0}\right\rangle$ with probability $\mathrm{Pr}_{\mathcal{M}}(s_{0},a,s)$.
+Then, at each round from position $\left\langle s,q\right\rangle$, the player chooses an action $a\in A_{s}$ and a state $q^{\prime}\in\delta(q,\tau(s))$, and then nature chooses a state $s^{\prime}\in S$ with probability $\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})$, and the game moves to $\left\langle s^{\prime},q^{\prime}\right\rangle$.
+The objective of the player is to on the one hand maximize the probability of generating a word in $L({\mathcal{N}})$ and on the other hand, to generate a rejecting run on words in $L({\mathcal{N}})$ with zero probability.
+When the automaton ${\mathcal{N}}$ is deterministic, the player has to only pick actions, as the state $q^{\prime}\in\delta(q,\tau(s))$ is unique.
+When the automaton is nondeterministic, the player may fail to generate an accepting run of ${\mathcal{N}}$ even when the actions he provides induce a word in $L({\mathcal{N}})$.
+Indeed, each nondeterministic choice may lead to accepting only a subset of the words in $L({\mathcal{N}})$.
+Intuitively, ${\mathcal{N}}$ is good-for-MDP, if for every MDP ${\mathcal{M}}$, the nondeterminism in ${\mathcal{N}}$ does not prevent the player from maximizing the probability of generating a word in $L({\mathcal{N}})$ while generating a rejecting run on words in $L({\mathcal{N}})$ with zero probability.
 We now formalize this intuition.
-A run of ${\mathcal{M}}$ is an infinite sequence of states $r=\langle s_{0},s_{1},s_{2},\ldots\rangle\in S^{\omega}$, such that for all $i\geq 0$ there exists $a\in A_{s_{i}}$ with $\mathrm{Pr}_{\mathcal{M}}(s_{i},a,s_{i+1})>0$.
+A run of ${\mathcal{M}}$ is an infinite sequence of states $r=\left\langle s_{0},s_{1},s_{2},\ldots\right\rangle\in S^{\omega}$, such that for all $i\geq 0$ there exists $a\in A_{s_{i}}$ with $\mathrm{Pr}_{\mathcal{M}}(s_{i},a,s_{i+1})>0$.
 A finite run is a prefix of such a sequence.
-Let $\Omega({\mathcal{M}})$ denote the set of runs of ${\mathcal{M}}$ from the initial state $s_{0}$, and let $\Pi({\mathcal{M}})$ denote the set of finite such runs $\langle s_{0},s_{1},\ldots,s_{k}\rangle\in S^{k+1}$, for some $k\geq 0$.
+Let $\Omega({\mathcal{M}})\subseteq S^{\omega}$ denote the set of runs of ${\mathcal{M}}$ from the initial state $s_{0}$, and let $\Pi({\mathcal{M}})\subseteq S^{*}$ denote the set of finite such runs.
 For a run $r\in\Omega({\mathcal{M}})$, we define the corresponding labeled run as $\tau(r)=\tau(s_{1}),\tau(s_{2}),\tau(s_{3}),\ldots\in\Sigma^{\omega}$.
@@
 Let $\Pi({\mathcal{M}},\mu)$ denote the subset of $\Pi({\mathcal{M}})$ that includes exactly all finite runs that agree with the strategy $\mu$.
-Formally, we have that $\langle s_{0}\rangle\in\Pi({\mathcal{M}},\mu)$, and for all $p^{\prime}=\langle s_{0},s_{1},\ldots,s_{k}\rangle\in\Pi({\mathcal{M}},\mu)$, and $s_{k+1}\in S$, we have that $p=p^{\prime}\cdot s_{k+1}\in\Pi({\mathcal{M}},\mu)$ iff there exists $a_{k+1}\in A_{s_{k}}$ such that $\mu(p^{\prime})(a_{k+1})>0$ and $\mathrm{Pr}_{{\mathcal{M}}}(s_{k},a_{k+1},s_{k+1})>0$.
+Formally, we have that $\left\langle s_{0}\right\rangle\in\Pi({\mathcal{M}},\mu)$, and for all $p^{\prime}=\left\langle s_{0},s_{1},\ldots,s_{k}\right\rangle\in\Pi({\mathcal{M}},\mu)$, and $s_{k+1}\in S$, we have that $p=p^{\prime}\cdot s_{k+1}\in\Pi({\mathcal{M}},\mu)$ iff there exists $a_{k+1}\in A_{s_{k}}$ such that $\mu(p^{\prime})(a_{k+1})>0$ and $\mathrm{Pr}_{{\mathcal{M}}}(s_{k},a_{k+1},s_{k+1})>0$.
 Thus, a run in $\Pi({\mathcal{M}},\mu)$ can only be extended using actions that have a positive probability according to $\mu$.
@@
 We say that a strategy $\mu$ is pure if for all $p\in\Pi({\mathcal{M}})$, the distribution $\mu(p)$ is a point distribution, and is mixed otherwise.
-We say that $\mu$ has finite-memory, if there exists a finite set $M$ of memories, with a distinguished initial memory $m_{0}\in M$, and a memory update function $\eta:M\times S\to M$ such that $\mu$ only depends on the current memory and MDP state.
-Formally, $\eta$ is extended to memories and finite sequences in $S^{*}$ as follows: $\eta:M\times S^{*}\to M$ is such that for every memory $m\in M$, finite sequence $p\in S^{*}$, and MDP state $s\in S$, we have that $\eta(m,\epsilon)=m$ and $\eta(m,p\cdot s)=\eta(\eta(m,p),s)$.
-Then, for all $p,p^{\prime}\in\Pi({\mathcal{M}},\mu)$ and $s\in S$ such that $p\cdot s,p^{\prime}\cdot s\in\Pi({\mathcal{M}},\mu)$, if $\eta(m_{0},p)=\eta(m_{0},p^{\prime})$ then $\mu(p\cdot s)=\mu(p^{\prime}\cdot s)$.
-Equivalently, $\mu$ can be thought as a function $\mu:M\times S\rightarrow\mathrm{dist}(A)\times M$, that given the current memory $m$ and the MDP state $s$, outputs a distribution on the available actions from $s$ and the next memory state to move to.
-Indeed, we say that $\langle m,s\rangle\in M\times S$ are feasible by $\mu$ if there exists $p\in\Pi({\mathcal{M}},\mu)$ such that $p\cdot s\in\Pi({\mathcal{M}},\mu)$ and $\eta(m_{0},p)=m$.
-Then for every feasible $\langle m,s\rangle\in M\times S$, let $p_{m,s}\in\Pi({\mathcal{M}},\mu)$ be a witness for feasibility.
-We can now define $\mu^{\prime}:M\times S\nrightarrow\mathrm{dist}(A)\times M$ as a partial function on all feasible $\langle m,s\rangle\in M\times S$ by $\mu^{\prime}(m,s)=\langle\mu(p_{m,s}\cdot s),\eta(m,s)\rangle$.
-Observe that $\mu(p\cdot s)$ only depends on $s$ and the memory $m=\eta(m_{0},p)$.
-That is, for all $p\cdot s\in\Pi({\mathcal{M}},\mu)$ we have that $\mu(p\cdot s)=\mu(p_{m,s}\cdot s)$ for $m=\eta(m_{0},p)$.
-Thus, if $\mu^{\prime}(m,s)=\langle d,m^{\prime}\rangle$, then $\mu(p\cdot s)=d$ for all $p\cdot s\in\Pi({\mathcal{M}},\mu)$ such that $\eta(m_{0},p)=m$.
+We say that $\mu$ has finite-memory, if there exists a finite set $M$ of memories, with a distinguished initial memory $m_{0}\in M$, and a memory update function $\eta:S\times M\to M$ such that $\mu$ only depends on the current memory and MDP state.
+Formally, $\eta$ is extended to finite sequences in $S^{*}$ as follows: $\eta^{*}:S^{*}\to M$ is such that for every finite sequence $p\in S^{*}$ and MDP state $s\in S$, we have that $\eta^{*}(\epsilon)=m_{0}$ and $\eta^{*}(p\cdot s)=\eta(s,\eta^{*}(p))$.
+Then, for all $s\in S$ and $p\cdot s,p^{\prime}\cdot s\in\Pi({\mathcal{M}},\mu)$, if $\eta^{*}(p\cdot s)=\eta^{*}(p^{\prime}\cdot s)$ then $\mu(p\cdot s)=\mu(p^{\prime}\cdot s)$.
+Equivalently, $\mu$ can be thought as a function $\mu:S\times M\rightarrow\mathrm{dist}(A)\times M$, that for a given the current MDP state $s$ and memory $m$, outputs a distribution on the available actions from $s$ and the next memory state to move to from $m$.
+Indeed, we show that every finite memory strategy $\mu$ with memory set $M$ can be represented as a strategy of the form $\mu^{\prime}:S\times M\rightarrow\mathrm{dist}(A)\times M$.
+We say that $\left\langle s,m\right\rangle\in S\times M$ is feasible by $\mu$ if there is a finite run $p_{m,s}$ such that $p_{m,s}\cdot s\in\Pi({\mathcal{M}},\mu)$ and $\eta^{*}(p_{m,s}\cdot s)=m$.
+In other words, there is a positive probability that when playing according to $\mu$ we reach the memory state $m$ when being in state $s$ in the MDP.
+We can now define $\mu^{\prime}:S\times M\nrightarrow\mathrm{dist}(A)\times M$ as a partial function on all feasible $\left\langle s,m\right\rangle\in S\times M$ by $\mu^{\prime}(s,m)=\left\langle\mu(p_{m,s}\cdot s),\eta(s,m)\right\rangle$.
+Observe that $\mu(p_{m,s}\cdot s)$ is independent of the choice of witness $p_{m,s}$.
+That is, for all $p\cdot s\in\Pi({\mathcal{M}},\mu)$ with $\eta^{*}(p\cdot s)=m$ we have that $\mu(p\cdot s)=\mu(p_{m,s}\cdot s)$.
+Thus, if $\mu^{\prime}(s,m)=\left\langle d,m^{\prime}\right\rangle$, then $\mu(p\cdot s)=d$ for all $p\cdot s\in\Pi({\mathcal{M}},\mu)$ such that $\eta^{*}(p\cdot s)=m$.
 We say that $\mu$ is memoryless if it has a finite memory $M$ with $|M|=1$.
@@
 Namely, if $\mu$ only depends on the current state, and not on the history beforehand.
-A Markov Chain (MC, for short) is an MDP with a dummy set of actions, $A_{s}=\{\ast\}$ for all $s\in S$.
-Formally, an MC is ${\mathcal{C}}=\langle S,s_{0},\mathrm{Pr}_{\mathcal{C}},\Sigma,\tau\rangle$, where $S$ is the set of states, $s_{0}\in S$ is the initial state, $\mathrm{Pr}_{\mathcal{C}}:S\times S\rightarrow[0,1]$ is a stochastic transition function: for every $s\in S$ it holds that $\sum_{s^{\prime}\in S}\mathrm{Pr}_{\mathcal{C}}(s,s^{\prime})=1$.
+A Markov Chain (MC, for short) is an MDP with a dummy set of actions, $A_{s}=\left\{\ast\right\}$ for all $s\in S$.
+Formally, an MC is ${\mathcal{C}}=\left\langle S,s_{0},\mathrm{Pr}_{\mathcal{C}},\Sigma,\tau\right\rangle$, where $S$ is the set of states, $s_{0}\in S$ is the initial state, $\mathrm{Pr}_{\mathcal{C}}:S\times S\rightarrow[0,1]$ is a stochastic transition function: for every $s\in S$ it holds that $\sum_{s^{\prime}\in S}\mathrm{Pr}_{\mathcal{C}}(s,s^{\prime})=1$.
 A random process of ${\mathcal{C}}$, is an infinite sequence of random variables $S_{0},S_{1},S_{2},\ldots$ with values in $S$, such that $S_{0}=s_{0}$ with probability 1, and for all $s_{1},s_{2},\ldots,s_{i+1}\in S$, $i\geq 0$, it holds that $\mathbb{P}_{\mathcal{C}}(S_{i+1}=s_{i+1}|S_{0}=s_{0},\ldots,S_{i}=s_{i})=\mathrm{Pr}_{\mathcal{C}}(s_{i},s_{i+1})$.
-Given a strategy $\mu$, we can obtain from ${\mathcal{M}}$ and $\mu$ the MC ${\mathcal{C}}({\mathcal{M}},\mu)=\langle\Pi({\mathcal{M}},\mu),s_{0},\mathrm{Pr}_{\mathcal{M}}^{\mu}\rangle$ in which the choice of actions is resolved according to $\mu$.
-Formally, if $p,p^{\prime}\in\Pi({\mathcal{M}},\mu)$ are such that there are $s,s^{\prime}\in S$, such that $p^{\prime}$ ends with $s^{\prime}$ and $p=p^{\prime}\cdot s$, then $\mathrm{Pr}_{\mathcal{M}}^{\mu}(p^{\prime},p)=\sum_{a\in A_{s^{\prime}}}\mu(p^{\prime})(a)\cdot\mathrm{Pr}_{\mathcal{M}}(s^{\prime},a,s)$.
-It is clear that when $\mu$ is memoryless, that is, $\mu$ only depends on the last position in $p^{\prime}$, then ${\mathrm{Pr}}_{\mathcal{M}}^{\mu}$ only depends on $s^{\prime}$ and $s$, and hence the projection of the chain ${\mathcal{C}}({\mathcal{M}},\mu)$ onto the last position in a path is a finite MC.
-Formally, let $\iota:\Pi({\mathcal{M}},\mu)\rightarrow S$ be the function that maps a finite path to its last state.
-Let $S^{\prime}=\iota(\Pi({\mathcal{M}},\mu))$ be the set of all states for which there is a positive probability to reach when playing with $\mu$, and for every $s\in S^{\prime}$ let $p_{s}\in\Pi({\mathcal{M}},\mu)$ be some path with $\iota(p_{s})=s$.
-Let ${\mathrm{Pr}^{\prime}}_{{\mathcal{M}}}^{\mu}:S^{\prime}\times S^{\prime}\rightarrow[0,1]$ be the stochastic transition function that is defined by ${\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}(s,s^{\prime})=\sum_{a\in A_{s}}\mu(p_{s})(a)\cdot\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})$.
-Then, if $\langle p_{0},p_{1},p_{2},\ldots\rangle\in(\Pi({\mathcal{M}},\mu))^{\omega}$ is an infinite sequence of increasing paths that are sampled according to ${\mathcal{C}}({\mathcal{M}},\mu)$, then the sequence $\langle\iota(p_{0}),\iota(p_{1}),\iota(p_{2}),\ldots\rangle\in S^{\omega}$ satisfies $\mathbb{P}(\iota(p_{i+1})=s^{\prime}|\iota(p_{i})=s)={\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}(s,s^{\prime})$ for all $s^{\prime},s\in S^{\prime}$.
-Let ${\mathcal{M}}=\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{\mathcal{M}},\Sigma,\tau\rangle$ be an MDP and ${\mathcal{A}}=\langle\Sigma,Q,q_{0},\delta,\alpha\rangle$ be a nondeterministic automaton.
+Given a strategy $\mu$, we can obtain from ${\mathcal{M}}$ and $\mu$ the MC ${\mathcal{C}}({\mathcal{M}},\mu)=\left\langle\Pi({\mathcal{M}},\mu),s_{0},\mathrm{Pr}_{\mathcal{M}}^{\mu}\right\rangle$ in which the choice of actions is resolved according to $\mu$.
+Formally, for all $p\in\Pi({\mathcal{M}},\mu)$ and $s_{1},s_{2}\in S$ such that both $p_{1}=p\cdot s_{1}$ and $p_{2}=p_{1}\cdot s_{2}$ belong to $\Pi({\mathcal{M}},\mu)$, then $\mathrm{Pr}_{\mathcal{M}}^{\mu}(p_{1},p_{2})=\sum_{a\in A_{s_{1}}}\mu(p_{1})(a)\cdot\mathrm{Pr}_{\mathcal{M}}(s_{1},a,s_{2})$.
+It is clear that when $\mu$ is memoryless, that is, $\mu$ only depends on the last position in $p_{1}$, then ${\mathrm{Pr}}_{\mathcal{M}}^{\mu}$ only depends on $s_{1}$ and $s_{2}$, and hence the projection of the chain ${\mathcal{C}}({\mathcal{M}},\mu)$ onto the last position in a path is a finite MC.
+Formally, let $\iota:\Pi({\mathcal{M}},\mu)\rightarrow S$ be the function that maps a finite path to its last state and let $d_{s}\in\mathrm{dist}({\mathcal{A}})$ be the unique distribution such that $\mu(p)=d_{s}$ for all $p\in\Pi({\mathcal{M}},\mu)$ with $\iota(p)=s$.
+Note that if no such $p\in\Pi({\mathcal{M}},\mu)$ with $\iota(p)=s$ exists, then we arbitrarily take $d_{s}$ to be some distribution in $A_{s}$.
+Let ${\mathrm{Pr}^{\prime}}_{{\mathcal{M}}}^{\mu}:S\times S\rightarrow[0,1]$ be the stochastic transition function that is defined by ${\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}(s,s^{\prime})=\sum_{a\in A_{s}}d_{s}(a)\cdot\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})$.
+Then, if $p_{0},p_{1},p_{2},\ldots\in(\Pi({\mathcal{M}},\mu))^{\omega}$ is an infinite sequence of increasing paths that are sampled according to ${\mathcal{C}}({\mathcal{M}},\mu)$, then it is not hard to see that the sequence $\iota(p_{0}),\iota(p_{1}),\iota(p_{2}),\ldots\in S^{\omega}$ satisfies $\mathbb{P}(\iota(p_{i+1})=s^{\prime}|\iota(p_{i})=s)={\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}(s,s^{\prime})$ for all $s^{\prime},s\in S$.
+In other words, the sequence $\iota(p_{0}),\iota(p_{1}),\iota(p_{2}),\ldots$ is sampled according to the finite markov chain $\left\langle S,s_{0},{\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}\right\rangle$.
+Let ${\mathcal{M}}=\left\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{\mathcal{M}},\Sigma,\tau\right\rangle$ be an MDP and ${\mathcal{N}}=\left\langle\Sigma,Q,q_{0},\delta,\alpha\right\rangle$ be a nondeterministic automaton.
 When the player takes actions in ${\mathcal{M}}$ the infinite word $\tau(r)\in\Sigma^{\omega}$ is generated.
-Assume that, simultaneously, the player also generates a run of ${\mathcal{A}}$.
-This process can be modeled by an MDP that is a product of ${\mathcal{M}}$ with ${\mathcal{A}}$, and in each step the player takes an action of ${\mathcal{M}}$ and chooses a successor state in ${\mathcal{A}}$ that extends the current run on the new letter drawn by ${\mathcal{M}}$.
-Formally, we define the MDP ×${\mathcal{M}}\times{\mathcal{A}}=\langle S^{\prime},\langle s_{0},\bot\rangle,(A_{\langle s,q\rangle})_{\langle s,q\rangle\in S^{\prime}},\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{A}}},\Sigma,\gamma\rangle$ that is defined as follows.
-The set of states of ×${\mathcal{M}}\times{\mathcal{A}}$ is $S^{\prime}=(S\times Q)\cup\{\langle s_{0},\bot\rangle\}$, for $\bot\notin Q$.
-For all $\langle s,q\rangle\in S\times Q$, the set of actions $A_{\langle s,q\rangle}$ available from $\langle s,q\rangle$ consists of all pairs $\langle a,q^{\prime}\rangle\in A_{s}\times\delta(q,\tau(s))$.
-In addition, the actions available from $\langle s_{0},\bot\rangle$ are $A_{\langle s_{0},\bot\rangle}=A_{s_{0}}\times\{q_{0}\}$.
-Then, for all $\langle s,\ast\rangle\in S^{\prime}$ and $\langle a,q^{\prime}\rangle\in A_{\langle s,\ast\rangle}$, we have $\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{A}}}(\langle s,\ast\rangle,\langle a,q^{\prime}\rangle,\langle s^{\prime},q^{\prime}\rangle)=\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$ and $\gamma(s,\ast)=\tau(s)$.
-Given a finite or infinite run $r\in\Pi({\mathcal{M}}\times{\mathcal{A}})\cup\Omega({\mathcal{M}}\times{\mathcal{A}})$, $r=\langle\langle s_{0},\bot\rangle,\langle s_{1},q_{0}\rangle,\langle s_{2},q_{1}\rangle,\ldots\rangle$, we define $\pi_{\mathcal{A}}(r)=\langle q_{0},q_{1},q_{2},\ldots\rangle$, and $\pi_{\mathcal{M}}(r)=\langle s_{0},s_{1},s_{2},\ldots\rangle$ to be the ${\mathcal{A}}$-run and ${\mathcal{M}}$-run that correspond to $r$ respectively.
-Conversely, for $r_{\mathcal{M}}=\langle s_{0},s_{1},s_{2},\ldots\rangle\in S^{*}\cup S^{\omega}$ and $r_{\mathcal{A}}=\langle q_{0},q_{1},q_{2},\ldots\rangle\in(Q_{\mathcal{A}})^{*}\cup(Q_{\mathcal{A}})^{\omega}$ of the same length, let $r_{\mathcal{M}}\oplus r_{\mathcal{A}}=\langle\langle s_{0},\bot\rangle,\langle s_{1},q_{0}\rangle,\langle s_{2},q_{1}\rangle,\ldots\rangle$.
-Notice that for all $r_{\mathcal{M}}\oplus r_{\mathcal{A}}\in\Pi({\mathcal{M}}\times{\mathcal{A}})\cup\Omega({\mathcal{M}}\times{\mathcal{A}})$, we have $\pi_{\mathcal{A}}(r_{\mathcal{M}}\oplus r_{\mathcal{A}})=r_{\mathcal{A}}$ and $\pi_{\mathcal{M}}(r_{\mathcal{M}}\oplus r_{\mathcal{A}})=r_{\mathcal{M}}$.
-The semantic satisfaction probability of a strategy $\mu\in{\mathcal{F}}({\mathcal{M}})$ with respect to ${\mathcal{A}}$, denoted $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)$, is defined as:
-$\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)={\mathrm{Pr}}_{\mathcal{M}}^{\mu}(r\in\Omega({\mathcal{M}},\mu)|\tau(r)\in L({\mathcal{A}})).$
-That is, $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)$ is the probability that when the player plays with $\mu$, the word $\tau(r)$ generated by ${\mathcal{M}}$ is in $L({\mathcal{A}})$.
-Then, the semantic satisfaction probability of an MDP ${\mathcal{M}}$ with respect to ${\mathcal{A}}$, denoted $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$, is defined by
-$\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}=\sup_{\mu\in{\mathcal{F}}({\mathcal{M}})}\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu).$
-That is, $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$ is the least upper bound on how good the player can do when playing with ${\mathcal{M}}$ and ${\mathcal{A}}$.
-It is well known that for every nondeterministic automaton ${\mathcal{A}}$ with an $\omega$-regular acceptance condition, the suprimum is attained by a pure finite-memory strategy[11].
-Similarly, the syntactic satisfaction probability of $\mu\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})$ and ${\mathcal{M}}$, denoted $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu)$ and $\mathrm{PSyn}^{\mathcal{A}}_{{\mathcal{M}}}$ respectively, are defined by
-$\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu)={\mathrm{Pr}}_{{\mathcal{M}}\times{\mathcal{A}}}^{\mu}(r\in\Omega({\mathcal{M}}\times{\mathcal{A}},\mu)|\pi_{\mathcal{A}}(r)\mbox{ is accepting}),$
+Assume that, simultaneously, the player also generates a run of ${\mathcal{N}}$.
+This process can be modeled by an MDP that is a product of ${\mathcal{M}}$ with ${\mathcal{N}}$, and in each step the player takes an action of ${\mathcal{M}}$ and chooses a successor state in ${\mathcal{N}}$ that extends the current run on the new letter drawn by ${\mathcal{M}}$.
+Formally, we define the MDP ×${\mathcal{M}}\times{\mathcal{N}}=\left\langle S^{\prime},\left\langle s_{0},\bot\right\rangle,A^{\prime},\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{N}}},\Sigma,\gamma\right\rangle$ that is defined as follows.
+The set of states of ×${\mathcal{M}}\times{\mathcal{N}}$ is $S^{\prime}=(S\times Q)\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}$, for $\bot\notin Q$.
+The set of actions $A^{\prime}=A\left\langle s_{0},\bot\right\rangle\cup\bigcup_{\left\langle s,q\right\rangle\in S\times Q}(A_{\left\langle s,q\right\rangle})$ is defined as follows.
+For all $\left\langle s,q\right\rangle\in S\times Q$, the set of actions $A_{\left\langle s,q\right\rangle}$ available from $\left\langle s,q\right\rangle$ consists of all pairs $\left\langle a,q^{\prime}\right\rangle\in A_{s}\times\delta(q,\tau(s))$.
+In addition, the actions available from $\left\langle s_{0},\bot\right\rangle$ are $A_{\left\langle s_{0},\bot\right\rangle}=A_{s_{0}}\times\left\{q_{0}\right\}$.
+Then, for all $\left\langle s,\ast\right\rangle\in S^{\prime}$ and $\left\langle a,q^{\prime}\right\rangle\in A_{\left\langle s,\ast\right\rangle}$, we have $\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{N}}}(\left\langle s,\ast\right\rangle,\left\langle a,q^{\prime}\right\rangle,\left\langle s^{\prime},q^{\prime}\right\rangle)=\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$ and $\gamma(s,\ast)=\tau(s)$.
+Given a finite or infinite run $r\in\Pi({\mathcal{M}}\times{\mathcal{N}})\cup\Omega({\mathcal{M}}\times{\mathcal{N}})$, $r=\left\langle\left\langle s_{0},\bot\right\rangle,\left\langle s_{1},q_{0}\right\rangle,\left\langle s_{2},q_{1}\right\rangle,\ldots\right\rangle$, we define $\pi_{\mathcal{N}}(r)=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle$, and $\pi_{\mathcal{M}}(r)=\left\langle s_{0},s_{1},s_{2},\ldots\right\rangle$ to be the ${\mathcal{N}}$-run and ${\mathcal{M}}$-run that correspond to $r$ respectively.
+Conversely, for $r_{\mathcal{M}}=\left\langle s_{0},s_{1},s_{2},\ldots\right\rangle\in S^{*}\cup S^{\omega}$ and $r_{\mathcal{N}}=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle\in(Q_{\mathcal{N}})^{*}\cup(Q_{\mathcal{N}})^{\omega}$ of appropriate lengths, let $r_{\mathcal{M}}\oplus r_{\mathcal{N}}=\left\langle\left\langle s_{0},\bot\right\rangle,\left\langle s_{1},q_{0}\right\rangle,\left\langle s_{2},q_{1}\right\rangle,\ldots\right\rangle$.
+Notice that for all $r_{\mathcal{M}}\oplus r_{\mathcal{N}}\in\Pi({\mathcal{M}}\times{\mathcal{N}})\cup\Omega({\mathcal{M}}\times{\mathcal{N}})$, we have $\pi_{\mathcal{N}}(r_{\mathcal{M}}\oplus r_{\mathcal{N}})=r_{\mathcal{N}}$ and $\pi_{\mathcal{M}}(r_{\mathcal{M}}\oplus r_{\mathcal{N}})=r_{\mathcal{M}}$.
+We can now formally define the notion an automaton being good-for-MDP.
+We use the formalism of [HPS+20].
+The semantic satisfaction probability of a strategy $\mu\in{\mathcal{F}}({\mathcal{M}})$ with respect to ${\mathcal{N}}$, denoted $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)$, is defined as:
+$\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)={\mathrm{Pr}}_{\mathcal{M}}^{\mu}(r\in\Omega({\mathcal{M}},\mu):\tau(r)\in L({\mathcal{N}})).$
+That is, $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)$ is the probability that when the player plays with $\mu$, the word $\tau(r)$ generated by ${\mathcal{M}}$ is in $L({\mathcal{N}})$.
+Then, the semantic satisfaction probability of an MDP ${\mathcal{M}}$ with respect to ${\mathcal{N}}$, denoted $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$, is defined by
+$\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}=\sup_{\mu\in{\mathcal{F}}({\mathcal{M}})}\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu).$
+That is, $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$ is the least upper bound on how good the player can do with respect to the objective of generating a word in $L({\mathcal{N}})$.
+Similarly, the syntactic satisfaction probability of $\mu\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ and ${\mathcal{M}}$, denoted $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu)$ and $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}$ respectively, are defined by
+$\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu)={\mathrm{Pr}}_{{\mathcal{M}}\times{\mathcal{N}}}^{\mu}(r\in\Omega({\mathcal{M}}\times{\mathcal{N}},\mu):\pi_{\mathcal{N}}(r)\mbox{ is accepting}),$
 and
-$\mathrm{PSyn}^{\mathcal{A}}_{{\mathcal{M}}}=\sup_{\mu\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})}\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu).$
-As with the semantic case, the value $\mathrm{PSyn}^{\mathcal{A}}_{{\mathcal{M}}}$ is attained. Moreover, it is guaranteed to be attained by a pure memoryless strategy [11].
-In general, $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}\leq\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$, but equality is not guaranteed, as the optimal resolution of nondeterministic choices may require access to future events.
-An automaton ${\mathcal{A}}$ is Good-For-MDPs (GFM, for short) if for every MDP ${\mathcal{M}}$, it holds that $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}=\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}$.
+$\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}=\sup_{\mu\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})}\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu).$
+Maximizing the syntactic satisfaction can be expressed as a simple stochastic parity game that is played on top of the MDP ×${\mathcal{M}}\times{\mathcal{N}}$, where the winning objective is inherited from ${\mathcal{N}}$.
+In  [CH03], the authors prove that simple stochastic parity games admit optimal pure memoryless strategies.
+Thus, the supremum in the definition of $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}$ is attained by a pure memoryless strategy.
+Note that $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}$ is evaluated with respect to the MDP ×${\mathcal{M}}\times{\mathcal{N}}$, and thus by pure memoryless we mean that there is a strategy $\mu^{*}_{\mathcal{N}}:S\times Q\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}\rightarrow A\times Q$ such that $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}(\mu^{*}_{\mathcal{N}})=\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}$.
+Similarly, the semantic satisfaction can be expressed as a simple stochastic parity game that augments ×${\mathcal{M}}\times{\mathcal{N}}$ with a deterministic automaton ${\mathcal{D}}$ for $L({\mathcal{N}})$.
+Thus, the supremum in the definition of $\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}$ is attained by a pure finite-memory strategy.
+Indeed, the considerations are as in the syntactic case, except that here, the current state of the deterministic automaton ${\mathcal{D}}$ is used as a memory for the player.
+In general, $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}\leq\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$.
+In the syntactic case, the player not only aims to generate a word in $L({\mathcal{N}})$, but also to simultaneously generate an accepting run of ${\mathcal{N}}$ on it.
+Therefore, any strategy for the syntactic case can be used for the semantic case by ignoring the component of ${\mathcal{N}}$, resulting in a potentially higher semantic satisfaction.
+Equality, however, is not guaranteed.
+Indeed, the same way not all automata are history deterministic, it may not be possible to resolve the nondeterminism in ${\mathcal{N}}$ in a way that maximizes the probability of generating a word in $L({\mathcal{N}})$ (thus reaching a maximal semantic satisfaction) while generating rejecting runs on words in $L({\mathcal{N}})$ with zero probability (and thus reaching an equal syntactic satisfaction).
+When the equality holds for all MDP ${\mathcal{M}}$, we say that ${\mathcal{N}}$ is good-for-MDP (GFM for short).
+Note that when ${\mathcal{N}}$ is GFM, it holds that for all MDP ${\mathcal{M}}=\left\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{{\mathcal{M}}},\Sigma,\tau\right\rangle$, there exists a pure-memroyless strategy for ×${\mathcal{M}}\times{\mathcal{N}}$, $\mu^{*}_{\mathcal{N}}:S\times Q\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}\rightarrow A\times Q$, such that $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}=\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}(\mu^{*}_{\mathcal{N}})$ and $\mu^{*}_{\mathcal{N}}$ generates a rejecting run on words in $L({\mathcal{N}})$ with zero probability.
+Note that $\mu^{*}_{\mathcal{N}}$ can be viewed as a pure finite-memory strategy for ${\mathcal{M}}$ for which $\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}=\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}(\mu^{*}_{\mathcal{N}})$.
 6.2. The nondeterminism hierarchy and GFMness
 In this section we locate GFMness in the nondeterminism hierarchy.
-For the semantic hierarchy, it is proved in [13] that GFM-NBWs are as expressive as NBWs.
+For the semantic hierarchy, it is proved in [HPS+20] that GFM-NBWs are as expressive as NBWs.
 More specifically, every NBW can be translated into an equivalent GFM-NBW.
 Unlike the case of co-Büchi automata, where nondeterministic and deterministic automata have the same expressive power, making the expressiveness equivalence of GFM-NCWs and NCW straightforward, NBWs are more expressive than DBWs, making the result about GFM-NBWs interesting.
-Another aspect that has been studied is deciding GFMness [30].
-In  [30], the authors prove an exponential-time upper bound as well as a polynomial-space lower bound for the problem of deciding whether an automaton is GFM-NBW.
-We study the syntactic hierarchy, we first point out that every HD automaton is GFM.
+Another aspect that has been studied is deciding GFMness [STZ22].
+In  [STZ22], the authors prove an exponential-time upper bound as well as a polynomial-space lower bound for the problem of deciding whether an automaton is GFM-NBW.
+We study the syntactic hierarchy, we first point out that every HD automaton is GFM [KMBK14].
 Then, by relating GFMness and almost-DBPness, we show that SD-NBWs are GFM, yet SD-NCWs need not be GFM.
@@
 Theorem 14.
-Every HD automaton is GFM.
+Every HD automaton is GFM [KMBK14].
 Proof 6.1.
-Consider an HD automaton ${\mathcal{A}}$.
-Let $f$ be an HD strategy for ${\mathcal{A}}$.
-Then, every strategy $\mu\in{\mathcal{F}}({\mathcal{M}})$ can be augmented with $f$ to obtain a strategy $\mu\times f\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})$ such that $\mathrm{PSem}^{\mathcal{A}}_{{\mathcal{M}}}(\mu)=\mathrm{PSem}^{\mathcal{A}}_{{\mathcal{M}}\times{\mathcal{A}}}(\mu\times f)=\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu\times f)$.
-Thus, $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}\leq\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}$, and so ${\mathcal{A}}$ is GFM.
+The theorem is proved in [KMBK14].
+We give a short proof for completeness.
+Consider an HD automaton ${\mathcal{N}}$.
+Let $f$ be an HD strategy for ${\mathcal{N}}$.
+Then, every strategy $\mu\in{\mathcal{F}}({\mathcal{M}})$ can be augmented with $f$ to obtain a strategy $\mu\times f\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ such that $\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}(\mu)=\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}\times{\mathcal{N}}}(\mu\times f)=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu\times f)$.
+Thus, $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}\leq\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}$, and so ${\mathcal{N}}$ is GFM.
 Theorem 15.
@@
 Proof 6.2.
-Consider an $\omega$-regular automaton ${\mathcal{A}}$ over an alphabet $\Sigma$, and consider the MDP ${\mathcal{M}}=\langle\Sigma,\sigma_{0},(A_{\sigma})_{\sigma\in\Sigma},\mathrm{Pr}_{{\mathcal{M}}},\mathsf{id}\rangle$, where $\sigma_{0}\in\Sigma$ is an arbitrary initial state, $A_{\sigma}=\{\ast\}$ is a dummy set of actions for each letter $\sigma\in\Sigma$, and $\mathrm{Pr}_{{\mathcal{M}}}(\sigma,\ast,\sigma^{\prime})=|\Sigma|^{-1}$, for all $\sigma,\sigma^{\prime}\in\Sigma$.
-By [11], there is a pure memoryless strategy $\mu^{*}:S\times Q_{\mathcal{A}}\cup\{\langle s_{0},\bot\rangle\}\rightarrow A\times Q_{\mathcal{A}}$ that attains $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}$.
+Consider an $\omega$-regular automaton ${\mathcal{N}}$ over an alphabet $\Sigma$, and consider the MDP ${\mathcal{M}}=\left\langle\Sigma,\sigma_{0},(A_{\sigma})_{\sigma\in\Sigma},\mathrm{Pr}_{{\mathcal{M}}},\mathsf{id}\right\rangle$, where $\sigma_{0}\in\Sigma$ is an arbitrary initial state, $A_{\sigma}=\left\{\ast\right\}$ is a dummy set of actions for each letter $\sigma\in\Sigma$, and $\mathrm{Pr}_{{\mathcal{M}}}(\sigma,\ast,\sigma^{\prime})=|\Sigma|^{-1}$, for all $\sigma,\sigma^{\prime}\in\Sigma$.
+By [CJH03], there is a pure memoryless strategy $\mu^{*}:S\times Q_{\mathcal{N}}\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}\rightarrow A\times Q_{\mathcal{N}}$ that attains $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}$.
 Observe that since $\mathrm{Pr}_{{\mathcal{M}}}(\sigma,\ast,\cdot)=(\sigma^{\prime}\mapsto|\Sigma|^{-1})$ is the uniform distribution on $\Sigma$ for all $\sigma\in\Sigma$, then ${\mathcal{M}}$ is essentially an MC that generates a random word $w=\sigma_{1},\sigma_{2},\sigma_{3},\ldots\in\Sigma^{\omega}$, such that all its letters are independent and distributed uniformly.
-Thus, the membership $w\in L({\mathcal{A}})$, where $w$ is the word generated by ${\mathcal{C}}({\mathcal{M}},\mu^{\prime})$ (or by ${\mathcal{C}}({\mathcal{M}}\times{\mathcal{A}},\mu^{\prime})$) for $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}})$ (respectively $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})$), is independent of the strategy $\mu^{\prime}$.
-In particular $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$ equals the probability that a random word $w$ is in $L({\mathcal{A}})$.
-Note that it must be that $\mu^{*}(\sigma_{0},\bot)=\langle\ast,q_{0}\rangle$; i.e., there is no choice at the initial state of ×${\mathcal{M}}\times{\mathcal{A}}$.
-Also, as $A_{\sigma}=\{\ast\}$, for all $\sigma\in\Sigma$, we can think of $\mu^{*}$ as a function $\mu^{*}:\Sigma\times Q_{\mathcal{A}}\rightarrow Q_{\mathcal{A}}$, with $\mu^{*}(\sigma,q)\in\delta_{\mathcal{A}}(q,\sigma)$ for all $\langle\sigma,q\rangle\in\Sigma\times Q_{\mathcal{A}}$.
-Thus, $\mu^{*}$ induces a pruning ${\mathcal{A}}_{\mu^{*}}$ of ${\mathcal{A}}$, and $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu^{*})$ equals to the probability that a random word is accepted by ${\mathcal{A}}_{\mu^{*}}$.
-Now, if ${\mathcal{A}}$ is GFM, then $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu^{*})=\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}=\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$.
-Hence, the probability that a random word $w\in\Sigma^{\omega}$ is accepted by ${\mathcal{A}}_{\mu^{*}}$ equals the probability that $w\in L({\mathcal{A}})$.
-That is, ${\mathcal{A}}_{\mu^{*}}$ is a deterministic pruning of ${\mathcal{A}}$ that loses only a set of words of measure zero.
-Thus, ${\mathcal{A}}$ is almost-DBP, and we are done.
+Thus, the membership $w\in L({\mathcal{N}})$, where $w$ is the word generated by ${\mathcal{C}}({\mathcal{M}},\mu^{\prime})$ (or by ${\mathcal{C}}({\mathcal{M}}\times{\mathcal{N}},\mu^{\prime})$) for $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}})$ (respectively $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$), is independent of the strategy $\mu^{\prime}$.
+In particular $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$ equals the probability that a random word $w$ is in $L({\mathcal{N}})$.
+Note that it must be that $\mu^{*}(\sigma_{0},\bot)=\left\langle\ast,q_{0}\right\rangle$; i.e., there is no choice at the initial state of ×${\mathcal{M}}\times{\mathcal{N}}$.
+Also, as $A_{\sigma}=\left\{\ast\right\}$, for all $\sigma\in\Sigma$, we can think of $\mu^{*}$ as a function $\mu^{*}:\Sigma\times Q_{\mathcal{N}}\rightarrow Q_{\mathcal{N}}$, with $\mu^{*}(\sigma,q)\in\delta_{\mathcal{N}}(q,\sigma)$ for all $\left\langle\sigma,q\right\rangle\in\Sigma\times Q_{\mathcal{N}}$.
+Thus, $\mu^{*}$ induces a pruning ${\mathcal{N}}_{\mu^{*}}$ of ${\mathcal{N}}$, and $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{*})$ equals to the probability that a random word is accepted by ${\mathcal{N}}_{\mu^{*}}$.
+Now, since $\mu^{*}$ is an optimal strategy for ×${\mathcal{M}}\times{\mathcal{N}}$ and ${\mathcal{N}}$ is GFM, it follows that $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{*})=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$.
+Hence, the probability that a random word $w\in\Sigma^{\omega}$ is accepted by ${\mathcal{N}}_{\mu^{*}}$ equals the probability that $w\in L({\mathcal{N}})$.
+That is, ${\mathcal{N}}_{\mu^{*}}$ is a deterministic pruning of ${\mathcal{N}}$ that loses only a set of words of measure zero and hence ${\mathcal{N}}$ is almost-DBP.
 Recall that in Theorem 11, we argued that there is an SD-NCW that is not almost-DBP.
 By Theorem 15, we can conclude that the SD-NCW described there, is also not GFM.
-Thus, while HDness implies GFMness, for co-Büchi automata, SDness need not imply GFMness.
+Thus, while HDness implies GFMness for all accaptence conditions [KMBK14], for co-Büchi automata, SDness need not imply GFMness.
 In the rest of this section we prove that for Büchi automata, SDness does imply GFMness, thus all SD-NBWs are GFM.
@@
 Proposition 16.
-If ${\mathcal{A}}$ is an SD-NBW and $\mu\in{\mathcal{F}}({\mathcal{M}})$ is a finite memory strategy, then there exists $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})$ such that $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)=\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu^{\prime})$.
+If ${\mathcal{N}}$ is an SD-NBW and $\mu\in{\mathcal{F}}({\mathcal{M}})$ is a finite memory strategy, then there exists $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ such that $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})$.
 Proof 6.3.
-Let $\mu:M\times S\rightarrow\mathrm{dist}(A)\times M$, $\mu\in{\mathcal{F}}({\mathcal{M}})$, be a finite memory strategy, and let ${\mathcal{C}}={\mathcal{C}}({\mathcal{M}},\mu)$ be the Markov chain where actions of ${\mathcal{M}}$ are resolved according to $\mu$.
-Notice that since $\mu$ uses $M$ as a memory structure, we may think of ${\mathcal{C}}$ as a Markov chain with positions $M\times S$, and a stochastic transition function $\mathrm{Pr}_{\mathcal{C}}:(M\times S)^{2}\rightarrow[0,1]$ that is defined $\mathrm{Pr}_{\mathcal{C}}(\langle m,s\rangle,\langle m^{\prime},s^{\prime}\rangle)=\sum_{a\in A_{s}}d(a)\cdot\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$ when $\mu(m,s)=\langle d,m^{\prime}\rangle$, and otherwise $\mathrm{Pr}_{\mathcal{C}}(\langle m,s\rangle,\langle m^{\prime},s^{\prime}\rangle)=0$.
-Let $r=\langle\langle m_{0},s_{0}\rangle,\langle m_{1},s_{1}\rangle,\langle m_{2},s_{2}\rangle,\ldots\rangle\in(M\times S)^{\omega}$ be a random process sampled according to ${\mathcal{C}}$.
-Let $r_{\mathcal{A}}=\langle q_{0},q_{1},q_{2},\ldots\rangle\in(Q_{\mathcal{A}})^{\omega}$ be a random run of ${\mathcal{A}}$ over the word $\tau(r)$ that is sampled by playing according to $\mu$ in ${\mathcal{M}}$.
-I.e., given $\langle m_{0},s_{0}\rangle,\langle m_{1},s_{1}\rangle,\ldots,\langle m_{k},s_{k}\rangle$, for $k\geq 1$, and $q_{0},q_{1},\ldots,q_{k-1}$, we set $q_{k}$ to be one of the $\tau(s_{k})$-successors of $q_{k-1}$ at random.
-Notice that the composition $r\oplus r_{\mathcal{A}}=\langle\langle m_{0},s_{0},q_{0}\rangle,\langle m_{1},s_{1},q_{1}\rangle,\ldots\rangle$ is a process of a finite Markov chain ${\mathcal{C}}_{\mathcal{A}}$ with a set of positions $M\times S\times Q_{\mathcal{A}}$.
-Indeed, the stochastic transition function of ${\mathcal{C}}_{\mathcal{A}}$ can be expressed as follows:
-[…]
-Since ${\mathcal{C}}_{\mathcal{A}}$ is a finite Markov chain, it holds that with probability 1 the random process $r\oplus r_{\mathcal{A}}$ of ${\mathcal{C}}_{\mathcal{A}}$ eventually enters an ergodic component and traverses all its states.
-We say that an ergodic component $C\subseteq M\times S\times Q_{\mathcal{A}}$ of ${\mathcal{C}}_{\mathcal{A}}$ is $\alpha_{\mathcal{A}}$-accepting if the projection of $C$ onto $Q_{\mathcal{A}}$ intersects $\alpha_{\mathcal{A}}$.
-I.e., $C$ is $\alpha_{\mathcal{A}}$-accepting if there exist $s\in S$, $m\in M$ and $q\in\alpha_{\mathcal{A}}$ such that $\langle m,s,q\rangle\in C$.
-Thus, if $r\oplus r_{\mathcal{A}}$ eventually enters an $\alpha_{\mathcal{A}}$-accepting ergodic component $C$, then with probability 1 the process traverses all the positions in $C$ infinitely often, and in particular with probability 1 $r_{\mathcal{A}}$ is an accepting run of ${\mathcal{A}}$ over the generated word $\tau(r)$.
-Thus, since with probability 1 the process of ${\mathcal{C}}_{\mathcal{A}}$ eventually enters an ergodic component, it follows that the probability that $\tau(r)\in L({\mathcal{A}})$ but $r_{\mathcal{A}}$ is rejecting equals the probability that $r\oplus r_{\mathcal{A}}$ enters an ergodic component that is not $\alpha_{\mathcal{A}}$-accepting and $\tau(r)\in L({\mathcal{A}})$.
-We prove that if $r\oplus r_{\mathcal{A}}$ enters an ergodic component $C$ that is not $\alpha_{\mathcal{A}}$-accepting then $\tau(r)\notin L({\mathcal{A}})$.
-Note that we claim that the implication $C$ not $\alpha_{\mathcal{A}}$-accepting implies that $\tau(r)\notin L({\mathcal{A}})$, is surely, and not only almost-surely.
-Consider an ergodic component $C$ that is not $\alpha_{\mathcal{A}}$-accepting, and assume that $r\oplus r_{\mathcal{A}}$ is such that $\langle m_{k},s_{k},q_{k}\rangle\in C$ for some $k\geq 0$.
-Assume by way of contradiction that $\tau(r)\in L({\mathcal{A}})$.
-Since ${\mathcal{A}}$ is SD, it is possible to extend the finite run $q_{0},q_{1},\ldots q_{k}$ of ${\mathcal{A}}$ over $\tau(s_{1}),\tau(s_{2}),\ldots,\tau(s_{k})$ into an accepting run of ${\mathcal{A}}$ over $\tau(r)$.
-In particular there exist $j>k$ and $q^{\prime}_{k+1},q^{\prime}_{k+2},\ldots,q^{\prime}_{j}\in(Q_{\mathcal{A}})^{*}$, such that $q^{\prime}_{j}\in\alpha_{\mathcal{A}}$ and $q_{0},\ldots,q_{k},q^{\prime}_{k+1},\ldots,q^{\prime}_{j}$ is a finite run of ${\mathcal{A}}$ over $\tau(s_{1}),\ldots,\tau(s_{j})$.
-By definition of the stochastic transition function of ${\mathcal{C}}_{\mathcal{A}}$, it follows that there is a positive probability to move to $\langle m_{k+1},s_{k+1},q^{\prime}_{k+1}\rangle$ from $\langle m_{k},s_{k},q_{k}\rangle$.
-Similarly, it follows by induction that for all $k+1<i\leq j$ there is a positive probability to move to $\langle m_{i},s_{i},q^{\prime}_{i}\rangle$ from $\langle m_{i-1},s_{i-1},q^{\prime}_{i-1}\rangle$.
-In particular $\langle m_{j},s_{j},q^{\prime}_{j}\rangle\in C$, in contradiction to the assumption that $C$ is not $\alpha_{\mathcal{A}}$-accepting.
-Let $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})$ be the strategy for ×${\mathcal{M}}\times{\mathcal{A}}$ that resolves the actions of ${\mathcal{M}}$ according to $\mu$ and in addition generates a random run of ${\mathcal{A}}$ over $\tau(r)$.
-It is not hard to see that ${\mathcal{C}}_{\mathcal{A}}$ induces the same distribution over infinite paths in $(S\times Q_{\mathcal{A}})^{\omega}$ as the Markov chain ${\mathcal{C}}({\mathcal{M}}\times{\mathcal{A}},\mu^{\prime})$.
-We proved that there is a zero probability that $\tau(r)\in L({\mathcal{A}})$ and $r_{\mathcal{A}}$ is rejecting.
-Thus, $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu^{\prime})=\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)$.
-We can now conclude with the desired connection between SDness and GFMness.
-We first examine whether SDness implies GFMness, and show that the answer depends on the acceptance condition.
+Let $\mu:S\times M\rightarrow\mathrm{dist}(A)\times M$, be a finite memory strategy in ${\mathcal{F}}({\mathcal{M}})$, and let ${\mathcal{C}}={\mathcal{C}}({\mathcal{M}},\mu)$ be the Markov chain induced when taking actions in ${\mathcal{M}}$ according to $\mu$.
+Note that since $\mu$ uses $M$ as a memory structure, we may think of ${\mathcal{C}}$ as a Markov chain with positions $S\times M$, and a stochastic transition function $\mathrm{Pr}_{\mathcal{C}}:(S\times M)^{2}\rightarrow[0,1]$ that is defined by $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=\sum_{a\in A_{s}}d(a)\cdot\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$, when $\mu(s,m)=\left\langle d,m^{\prime}\right\rangle$, and $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=0$, otherwise.
+Let $r=\left\langle\left\langle s_{0},m_{0}\right\rangle,\left\langle s_{1},m_{1}\right\rangle,\left\langle s_{2},m_{2}\right\rangle,\ldots\right\rangle\in(S\times M)^{\omega}$ be a random process sampled according to ${\mathcal{C}}$.
+Let $r_{\mathcal{N}}=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle\in(Q_{\mathcal{N}})^{\omega}$ be a random run of ${\mathcal{N}}$ over the word $\tau(r)$ that is sampled when playing in ${\mathcal{M}}$ with the startegy $\mu$.
+I.e., given $\left\langle s_{0},m_{0}\right\rangle,\left\langle s_{1},m_{1}\right\rangle,\ldots,\left\langle s_{k},m_{k}\right\rangle$, for $k\geq 1$, and $q_{0},q_{1},\ldots,q_{k-1}$, we set $q_{k}$ to be one of the $\tau(s_{k})$-successors of $q_{k-1}$ at random.
+Notice that the composition $r\oplus r_{\mathcal{N}}=\left\langle\left\langle s_{0},m_{0},q_{0}\right\rangle,\left\langle s_{1},m_{1},q_{1}\right\rangle,\ldots\right\rangle$ is a process of a finite Markov chain ${\mathcal{C}}_{\mathcal{N}}$ with a set of positions $S\times M\times Q_{\mathcal{N}}$.
+Indeed, the stochastic transition function of ${\mathcal{C}}_{\mathcal{N}}$ can be expressed as follows:
+$\mathrm{Pr}_{{\mathcal{C}}_{\mathcal{N}}}(\left\langle s,m,q\right\rangle,\left\langle s^{\prime},m^{\prime},q^{\prime}\right\rangle)=\begin{cases}|\delta_{\mathcal{N}}(q,\tau(s^{\prime}))|^{-1}\cdot\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)&:q^{\prime}\in\delta_{\mathcal{N}}(q,\tau(s^{\prime}))\\ 0&:\mbox{otherwise}\end{cases}$
+Since ${\mathcal{C}}_{\mathcal{N}}$ is a finite Markov chain, it holds that with probability $1$, the random process $r\oplus r_{\mathcal{N}}$ of ${\mathcal{C}}_{\mathcal{N}}$ eventually enters an ergodic component and traverses all its states.
+We say that an ergodic component $C\subseteq S\times M\times Q_{\mathcal{N}}$ of ${\mathcal{C}}_{\mathcal{N}}$ is $\alpha_{\mathcal{N}}$-accepting if the projection of $C$ onto $Q_{\mathcal{N}}$ intersects $\alpha_{\mathcal{N}}$.
+I.e., $C$ is $\alpha_{\mathcal{N}}$-accepting if there exist $s\in S$, $m\in M$ and $q\in\alpha_{\mathcal{N}}$ such that $\left\langle s,m,q\right\rangle\in C$.
+Thus, if $r\oplus r_{\mathcal{N}}$ eventually enters an $\alpha_{\mathcal{N}}$-accepting ergodic component $C$, then with probability 1 the process traverses all the positions in $C$ infinitely often, and in particular with probability 1 $r_{\mathcal{N}}$ is an accepting run of ${\mathcal{N}}$ over the generated word $\tau(r)$.
+Since with probability 1 the process of ${\mathcal{C}}_{\mathcal{N}}$ eventually enters an ergodic component, it follows that the probability that $\tau(r)\in L({\mathcal{N}})$ but $r_{\mathcal{N}}$ is rejecting equals the probability that $r\oplus r_{\mathcal{N}}$ enters an ergodic component that is not $\alpha_{\mathcal{N}}$-accepting and $\tau(r)\in L({\mathcal{N}})$.
+We prove that if $r\oplus r_{\mathcal{N}}$ enters an ergodic component $C$ that is not $\alpha_{\mathcal{N}}$-accepting then $\tau(r)\notin L({\mathcal{N}})$.
+Note that we claim that the implication $C$ not $\alpha_{\mathcal{N}}$-accepting implies that $\tau(r)\notin L({\mathcal{N}})$, is surely, and not only almost-surely.
+Consider an ergodic component $C$ that is not $\alpha_{\mathcal{N}}$-accepting, and assume that $r\oplus r_{\mathcal{N}}$ is such that $\left\langle s_{k},m_{k},q_{k}\right\rangle\in C$ for some $k\geq 0$.
+Assume by way of contradiction that $\tau(r)\in L({\mathcal{N}})$.
+Since ${\mathcal{N}}$ is SD, it is possible to extend the finite run $q_{0},q_{1},\ldots q_{k}$ of ${\mathcal{N}}$ over $\tau(s_{1}),\tau(s_{2}),\ldots,\tau(s_{k})$ into an accepting run of ${\mathcal{N}}$ over $\tau(r)$.
+In particular there exist $j>k$ and $q^{\prime}_{k+1},q^{\prime}_{k+2},\ldots,q^{\prime}_{j}\in(Q_{\mathcal{N}})^{*}$, such that $q^{\prime}_{j}\in\alpha_{\mathcal{N}}$ and $q_{0},\ldots,q_{k},q^{\prime}_{k+1},\ldots,q^{\prime}_{j}$ is a finite run of ${\mathcal{N}}$ over $\tau(s_{1}),\ldots,\tau(s_{j})$.
+By definition of the stochastic transition function of ${\mathcal{C}}_{\mathcal{N}}$, it follows that there is a positive probability to move to $\left\langle s_{k+1},m_{k+1},q^{\prime}_{k+1}\right\rangle$ from $\left\langle s_{k},m_{k},q_{k}\right\rangle$.
+Similarly, it follows by induction that for all $k+1<i\leq j$ there is a positive probability to move to $\left\langle s_{i},m_{i},q^{\prime}_{i}\right\rangle$ from $\left\langle s_{i-1},m_{i-1},q^{\prime}_{i-1}\right\rangle$.
+In particular $\left\langle s_{j},m_{j},q^{\prime}_{j}\right\rangle\in C$, in contradiction to the assumption that $C$ is not $\alpha_{\mathcal{N}}$-accepting.
+Let $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ be the strategy for ×${\mathcal{M}}\times{\mathcal{N}}$ that resolves the actions of ${\mathcal{M}}$ according to $\mu$ and in addition generates a random run of ${\mathcal{N}}$ over $\tau(r)$.
+It is not hard to see that ${\mathcal{C}}_{\mathcal{N}}$ induces the same distribution over infinite paths in $(S\times Q_{\mathcal{N}})^{\omega}$ as the Markov chain ${\mathcal{C}}({\mathcal{M}}\times{\mathcal{N}},\mu^{\prime})$.
+We proved that there is a zero probability that $\tau(r)\in L({\mathcal{N}})$ and $r_{\mathcal{N}}$ is rejecting.
+Thus, $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)$.
+We can now conclude with the desired connection between SDness and GFMness. We first examine whether SDness implies GFMness, and show that the answer depends on the acceptance condition.
 Theorem 17.
@@
 The co-Büchi case follows from Theorems 11 and 15.
-For Büchi automata, consider an SD-NBW ${\mathcal{A}}=\langle\Sigma,Q_{\mathcal{A}},q_{0},\delta_{\mathcal{A}},\alpha_{\mathcal{A}}\rangle$, and a $\Sigma$-labeled MDP ${\mathcal{M}}$.
-Let $\mu$ be a finite memory strategy such that $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)=\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$ given by [11].
-Then by Proposition 16 it holds that there exists $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})$ such that $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu^{\prime})=\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)$, and we get $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}=\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu^{\prime})\leq\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}$.
-Since $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}\leq\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$ holds for all ${\mathcal{A}}$ and ${\mathcal{M}}$, we have the equality $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}=\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$.
+For Büchi automata, consider an SD-NBW ${\mathcal{N}}=\left\langle\Sigma,Q_{\mathcal{N}},q_{0},\delta_{\mathcal{N}},\alpha_{\mathcal{N}}\right\rangle$, and a $\Sigma$-labeled MDP ${\mathcal{M}}$.
+Let $\mu$ be a finite memory strategy such that $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$ given by [CJH03].
+Then by Proposition 16 it holds that there exists $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ such that $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)$, and we get $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})\leq\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}$.
+Since $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}\leq\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$ holds for all ${\mathcal{N}}$ and ${\mathcal{M}}$, we have the equality $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$.
 We continue and check whether GFMness implies SDness, and show that the answer is negative even for weak automata.
@@
 […]
+We show that ${\mathcal{N}}$ is GFM.
+Consider an MDP ${\mathcal{M}}$, and let $\mu:S\times M\to dist(A)\times M$ be a finite memory strategy such that $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)=\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}$.
+Consider the MC ${\mathcal{C}}={\mathcal{C}}({\mathcal{M}},\mu)$, and note that since $\mu$ uses $M$ as a memory structure, we may think of ${\mathcal{C}}$ as a finite MC with positions $S\times M$, and a stochastic transition function $\mathrm{Pr}_{\mathcal{C}}:(S\times M)^{2}\rightarrow[0,1]$ that is defined $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=\sum_{a\in A_{s}}d(a)\cdot\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$ when $\mu(s,m)=\left\langle d,m^{\prime}\right\rangle$, and otherwise $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=0$.
+To conclude the GFMness of ${\mathcal{N}}$, we show that there is a strategy $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ such that $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})$.
+Recall that a strategy for ×${\mathcal{M}}\times{\mathcal{N}}$ needs to take actions in ${\mathcal{M}}$ and in ${\mathcal{N}}$ simultaneously.
+In the first it needs to choose an action $a\in A_{s}$ where $s\in S$ is the current state in ${\mathcal{M}}$, while in the latter it should choose a $\tau(s)$-succsessor of the current state $q$ in ${\mathcal{N}}$.
+The strategy $\mu^{\prime}$ resolves the first according to to $\mu$, and hence ignores the ${\mathcal{N}}$ component.
+Then, for the ${\mathcal{N}}$ component, $\mu^{\prime}$ generates a run $r_{\mathcal{N}}$ of ${\mathcal{N}}$ on $w_{\mathcal{C}}$ as follows.
+The run $r_{\mathcal{N}}$ waits at $q_{0}$ (possibly forever), and moves to $q_{acc}$ only when an ergodic SCC that does not contain an $a$-labeled state is reached in ${\mathcal{C}}$; in particular, if the ergodic SCC that is reached in ${\mathcal{C}}$ contains an $a$-labeled state, then $r_{\mathcal{N}}$ does not leave $q_{0}$.
+In other words, the strategy $\mu^{\prime}$ moves from $q_{0}$ to $q_{acc}$ only when it is certain that there are no more $a$’s in the suffix to be read.
+As the MC ${\mathcal{C}}$ is finite, then with probability 1 an ergodic component of ${\mathcal{C}}$ is reached and all of its states are traversed infinitely often.
+Hence, with probability 1, we have one of the following outcomes.
+Either the ergodic component that is reached does not contains an $a$-labeled state, or it contains an $a$-labeled state.
+In the former case, the run $r_{\mathcal{N}}$ is accepting.
+Indeed, according to $\mu^{\prime}$, the run $r_{\mathcal{N}}$ moves to $q_{acc}$ once the ergodic component that has no $a$-labeled states is reached, and thus $r_{\mathcal{N}}$ eventually gets stuck at $q_{acc}$ while reading a suffix of $w_{\mathcal{C}}$ that has no $a$’s.
+In the latter case, $r_{\mathcal{N}}$ is stuck in $q_{0}$ and is thus rejecting, and with probability 1 $w_{\mathcal{C}}$ has infinitely many $a$’s.
+Hence, in the latter case there is a zero probability that $w_{\mathcal{C}}\in L({\mathcal{N}})$ and $r_{\mathcal{N}}$ is rejecting, and we are done.
 Figure 6. A GFM-NWW that is not SD. Missing transitions lead to a rejecting sink.
@@
 Proposition 19.
-Consider states $q\in Q$ and $S\in Q^{\prime}$, a letter $\sigma\in\Sigma$, and transitions $\langle q,\sigma,q^{\prime}\rangle$ and $\langle S,\sigma,S^{\prime}\rangle$ of ${\mathcal{A}}$ and ${\mathcal{A}}^{\prime}$, respectively.
+Consider states $q\in Q$ and $S\in Q^{\prime}$, a letter $\sigma\in\Sigma$, and transitions $\left\langle q,\sigma,q^{\prime}\right\rangle$ and $\left\langle S,\sigma,S^{\prime}\right\rangle$ of ${\mathcal{A}}$ and ${\mathcal{A}}^{\prime}$, respectively.
 If $q$ is ${\mathcal{A}}$-equivalent to the states in $S$, then $q^{\prime}$ is ${\mathcal{A}}$-equivalent to the states in $S^{\prime}$.
@@
 Consider a state $S$ of ${\mathcal{A}}^{\prime}$ and a letter $\sigma\in\Sigma$.
-If $\langle S,\sigma,S^{\prime}\rangle\in\Delta^{\prime}$ and $S^{\prime}\notin\alpha^{\prime}$, then all the $\sigma$-successors of a state $s\in S$ are in $S^{\prime}\setminus\alpha$.
+If $\left\langle S,\sigma,S^{\prime}\right\rangle\in\Delta^{\prime}$ and $S^{\prime}\notin\alpha^{\prime}$, then all the $\sigma$-successors of a state $s\in S$ are in $S^{\prime}\setminus\alpha$.
 We can now prove the correctness of the construction:
@@
 We also define $j_{0}=0$.
-Consider the DAG $G=\langle V,E\rangle$, where
+Consider the DAG $G=\left\langle V,E\right\rangle$, where
   • $V\subseteq Q\times\mathbb{N}$ is the union $\bigcup_{i\geq 0}(S_{j_{i}}\times\{i\})$.
-  • $E\subseteq\bigcup_{i\geq 0}(S_{j_{i}}\times\{i\})\times(S_{j_{i+1}}\times\{i+1\})$ is such that for all $i\geq 0$, it holds that $E(\langle s^{\prime},i\rangle,\langle s,i+1\rangle)$ iff there is a finite run from $s^{\prime}$ to $s$ over $w[j_{i}+1,j_{i+1}]$.
+  • $E\subseteq\bigcup_{i\geq 0}(S_{j_{i}}\times\{i\})\times(S_{j_{i+1}}\times\{i+1\})$ is such that for all $i\geq 0$, it holds that $E(\left\langle s^{\prime},i\right\rangle,\left\langle s,i+1\right\rangle)$ iff there is a finite run from $s^{\prime}$ to $s$ over $w[j_{i}+1,j_{i+1}]$.
 Then, we label this edge by the run from $s^{\prime}$ to $s$.
-By the definition of ${\mathcal{A}}^{\prime}$, for every $j\geq 0$ and state $s_{j+1}\in S_{j+1}$, there is a state $s_{j}\in S_{j}$ such that $\langle s_{j},\sigma_{j},s_{j+1}\rangle\in\Delta$.
+By the definition of ${\mathcal{A}}^{\prime}$, for every $j\geq 0$ and state $s_{j+1}\in S_{j+1}$, there is a state $s_{j}\in S_{j}$ such that $\left\langle s_{j},\sigma_{j},s_{j+1}\right\rangle\in\Delta$.
 Thus, it follows by induction that for every $i\geq 0$ and state $s_{i+1}\in S_{j_{i+1}}$, there is a state $s_{i}\in S_{j_{i}}$ such that there is a finite run from $s_{i}$ to $s_{i+1}$ on $w[j_{i}+1,j_{i+1}]$.
-Thus, the DAG $G$ has infinitely many reachable vertices from the vertex $\langle q_{0},0\rangle$.
+Thus, the DAG $G$ has infinitely many reachable vertices from the vertex $\left\langle q_{0},0\right\rangle$.
 Also, as the nondeterminism degree of ${\mathcal{A}}$ is finite, so is the branching degree of $G$.
@@
 We describe the technical details of the construction of ${\mathcal{A}}$.
-Let $T=\langle\Gamma,Q,\rightarrow,q_{0},q_{acc},q_{rej}\rangle$, where $\Gamma$ is the working alphabet, $Q$ is the set of states, $\rightarrow\subseteq Q\times\Gamma\times Q\times\Gamma\times\{L,R\}$ is the transition relation (we use $(q,a)\rightarrow(q^{\prime},b,\Delta)$ to indicate that when $T$ is in state $q$ and it reads the input $a$ in the current tape cell, it moves to state $q^{\prime}$, writes $b$ in the current tape cell, and its reading head moves one cell to the left/right, according to $\Delta$), $q_{0}$ is the initial state, $q_{acc}$ is the accepting state, and $q_{rej}$ is the rejecting one.
+Let $T=\left\langle\Gamma,Q,\rightarrow,q_{0},q_{acc},q_{rej}\right\rangle$, where $\Gamma$ is the working alphabet, $Q$ is the set of states, $\rightarrow\subseteq Q\times\Gamma\times Q\times\Gamma\times\{L,R\}$ is the transition relation (we use $(q,a)\rightarrow(q^{\prime},b,\Delta)$ to indicate that when $T$ is in state $q$ and it reads the input $a$ in the current tape cell, it moves to state $q^{\prime}$, writes $b$ in the current tape cell, and its reading head moves one cell to the left/right, according to $\Delta$), $q_{0}$ is the initial state, $q_{acc}$ is the accepting state, and $q_{rej}$ is the rejecting one.
 The transitions function $\rightarrow$ is defined also for the final states $q_{acc}$ and $q_{rej}$: when a computation of $T$ reaches them, it erases the tape, goes to the leftmost cell in the tape, and moves to the initial state $q_{0}$.
@@
 We also set $\sigma_{0}$, $\sigma^{\prime}_{0}$, and $\sigma_{n_{0}+1}$ to ##\##.
-For each triple $\langle\sigma_{i-1},\sigma_{i},\sigma_{i+1}\rangle$ with $1\leq i\leq n_{0}$, we know, by the transition relation of $T$, what $\sigma^{\prime}_{i}$ should be.
+For each triple $\left\langle\sigma_{i-1},\sigma_{i},\sigma_{i+1}\right\rangle$ with $1\leq i\leq n_{0}$, we know, by the transition relation of $T$, what $\sigma^{\prime}_{i}$ should be.
 In addition, the letter ##\## should repeat exactly every $n_{0}+1$ letters.
@@
   • $next(\sigma_{i-1},\sigma_{i},\sigma_{i+1})=\sigma_{i}$ if $\sigma_{i}=\#$, or if non of $\sigma_{i-1},\sigma_{i}$ and $\sigma_{i+1}$ are in $Q\times\Gamma$.
-  • $next(\sigma_{i-1},\sigma_{i},\sigma_{i+1})=\#$ if |$|\{\sigma_{i-1},\sigma_{i},\sigma_{i+1}\}\cap Q\times\Gamma|\geq 2$, or if |$|\{\sigma_{i-1},\sigma_{i},\sigma_{i+1}\}\cap\{\#\}|\geq 2$  [^10] These case describes an illegal encoding of a configuration, and there is no correct expectation of the letter that comes in the same position in the next configuration.
+  • $next(\sigma_{i-1},\sigma_{i},\sigma_{i+1})=\#$ if |$|\left\{\sigma_{i-1},\sigma_{i},\sigma_{i+1}\right\}\cap Q\times\Gamma|\geq 2$, or if |$|\left\{\sigma_{i-1},\sigma_{i},\sigma_{i+1}\right\}\cap\left\{\#\right\}|\geq 2$  [^10] These case describes an illegal encoding of a configuration, and there is no correct expectation of the letter that comes in the same position in the next configuration.
 Thus, the choice $next(\sigma_{i-1},\sigma_{i},\sigma_{i+1})=\#$ is arbitrary..
@@
 In order to accept words that satisfy C1, namely detect a violation of the encoding, we distinguish between two types of violations: a violation of a single encoded configuration, and a violation of $next$.
-In order to detect a violation of $next$, the NWW ${\mathcal{A}}$ use its nondeterminism and guesses a triple $\langle\sigma_{i-1},\sigma_{i},\sigma_{i+1}\rangle\in\Sigma^{3}$ and guesses a position in the word, where it checks whether the three letters to be read starting this position are $\sigma_{i-1},\sigma_{i}$, and $\sigma_{i+1}$, and checks whether $next(\sigma_{i-1},\sigma_{i},\sigma_{i+1})$ is not the letter to come $n_{0}+1$ letters later.
+In order to detect a violation of $next$, the NWW ${\mathcal{A}}$ use its nondeterminism and guesses a triple $\left\langle\sigma_{i-1},\sigma_{i},\sigma_{i+1}\right\rangle\in\Sigma^{3}$ and guesses a position in the word, where it checks whether the three letters to be read starting this position are $\sigma_{i-1},\sigma_{i}$, and $\sigma_{i+1}$, and checks whether $next(\sigma_{i-1},\sigma_{i},\sigma_{i+1})$ is not the letter to come $n_{0}+1$ letters later.
 Once ${\mathcal{A}}$ sees such a violation, it goes to an accepting sink.
@@
 Thus, we may take $z$ to be the minimal prefix of $v$ that is in $R_{n,m}$.
-Recall that an HD strategy $g:\Sigma^{*}\rightarrow Q$ has to agree with the the transitions of ${\mathcal{A}}_{\varphi}$.
+Recall that an HD strategy $g:\Sigma^{*}\rightarrow Q$ has to agree with the transitions of ${\mathcal{A}}_{\varphi}$.
 That is, for all $w\in(X\cdot C)^{*}$, $x_{k}\in X$, and $c_{j}\in C$, it holds that $(g(w),x_{k},g(w\cdot x_{k}))$ and $(g(w\cdot x_{k}),c_{j},g(x\cdot x_{k}\cdot c_{j}))$ are in $\Delta_{\varphi}$.
@@
 It is easy to see that there is a correspondence between assignments to the variables in $X$ and deterministic prunnings of ${\mathcal{A}}_{\varphi}$.
-Indeed, a pruning of $p$ amounts to choosing, for each $k\in[n]$, a value $i_{k}\in{\{0,1\}}$: the assignment $x_{k}=i_{k}$ corresponds to keeping the transition $\langle p,x_{k},q^{i_{k}}_{k}\rangle$ and removing the transition $\langle p,x_{k},q^{\lnot i_{k}}_{k}\rangle$.
+Indeed, a pruning of $p$ amounts to choosing, for each $k\in[n]$, a value $i_{k}\in{\{0,1\}}$: the assignment $x_{k}=i_{k}$ corresponds to keeping the transition $\left\langle p,x_{k},q^{i_{k}}_{k}\right\rangle$ and removing the transition $\left\langle p,x_{k},q^{\lnot i_{k}}_{k}\right\rangle$.
 For an assignment $a:X\rightarrow{\{0,1\}}$, we denote by ${\mathcal{A}}_{\varphi}^{a}$ the deterministic pruning of ${\mathcal{A}}_{\varphi}$ that is associated with $a$.

### v5

@@
 Approximated determinization by pruning
-We say that a nondeterministic automaton ${\mathcal{A}}$ is almost-DBP if ${\mathcal{A}}$ embodies a deterministic automaton ${\mathcal{A}}^{\prime}$ such that the probability of a random word to be in $L({\mathcal{A}})\setminus L({\mathcal{A}}^{\prime})$ is 00.
+We say that a nondeterministic automaton $\mathcal{A}$} is almost-DBP if $\mathcal{A}$} embodies a deterministic automaton $\mathcal{A}$$^{\prime}$ such that the probability of a random word to be in $L({\mathcal{A}})\setminus L({\mathcal{A}}^{\prime})$ is 0.
 Clearly, if $\mathcal{A}$} is DBP, then it is almost-DBP.
@@
 Two states $q,s\in Q$ are equivalent, denoted $q\sim_{{\mathcal{A}}}s$, if $L({\mathcal{A}}^{q})=L({\mathcal{A}}^{s})$.
-[…]
+The acceptance condition $\alpha$ determines which runs are “good”.
+We consider here the Büchi and co-Büchi acceptance conditions, where $\alpha\subseteq Q$ is a subset of states.
+We use the terms $\alpha$-states and $\bar{\alpha}$-states to refer to states in $\alpha$ and in $Q\setminus\alpha$, respectively.
+For a run $r$, let ${\it inf}(r)\subseteq Q$ be the set of states that $r$ traverses infinitely often.
+Thus, ${\it inf}(r)=\{q\in Q:q=r_{i}\text{ for infinitely many $i$'s}\}$.
+A run $r$ of a Büchi automaton is accepting iff it visits states in $\alpha$ infinitely often, thus ${\it inf}(r)\cap\alpha\neq\emptyset$.
+Dually, a run $r$ of a co-Büchi automaton is accepting iff it visits states in $\alpha$ only finitely often, thus ${\it inf}(r)\cap\alpha=\emptyset$.
+A run that is not accepting is rejecting.
+Note that as $\mathcal{A}$} is nondeterministic, it may have several runs on a word $w$.
+The word $w$ is accepted by $\mathcal{A}$} if there is an accepting run of $\mathcal{A}$} on $w$.
+The language of $\mathcal{A}$}, denoted $L({\mathcal{A}})$, is the set of words that $\mathcal{A}$} accepts.
+Two automata are equivalent if their languages are equivalent.
 Consider a directed graph $G=\langle V,E\rangle$.
@@
 Theorem 2.
-[Semantic Hierarchy] For $X\in\{B,W\}$, we have that DXW $=$ DBP-NXW $=$ HD-NXW $=$ SD-NXW $<$ NXW.
+[Semantic Hierarchy] For $X\in\{B,W\}$, we have that DXW == DBP-NXW == HD-NXW == SD-NXW << NXW.
 Proof 3.2.
@@
 Moreover, it preserves weakness.
-Thus, DBW$=$SD-NBW and DWW$=$SD-NWW.
-Also, the last inequality follows from the fact DBW$<$NBW and DWW$<$NWW [Lan69].
-Given an NBW ${\mathcal{A}}=\left\langle\Sigma,Q,q_{0},\delta,\alpha\right\rangle$, the DBW generated in [KS15] [^4] The construction in [KS15] assumes automata with transition-based acceptance, and (regardless of this) is slightly different: when $\alpha$ is visited, ${\mathcal{A}}^{\prime}$ continues with a single state from the set of successors.
+Thus, DBW==SD-NBW and DWW==SD-NWW.
+Also, the last inequality follows from the fact DBW<<NBW and DWW<<NWW [Lan69].
+Given an NBW ${\mathcal{A}}=\left\langle\Sigma,Q,q_{0},\delta,\alpha\right\rangle$, the DBW ${\mathcal{A}}^{\prime}=\left\langle\Sigma,Q^{\prime},q^{\prime}_{0},\delta^{\prime},\alpha^{\prime}\right\rangle$ generated in [KS15] [^4] The construction in [KS15] assumes automata with transition-based acceptance, and (regardless of this) is slightly different: when $\alpha$ is visited, $\mathcal{A}$$^{\prime}$ continues with a single state from the set of successors.
 However, the key point remains the same: $\mathcal{A}$} being SD enables $\mathcal{A}$$^{\prime}$ to maintain only subsets of states, rather than Safra trees, which makes determinization much easier.
-is ${\mathcal{A}}^{\prime}=\left\langle\Sigma,Q^{\prime},q^{\prime}_{0},\delta^{\prime},\alpha^{\prime}\right\rangle$, where $Q^{\prime}=2^{Q}$, $q^{\prime}_{0}=\{q_{0}\}$, $\alpha^{\prime}=\{S\in 2^{Q}:S\subseteq\alpha\}$, and the transition function $\delta^{\prime}$ is defined for every subset $S\in 2^{Q}$ and letter $\sigma\in\Sigma$ as follows.
+is such that $Q^{\prime}=2^{Q}$, $q^{\prime}_{0}=\{q_{0}\}$, $\alpha^{\prime}=\{S\in 2^{Q}:S\subseteq\alpha\}$, and the transition function $\delta^{\prime}$ is defined for every subset $S\in 2^{Q}$ and letter $\sigma\in\Sigma$ as follows.
 If $\delta(S,\sigma)\cap\alpha=\emptyset$, then $\delta^{\prime}(S,\sigma)=\delta(S,\sigma)$.
@@
 Consider a SAT instance $\varphi$ over the variable set $X=\{x_{1},\ldots,x_{n}\}$ and with $m\geq 1$ clauses $C=\{c_{1},\ldots,c_{m}\}$.
-For $n\geq 1$, let $[n]=\{1,2,\ldots,n\}$.
+For $n\geq 1$, let $n]=\{1,2,\ldots,n\}$.
 For a variable $x_{k}\in X$, let $C^{0}_{k}\subseteq C$ be the set of clauses in which $x_{k}$ appears negatively, and let $C^{1}_{k}\subseteq C$ be the set of clauses in which $x_{k}$ appears positively.
@@
 For the lower bound, we analyze the reduction in [KM18] and prove that the NCW constructed there is HD.
-The reduction is from the Hamiltonian-cycle problem: given a connected graph $G=\left\langle[n],E\right\rangle$, the reduction outputs an NCW ${\mathcal{A}}_{G}$ over the alphabet $[n]$ that is obtained from $G$ by adding self loops to all vertices, labelling the loop at a vertex $i$ by the letter $i$, and labelling the edges from vertex $i$ to all its neighbours in $G$ by every letter $j\neq i$.
-Then, the co-Büchi condition requires a run to eventually get stuck at a self-loop [^7] The exact reduction is more complicated and involves an additional letter ##\## that forces each deterministic pruning of ${\mathcal{A}}_{G}$ to proceed to the same neighbour of $i$ upon reading a letter $j\neq i$ from the vertex $i$..
+The reduction is from the Hamiltonian-cycle problem: given a connected graph $G=\left\langle[n],E\right\rangle$, the reduction outputs an NCW ${\mathcal{A}}_{G}$ over the alphabet $[n]$ that is obtained from $G$ by adding self loops to all vertices, labelling the loop at a vertex $i$ by the letter $i$, and labelling the edges from vertex $i$ to all its neighbors in $G$ by every letter $j\neq i$.
+Then, the co-Büchi condition requires a run to eventually get stuck at a self-loop [^7] The exact reduction is more complicated and involves an additional letter #\# that forces each deterministic pruning of ${\mathcal{A}}_{G}$ to proceed to the same neighbor of $i$ upon reading a letter $j\neq i$ from the vertex $i$..
 Accordingly, $L({\mathcal{A}}_{G})=[n]^{*}\cdot\bigcup_{i\in[n]}i^{\omega}$.
 It is not hard to see that ${\mathcal{A}}_{G}$ is HD.
-Indeed, an HD strategy can decide to which neighbour of $i$ to proceed with a letter $j\neq i$ by following a cycle $c$ that traverses all the vertices of the graph $G$.
-Since when we read $j\neq i$ at vertex $i$ we move to a neighbour state, then by following the cycle $c$ upon reading $i^{\omega}$, we eventually reach the vertex $i$ and get stuck at the $i$-labeled loop.
+Indeed, an HD strategy can decide to which neighbor of $i$ to proceed with a letter $j\neq i$ by following a cycle $c$ that traverses all the vertices of the graph $G$.
+Since when we read $j\neq i$ at vertex $i$ we move to a neighbor state, then by following the cycle $c$ upon reading $i^{\omega}$, we eventually reach the vertex $i$ and get stuck at the $i$-labeled loop.
 Thus, the Hamiltonian-cycle problem is reduced to DBPness of an HD-NCW, and we are done.
@@
 Thus, there is an SD-NCW that is not almost-DBP, yet all HD-NCWs are almost-DBP.
-[…]
+We start with the positive result about Büchi automata.
+Consider an NBW ${\mathcal{A}}=\left\langle\Sigma,Q,q_{0},\delta,\alpha\right\rangle$.
+We define a simple stochastic Büchi game $\mathcal{G}_{\mathcal{A}}$ as follows.
+[^9] In [CJH03] these games are called simple $1\frac{1}{2}$ -player games with Büchi winning objectives and almost-sure winning criterion.
+The game is played between Random and Eve.
+The set of positions of Random are $Q$, and the set of positions of Eve are $Q\times\Sigma$.
+The game starts from position $q_{0}$.
+A round in the game starts at some position $q\in Q$ and proceeds as follows.
   1. (1)
@@
 Otherwise, $r_{g}$ reaches a non $\alpha$-free ergodic SCC, and with probability $1$, it visits all the states in that SCC.
-Thus, with probability $1$, we have $\inf(r_{g})\cap\alpha\neq 0$, and $r_{g}$ is winning for Eve.
+Thus, with probability 11, we have $\inf(r_{g})\cap\alpha\neq 0$, and $r_{g}$ is winning for Eve.
 Overall, Eve wins ${\mathcal{G}}_{\mathcal{A}}$ with probability $1$ when playing according to $g$.
@@
 Since $f$ is almost-sure winning for Eve, we have that with probability 1, either $r_{f}$ is an accepting run or $\inf(r_{f})\subseteq Q_{rej}$.
-Thus, it is sufficient to prove that if $\inf(r_{f})\subseteq Q_{rej}$, then $w_{f}\in L({\mathcal{A}})$ with probability 00, since otherwise, almost surely $r_{f}$ is accepting, and in particular is correct.
+Thus, it is sufficient to prove that if $\inf(r_{f})\subseteq Q_{rej}$, then $w_{f}\in L({\mathcal{A}})$ with probability 0, since otherwise, almost surely $r_{f}$ is accepting, and in particular is correct.
 By semantic determinism, for all $i\geq 0$, it holds that $w_{f}\in L({\mathcal{A}})$ iff $w_{f}[i+1,\infty]\in L({\mathcal{A}}^{q_{f}^{i}})$.
@@
 Thus, ${\mathcal{A}}_{2}$ is not almost-DBP.
-[…]
+Consider a language $L\subseteq\Sigma^{\omega}$ of infinite words.
+We say that a finite word $x\in\Sigma^{*}$ is a good prefix for $L$ if $x\cdot\Sigma^{\omega}\subseteq L$.
+Then, $L$ is a co-safety language if every word in $L$ has a good prefix [AS87].
+Let $\mbox{\it co-safe}(L)=\{x\cdot w\in\Sigma^{\omega}:\text{$x$ is a good prefix of }L\}$.
+Clearly, $\mbox{\it co-safe}(L)\subseteq L$.
+The other direction is not necessarily true.
+For example, if $L\subseteq\{a,b\}^{\omega}$ is the set of all words with infinitely many $a$’s, then $\mbox{\it co-safe}(L)=\emptyset$.
+In fact, $\mbox{\it co-safe}(L)=L$ iff $L$ is a co-safety language.
+As we show now, when $L$ is NCW-recognizable, we can relate $L$ and $\mbox{\it co-safe}(L)$ as follows.
 Lemma 12.
@@
 Since the later happens with probability 0, we have that $\mathbb{P}(L\setminus\mbox{\it co-safe}(L))=0$.
-[…]
+By Lemma 12, pruning an NCW in a way that would make it recognize $\mbox{\it co-safe}(L({\mathcal{A}}))$ results in a DCW that approximates $\mathcal{A}$}, and thus witnesses that $\mathcal{A}$} is almost-DBP.
+We now show that for HD-NCWs, such a pruning is possible, and conclude that HD-NCWs are almost-DBP.
 Theorem 13.
@@
 Thus, $L({\mathcal{A}}^{\prime})\subseteq\mbox{\it co-safe}(L({\mathcal{A}}))$.
-[…]
+In order to see that $\mathcal{A}$$^{\prime}$ is HD and that $\mbox{\it co-safe}(L({\mathcal{A}}))\subseteq L({\mathcal{A}}^{\prime})$, consider an HD strategy $f$ of $\mathcal{A}$}.
+We claim that $f$ is also an HD strategy for $\mathcal{A}$$^{\prime}$, and for all words $w\in\mbox{\it co-safe}(L({\mathcal{A}}))$, the run $r$ that $f$ generates on $w$ eventually visits only states $q\notin\alpha^{\prime}$.
+Since $\mbox{\it co-safe}(L({\mathcal{A}}))\subseteq L({\mathcal{A}})$, we know that $\inf(r)\cap\alpha=\emptyset$.
+It is left to prove that $r$ visits only finitely many states $q\in Q$ with $L({\mathcal{A}}^{q})\neq\Sigma^{\omega}$.
 Since $w\in\mbox{\it co-safe}(L({\mathcal{A}}))$, it has a good prefix $x\in\Sigma^{*}$.
@@
 Thus, $\inf(r)\cap\alpha^{\prime}=\emptyset$, and we are done.
-[…]
-[…]
+So, $\mathcal{A}$$^{\prime}$ is an HD-NCW with $L({\mathcal{A}}^{\prime})=\mbox{\it co-safe}(L({\mathcal{A}}))$.
+Since $\mbox{\it co-safe}(L({\mathcal{A}}))$ is co-safe, it is DWW-recognizable [Sis94].
+By [BKS17], HD-NCWs whose language is DWW-realizable are DBP.
+Let $\delta^{\prime}$ be the restriction of $\delta$ to a deterministic transition function such that ${\mathcal{D}}^{\prime}=\left\langle\Sigma,Q,q_{0},\delta^{\prime},\alpha^{\prime}\right\rangle$ is a DCW with $L({\mathcal{D}}^{\prime})=L({\mathcal{A}}^{\prime})=\mbox{\it co-safe}(L({\mathcal{A}}))$.
+Consider now the DCW ${\mathcal{D}}=\left\langle\Sigma,Q,q_{0},\delta^{\prime},\alpha\right\rangle$ that is obtained from $\mathcal{D}$$^{\prime}$ by replacing $\alpha^{\prime}$ with $\alpha$.
+It is clear that $\mathcal{D}$} is a pruning of $\mathcal{A}$}.
+Note that, $\alpha\subseteq\alpha^{\prime}$, and hence $\mbox{\it co-safe}(L({\mathcal{A}}))=L({\mathcal{D}}^{\prime})\subseteq L({\mathcal{D}})$.
+That is, $\mathcal{D}$} is a pruning of $\mathcal{A}$} that approximates $L({\mathcal{A}})$ up to a negligible set, and $\mathcal{A}$} is almost-DBP.
 6. Nondeterminism and Reasoning about MDPs
@@
 This process can be modeled by an MDP that is a product of $\mathcal{M}$} with $\mathcal{N}$}, and in each step the player takes an action of $\mathcal{M}$} and chooses a successor state in $\mathcal{N}$} that extends the current run on the new letter drawn by $\mathcal{M}$}.
-Formally, we define the MDP ×${\mathcal{M}}\times{\mathcal{N}}=\left\langle S^{\prime},\left\langle s_{0},\bot\right\rangle,A^{\prime},\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{N}}},\Sigma,\gamma\right\rangle$ that is defined as follows.
-The set of states of ×${\mathcal{M}}\times{\mathcal{N}}$ is $S^{\prime}=(S\times Q)\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}$, for $\bot\notin Q$.
+Formally, we define the MDP ${\mathcal{M}}\times{\mathcal{N}}=\left\langle S^{\prime},\left\langle s_{0},\bot\right\rangle,A^{\prime},\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{N}}},\Sigma,\gamma\right\rangle$ that is defined as follows.
+The set of states of ${\mathcal{M}}\times{\mathcal{N}}$ is $S^{\prime}=(S\times Q)\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}$, for $\bot\notin Q$.
 The set of actions $A^{\prime}=A\left\langle s_{0},\bot\right\rangle\cup\bigcup_{\left\langle s,q\right\rangle\in S\times Q}(A_{\left\langle s,q\right\rangle})$ is defined as follows.
@@
 $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}=\sup_{\mu\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})}\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu).$
-Maximizing the syntactic satisfaction can be expressed as a simple stochastic parity game that is played on top of the MDP ×${\mathcal{M}}\times{\mathcal{N}}$, where the winning objective is inherited from ${\mathcal{N}}$.
+Maximizing the syntactic satisfaction can be expressed as a simple stochastic parity game that is played on top of the MDP ${\mathcal{M}}\times{\mathcal{N}}$, where the winning objective is inherited from $\mathcal{N}$}.
 In  [CH03], the authors prove that simple stochastic parity games admit optimal pure memoryless strategies.
 Thus, the supremum in the definition of $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}$ is attained by a pure memoryless strategy.
-Note that $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}$ is evaluated with respect to the MDP ×${\mathcal{M}}\times{\mathcal{N}}$, and thus by pure memoryless we mean that there is a strategy $\mu^{*}_{\mathcal{N}}:S\times Q\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}\rightarrow A\times Q$ such that $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}(\mu^{*}_{\mathcal{N}})=\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}$.
-Similarly, the semantic satisfaction can be expressed as a simple stochastic parity game that augments ×${\mathcal{M}}\times{\mathcal{N}}$ with a deterministic automaton ${\mathcal{D}}$ for $L({\mathcal{N}})$.
+Note that $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}$ is evaluated with respect to the MDP ${\mathcal{M}}\times{\mathcal{N}}$, and thus by pure memoryless we mean that there is a strategy $\mu^{*}_{\mathcal{N}}:S\times Q\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}\rightarrow A\times Q$ such that $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}(\mu^{*}_{\mathcal{N}})=\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}$.
+Similarly, the semantic satisfaction can be expressed as a simple stochastic parity game that augments ${\mathcal{M}}\times{\mathcal{N}}$ with a deterministic automaton $\mathcal{D}$} for $L({\mathcal{N}})$.
 Thus, the supremum in the definition of $\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}$ is attained by a pure finite-memory strategy.
@@
 When the equality holds for all MDP $\mathcal{M}$}, we say that $\mathcal{N}$} is good-for-MDP (GFM for short).
-Note that when ${\mathcal{N}}$ is GFM, it holds that for all MDP ${\mathcal{M}}=\left\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{{\mathcal{M}}},\Sigma,\tau\right\rangle$, there exists a pure-memroyless strategy for ×${\mathcal{M}}\times{\mathcal{N}}$, $\mu^{*}_{\mathcal{N}}:S\times Q\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}\rightarrow A\times Q$, such that $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}=\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}(\mu^{*}_{\mathcal{N}})$ and $\mu^{*}_{\mathcal{N}}$ generates a rejecting run on words in $L({\mathcal{N}})$ with zero probability.
+Note that when $\mathcal{N}$} is GFM, it holds that for all MDP ${\mathcal{M}}=\left\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{{\mathcal{M}}},\Sigma,\tau\right\rangle$, there exists a pure-memroyless strategy for ${\mathcal{M}}\times{\mathcal{N}}$, $\mu^{*}_{\mathcal{N}}:S\times Q\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}\rightarrow A\times Q$, such that $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}=\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}(\mu^{*}_{\mathcal{N}})$ and $\mu^{*}_{\mathcal{N}}$ generates a rejecting run on words in $L({\mathcal{N}})$ with zero probability.
 Note that $\mu^{*}_{\mathcal{N}}$ can be viewed as a pure finite-memory strategy for $\mathcal{M}$} for which $\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}=\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}(\mu^{*}_{\mathcal{N}})$.
@@
 Thus, with respect to the syntactic hierarchy, GFM-NBWs are between SD and nondeterministic Büchi automata, while GFM-NCWs are incomparable to SD-NCWs.
-Theorem 14.
-Every HD automaton is GFM [KMBK14].
+{thmC}
+[[KMBK14]] Every HD automaton is GFM.
 Proof 6.1.
@@
 In particular $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$ equals the probability that a random word $w$ is in $L({\mathcal{N}})$.
-Note that it must be that $\mu^{*}(\sigma_{0},\bot)=\left\langle\ast,q_{0}\right\rangle$; i.e., there is no choice at the initial state of ×${\mathcal{M}}\times{\mathcal{N}}$.
+Note that it must be that $\mu^{*}(\sigma_{0},\bot)=\left\langle\ast,q_{0}\right\rangle$; i.e., there is no choice at the initial state of ${\mathcal{M}}\times{\mathcal{N}}$.
 Also, as $A_{\sigma}=\left\{\ast\right\}$, for all $\sigma\in\Sigma$, we can think of $\mu^{*}$ as a function $\mu^{*}:\Sigma\times Q_{\mathcal{N}}\rightarrow Q_{\mathcal{N}}$, with $\mu^{*}(\sigma,q)\in\delta_{\mathcal{N}}(q,\sigma)$ for all $\left\langle\sigma,q\right\rangle\in\Sigma\times Q_{\mathcal{N}}$.
 Thus, $\mu^{*}$ induces a pruning ${\mathcal{N}}_{\mu^{*}}$ of $\mathcal{N}$}, and $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{*})$ equals to the probability that a random word is accepted by ${\mathcal{N}}_{\mu^{*}}$.
-Now, since $\mu^{*}$ is an optimal strategy for ×${\mathcal{M}}\times{\mathcal{N}}$ and ${\mathcal{N}}$ is GFM, it follows that $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{*})=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$.
+Now, since $\mu^{*}$ is an optimal strategy for ${\mathcal{M}}\times{\mathcal{N}}$ and $\mathcal{N}$} is GFM, it follows that $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{*})=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$.
 Hence, the probability that a random word $w\in\Sigma^{\omega}$ is accepted by ${\mathcal{N}}_{\mu^{*}}$ equals the probability that $w\in L({\mathcal{N}})$.
@@
 By Theorem 14, we can conclude that the SD-NCW described there, is also not GFM.
-Thus, while HDness implies GFMness for all accaptence conditions [KMBK14], for co-Büchi automata, SDness need not imply GFMness.
+Thus, while HDness implies GFMness for all acceptance conditions [KMBK14], for co-Büchi automata, SDness need not imply GFMness.
 In the rest of this section we prove that for Büchi automata, SDness does imply GFMness, thus all SD-NBWs are GFM.
@@
 Let $r=\left\langle\left\langle s_{0},m_{0}\right\rangle,\left\langle s_{1},m_{1}\right\rangle,\left\langle s_{2},m_{2}\right\rangle,\ldots\right\rangle\in(S\times M)^{\omega}$ be a random process sampled according to $\mathcal{C}$}.
-Let $r_{\mathcal{N}}=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle\in(Q_{\mathcal{N}})^{\omega}$ be a random run of ${\mathcal{N}}$ over the word $\tau(r)$ that is sampled when playing in ${\mathcal{M}}$ with the startegy $\mu$.
+Let $r_{\mathcal{N}}=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle\in(Q_{\mathcal{N}})^{\omega}$ be a random run of $\mathcal{N}$} over the word $\tau(r)$ that is sampled when playing in $\mathcal{M}$} with the strategy $\mu$.
 I.e., given $\left\langle s_{0},m_{0}\right\rangle,\left\langle s_{1},m_{1}\right\rangle,\ldots,\left\langle s_{k},m_{k}\right\rangle$, for $k\geq 1$, and $q_{0},q_{1},\ldots,q_{k-1}$, we set $q_{k}$ to be one of the $\tau(s_{k})$-successors of $q_{k-1}$ at random.
@@
 In particular $\left\langle s_{j},m_{j},q^{\prime}_{j}\right\rangle\in C$, in contradiction to the assumption that $C$ is not $\alpha_{\mathcal{N}}$-accepting.
-Let $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ be the strategy for ×${\mathcal{M}}\times{\mathcal{N}}$ that resolves the actions of ${\mathcal{M}}$ according to $\mu$ and in addition generates a random run of ${\mathcal{N}}$ over $\tau(r)$.
+Let $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ be the strategy for ${\mathcal{M}}\times{\mathcal{N}}$ that resolves the actions of $\mathcal{M}$} according to $\mu$ and in addition generates a random run of $\mathcal{N}$} over $\tau(r)$.
 It is not hard to see that ${\mathcal{C}}_{\mathcal{N}}$ induces the same distribution over infinite paths in $(S\times Q_{\mathcal{N}})^{\omega}$ as the Markov chain ${\mathcal{C}}({\mathcal{M}}\times{\mathcal{N}},\mu^{\prime})$.
@@
 Proof 6.4.
-The co-Büchi case follows from Theorems 11 and 15.
+The co-Büchi case follows from Theorems 11 and 14.
 For Büchi automata, consider an SD-NBW ${\mathcal{N}}=\left\langle\Sigma,Q_{\mathcal{N}},q_{0},\delta_{\mathcal{N}},\alpha_{\mathcal{N}}\right\rangle$, and a $\Sigma$-labeled MDP $\mathcal{M}$}.
@@
 To conclude the GFMness of $\mathcal{N}$}, we show that there is a strategy $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ such that $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})$.
-Recall that a strategy for ×${\mathcal{M}}\times{\mathcal{N}}$ needs to take actions in ${\mathcal{M}}$ and in ${\mathcal{N}}$ simultaneously.
-In the first it needs to choose an action $a\in A_{s}$ where $s\in S$ is the current state in ${\mathcal{M}}$, while in the latter it should choose a $\tau(s)$-succsessor of the current state $q$ in ${\mathcal{N}}$.
-The strategy $\mu^{\prime}$ resolves the first according to to $\mu$, and hence ignores the ${\mathcal{N}}$ component.
+Recall that a strategy for ${\mathcal{M}}\times{\mathcal{N}}$ needs to take actions in $\mathcal{M}$} and in $\mathcal{N}$} simultaneously.
+In the first it needs to choose an action $a\in A_{s}$ where $s\in S$ is the current state in $\mathcal{M}$}, while in the latter it should choose a $\tau(s)$-successor of the current state $q$ in $\mathcal{N}$}.
+The strategy $\mu^{\prime}$ resolves the first according to $\mu$, and hence ignores the $\mathcal{N}$} component.
 Then, for the $\mathcal{N}$} component, $\mu^{\prime}$ generates a run $r_{\mathcal{N}}$ of $\mathcal{N}$} on $w_{\mathcal{C}}$ as follows.
@@
 Hence, with probability 1, we have one of the following outcomes.
-Either the ergodic component that is reached does not contains an $a$-labeled state, or it contains an $a$-labeled state.
+Either the ergodic component that is reached does not contain an $a$-labeled state, or it contains an $a$-labeled state.
 In the former case, the run $r_{\mathcal{N}}$ is accepting.
@@
 Thus, when $T$ runs on the empty tape, it uses at most $n_{0}=s(0)$ cells.
-We use strings of the form #$\#\gamma_{1}\gamma_{2}\ldots(q,\gamma_{i})\ldots\gamma_{n_{0}}$ to encode a configuration of $T$ on a word of length at most $n_{0}$.
-That is, a configuration starts with ##\##, and all its other letters are in $\Gamma$, except for one letter in $Q\times\Gamma$.
+We use strings of the form $\gamma_{1}\gamma_{2}\ldots(q,\gamma_{i})\ldots\gamma_{n_{0}}$ to encode a configuration of $T$ on a word of length at most $n_{0}$.
+That is, a configuration starts with #\#, and all its other letters are in $\Gamma$, except for one letter in $Q\times\Gamma$.
 The meaning of such a configuration is that the $j$’th cell in $T$, for $1\leq j\leq n_{0}$, is labeled $\gamma_{j}$, the reading head points at cell $i$, and $T$ is in state $q$.
-For example, the initial configuration of $T$ is #$\#(q_{0},b)b\ldots b$ (with $n_{0}-1$ occurrences of $b$’s) where $b$ stands for an empty cell.
+For example, the initial configuration of $T$ is $(q_{0},b)b\ldots b$ (with $n_{0}-1$ occurrences of $b$’s) where $b$ stands for an empty cell.
 We can now encode a computation of $T$ by a sequence of configurations.
-Let $\Sigma=\{\#\}\cup\Gamma\cup(Q\times\Gamma)$ and let #$\#\sigma_{1}\ldots\sigma_{n_{0}}\#\sigma^{\prime}_{1}\ldots\sigma^{\prime}_{n_{0}}$ be two successive configurations of $T$.
-We also set $\sigma_{0}$, $\sigma^{\prime}_{0}$, and $\sigma_{n_{0}+1}$ to ##\##.
+Let $\Sigma=\{\#\}\cup\Gamma\cup(Q\times\Gamma)$ and let $\sigma_{1}\ldots\sigma_{n_{0}}\#\sigma^{\prime}_{1}\ldots\sigma^{\prime}_{n_{0}}$ be two successive configurations of $T$.
+We also set $\sigma_{0}$, $\sigma^{\prime}_{0}$, and $\sigma_{n_{0}+1}$ to #\#.
 For each triple $\left\langle\sigma_{i-1},\sigma_{i},\sigma_{i+1}\right\rangle$ with $1\leq i\leq n_{0}$, we know, by the transition relation of $T$, what $\sigma^{\prime}_{i}$ should be.
-In addition, the letter ##\## should repeat exactly every $n_{0}+1$ letters.
+In addition, the letter #\# should repeat exactly every $n_{0}+1$ letters.
 Let $next(\sigma_{i-1},\sigma_{i},\sigma_{i+1})$ denote our expectation for $\sigma^{\prime}_{i}$.
@@
   • $next(\sigma_{i-1},\sigma_{i},\sigma_{i+1})=\sigma_{i}$ if $\sigma_{i}=\#$, or if non of $\sigma_{i-1},\sigma_{i}$ and $\sigma_{i+1}$ are in $Q\times\Gamma$.
-  • $next(\sigma_{i-1},\sigma_{i},\sigma_{i+1})=\#$ if |$|\left\{\sigma_{i-1},\sigma_{i},\sigma_{i+1}\right\}\cap Q\times\Gamma|\geq 2$, or if |$|\left\{\sigma_{i-1},\sigma_{i},\sigma_{i+1}\right\}\cap\left\{\#\right\}|\geq 2$  [^10] These case describes an illegal encoding of a configuration, and there is no correct expectation of the letter that comes in the same position in the next configuration.
+  • $next(\sigma_{i-1},\sigma_{i},\sigma_{i+1})=\#$ if $|\left\{\sigma_{i-1},\sigma_{i},\sigma_{i+1}\right\}\cap Q\times\Gamma|\geq 2$, or if $|\left\{\sigma_{i-1},\sigma_{i},\sigma_{i+1}\right\}\cap\left\{\#\right\}|\geq 2$  [^10] These case describes an illegal encoding of a configuration, and there is no correct expectation of the letter that comes in the same position in the next configuration.
 Thus, the choice $next(\sigma_{i-1},\sigma_{i},\sigma_{i+1})=\#$ is arbitrary..
@@
   • $next(\gamma_{i-1},(q,\gamma_{i}),\gamma_{i+1})=next(\gamma_{i-1},(q,\gamma_{i}),\#)=\gamma^{\prime}_{i}$ where $(q,\gamma_{i})\rightarrow(q^{\prime},\gamma^{\prime}_{i},\Delta)$.
-Consistency with $next$ now gives us a necessary condition for a word to encode a legal computation that uses $n_{0}$ tape cells.
-In order to accept words that satisfy C1, namely detect a violation of the encoding, we distinguish between two types of violations: a violation of a single encoded configuration, and a violation of $next$.
-In order to detect a violation of $next$, the NWW ${\mathcal{A}}$ use its nondeterminism and guesses a triple $\left\langle\sigma_{i-1},\sigma_{i},\sigma_{i+1}\right\rangle\in\Sigma^{3}$ and guesses a position in the word, where it checks whether the three letters to be read starting this position are $\sigma_{i-1},\sigma_{i}$, and $\sigma_{i+1}$, and checks whether $next(\sigma_{i-1},\sigma_{i},\sigma_{i+1})$ is not the letter to come $n_{0}+1$ letters later.
+Consistency with nextnext now gives us a necessary condition for a word to encode a legal computation that uses $n_{0}$ tape cells.
+In order to accept words that satisfy C1, namely detect a violation of the encoding, we distinguish between two types of violations: a violation of a single encoded configuration, and a violation of nextnext.
+In order to detect a violation of nextnext, the NWW $\mathcal{A}$} use its nondeterminism and guesses a triple $\left\langle\sigma_{i-1},\sigma_{i},\sigma_{i+1}\right\rangle\in\Sigma^{3}$ and guesses a position in the word, where it checks whether the three letters to be read starting this position are $\sigma_{i-1},\sigma_{i}$, and $\sigma_{i+1}$, and checks whether $next(\sigma_{i-1},\sigma_{i},\sigma_{i+1})$ is not the letter to come $n_{0}+1$ letters later.
 Once $\mathcal{A}$} sees such a violation, it goes to an accepting sink.
-If $next$ is respected, or if the guessed triple and position is not successful, then ${\mathcal{A}}$ returns to its initial state.
+If nextnext is respected, or if the guessed triple and position is not successful, then $\mathcal{A}$} returns to its initial state.
 Also, at any point that $\mathcal{A}$} still waits to guess a position of a triple, it can guess to return back to the initial state.
-In order to detect a violation of a single encoded configuration, the NWW ${\mathcal{A}}$ accepts words that include in them a subword $x$ of length $n_{0}+1$, which is either ##\##-free, or is of the form #$\#\cdot y$, for a finite word $y$ that does not encode a legal configuration that uses $n_{0}$ cells.
-Specifically, $y$ is not of the form $\Gamma^{i}\cdot(Q\times\Gamma)\cdot\Gamma^{n_{0}-(i+1)}$, for some $10\leq i\leq n_{0}-$ 1.
-Indeed, such words either include a long subword with no ##\##, or have some block of size $n_{0}$ that follows a ##\##, that does not encode a legal configuration of $T$ when it uses $n_{0}$ cells.
+In order to detect a violation of a single encoded configuration, the NWW $\mathcal{A}$} accepts words that include in them a subword $x$ of length $n_{0}+1$, which is either #\#-free, or is of the form $\cdot y$, for a finite word $y$ that does not encode a legal configuration that uses $n_{0}$ cells.
+Specifically, $y$ is not of the form $\Gamma^{i}\cdot(Q\times\Gamma)\cdot\Gamma^{n_{0}-(i+1)}$, for some $0\leq i\leq n_{0}-1$.
+Indeed, such words either include a long subword with no #\#, or have some block of size $n_{0}$ that follows a #\#, that does not encode a legal configuration of $T$ when it uses $n_{0}$ cells.
 If $\mathcal{A}$} succeeds to detect such a violation, it goes to an accepting sink, otherwise, $\mathcal{A}$} returns to its initial state.
-In order to accept words that satisfy C2, namely detect an encoding of an accepting configuration of $T$, the NWW ${\mathcal{A}}$ guesses a position where it checks if the next $n_{0}+1$ letters are of the form #$\#\cdot x\cdot(q_{acc},\gamma)\cdot y$, for $x,y\in\Gamma^{*}$ and $\gamma\in\Gamma$.
+In order to accept words that satisfy C2, namely detect an encoding of an accepting configuration of $T$, the NWW $\mathcal{A}$} guesses a position where it checks if the next $n_{0}+1$ letters are of the form $\cdot x\cdot(q_{acc},\gamma)\cdot y$, for $x,y\in\Gamma^{*}$ and $\gamma\in\Gamma$.
 If the guess succeeded, then $\mathcal{A}$} goes to the accepting sink, and otherwise it goes back to the initial state.
@@
 That is, for all $w\in(X\cdot C)^{*}$, $x_{k}\in X$, and $c_{j}\in C$, it holds that $(g(w),x_{k},g(w\cdot x_{k}))$ and $(g(w\cdot x_{k}),c_{j},g(x\cdot x_{k}\cdot c_{j}))$ are in $\Delta_{\varphi}$.
-In addition, if $g$ satisfies the conditions in Lemmas 23 and 24, we say that $g$ supports a $(p,q_{0})$-circle.
+In addition, if $g$ satisfies the conditions in Lemmas 22 and 23, we say that $g$ supports a $(p,q_{0})$-circle.
 Lemma 24.
@@
 We conclude that $q_{0}\in\inf(g(w))$, and hence $g(w)$ is accepting.
-It is easy to see that there is a correspondence between assignments to the variables in $X$ and deterministic prunnings of ${\mathcal{A}}_{\varphi}$.
+It is easy to see that there is a correspondence between assignments to the variables in $X$ and deterministic prunings of ${\mathcal{A}}_{\varphi}$.
 Indeed, a pruning of $p$ amounts to choosing, for each $k\in[n]$, a value $i_{k}\in{\{0,1\}}$: the assignment $x_{k}=i_{k}$ corresponds to keeping the transition $\left\langle p,x_{k},q^{i_{k}}_{k}\right\rangle$ and removing the transition $\left\langle p,x_{k},q^{\lnot i_{k}}_{k}\right\rangle$.
@@
 We prove that $a$ satisfies $\varphi$ iff ${\mathcal{A}}_{\varphi}^{a}$ is equivalent to ${\mathcal{A}}_{\varphi}$.
-Thus, the number of deterministic prunnings of ${\mathcal{A}}_{\varphi}$ that result in a DBW equivalent to ${\mathcal{A}}_{\varphi}$, equals to the number of assignments that satisfy $\varphi$.
+Thus, the number of deterministic prunings of ${\mathcal{A}}_{\varphi}$ that result in a DBW equivalent to ${\mathcal{A}}_{\varphi}$, equals to the number of assignments that satisfy $\varphi$.
 In particular, $\varphi$ is satisfiable iff ${\mathcal{A}}_{\varphi}$ is DBP.

### v6

@@
 Let $A=\bigcup_{s\in S}A_{s}$ be the set of all actions.
-Then, $\mathrm{Pr}_{\mathcal{M}}:S\times A\times S\nrightarrow[0,1]$ is a (partial) stochastic transition function: for every two states $s,s^{\prime}\in S$ and action $a\in A_{s}$, we have that $\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})$ is the probability of moving from $s$ to $s^{\prime}$ when action $a$ is taken.
-Accordingly, for every $s\in S$ and $a\in A_{s}$, we have $\Sigma_{s^{\prime}\in S}\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})=1$.
+Then, $\mathrm{Pr}_{\mathcal{M}}:S\times A\times S\nrightarrow[0,1]$ is a (partial) stochastic transition function: for every two states $s,s^{\prime}\in S$ and action $c\in A_{s}$, we have that $\mathrm{Pr}_{\mathcal{M}}(s,c,s^{\prime})$ is the probability of moving from $s$ to $s^{\prime}$ when action $c$ is taken.
+Accordingly, for every $s\in S$ and $c\in A_{s}$, we have $\Sigma_{s^{\prime}\in S}\mathrm{Pr}_{\mathcal{M}}(s,c,s^{\prime})=1$.
 Finally, $\Sigma$ is a finite alphabet, and $\tau:S\rightarrow\Sigma$ is a labeling function on the states.
@@
 The positions of the game are pairs $\left\langle s,q\right\rangle$, where $s\in S$ is a state of $\mathcal{M}$}, and $q\in Q$ is a state of $\mathcal{N}$}.
-The game starts by the player picking an action $a\in A_{s_{0}}$ available from the initial state of $\mathcal{M}$}, and the game moves to $\left\langle s,q_{0}\right\rangle$ with probability $\mathrm{Pr}_{\mathcal{M}}(s_{0},a,s)$.
-Then, at each round from position $\left\langle s,q\right\rangle$, the player chooses an action $a\in A_{s}$ and a state $q^{\prime}\in\delta(q,\tau(s))$, and then nature chooses a state $s^{\prime}\in S$ with probability $\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})$, and the game moves to $\left\langle s^{\prime},q^{\prime}\right\rangle$.
+The game starts by the player picking an action $c\in A_{s_{0}}$ available from the initial state of $\mathcal{M}$}, and the game moves to $\left\langle s,q_{0}\right\rangle$ with probability $\mathrm{Pr}_{\mathcal{M}}(s_{0},c,s)$.
+Then, at each round from position $\left\langle s,q\right\rangle$, the player chooses an action $c\in A_{s}$ and a state $q^{\prime}\in\delta(q,\tau(s))$, and then nature chooses a state $s^{\prime}\in S$ with probability $\mathrm{Pr}_{\mathcal{M}}(s,c,s^{\prime})$, and the game moves to $\left\langle s^{\prime},q^{\prime}\right\rangle$.
 The objective of the player is to on the one hand maximize the probability of generating a word in $L({\mathcal{N}})$ and on the other hand, to generate a rejecting run on words in $L({\mathcal{N}})$ with zero probability.
@@
 We now formalize this intuition.
-A run of $\mathcal{M}$} is an infinite sequence of states $r=\left\langle s_{0},s_{1},s_{2},\ldots\right\rangle\in S^{\omega}$, such that for all $i\geq 0$ there exists $a\in A_{s_{i}}$ with $\mathrm{Pr}_{\mathcal{M}}(s_{i},a,s_{i+1})>0$.
+A run of $\mathcal{M}$} is an infinite sequence of states $r=\left\langle s_{0},s_{1},s_{2},\ldots\right\rangle\in S^{\omega}$, such that for all $i\geq 0$ there exists $c\in A_{s_{i}}$ with $\mathrm{Pr}_{\mathcal{M}}(s_{i},c,s_{i+1})>0$.
 A finite run is a prefix of such a sequence.
@@
 Let $\mathrm{dist}(A)$ be the set of distributions over $A$.
-A distribution $d\in\mathrm{dist}(A)$ is a point distribution if there exists $a\in A$ such that $d(a)=1$.
+A distribution $d\in\mathrm{dist}(A)$ is a point distribution if there exists $c\in A$ such that $d(c)=1$.
 A strategy for an MDP $\mathcal{M}$} is a function $\mu:\Pi({\mathcal{M}})\rightarrow\mathrm{dist}(A)$, which suggests to the player a distribution on the available actions given the history of the game so far.
-The strategy should suggest an available action, thus $\mu(s_{0},s_{1},s_{2},\ldots,s_{k})(a)>0$ implies that $a\in A_{s_{k}}$.
+The strategy should suggest an available action, thus $\mu(s_{0},s_{1},s_{2},\ldots,s_{k})(c)>0$ implies that $c\in A_{s_{k}}$.
 Let $\Pi({\mathcal{M}},\mu)$ denote the subset of $\Pi({\mathcal{M}})$ that includes exactly all finite runs that agree with the strategy $\mu$.
-Formally, we have that $\left\langle s_{0}\right\rangle\in\Pi({\mathcal{M}},\mu)$, and for all $p^{\prime}=\left\langle s_{0},s_{1},\ldots,s_{k}\right\rangle\in\Pi({\mathcal{M}},\mu)$, and $s_{k+1}\in S$, we have that $p=p^{\prime}\cdot s_{k+1}\in\Pi({\mathcal{M}},\mu)$ iff there exists $a_{k+1}\in A_{s_{k}}$ such that $\mu(p^{\prime})(a_{k+1})>0$ and $\mathrm{Pr}_{{\mathcal{M}}}(s_{k},a_{k+1},s_{k+1})>0$.
+Formally, we have that $\left\langle s_{0}\right\rangle\in\Pi({\mathcal{M}},\mu)$, and for all $p^{\prime}=\left\langle s_{0},s_{1},\ldots,s_{k}\right\rangle\in\Pi({\mathcal{M}},\mu)$, and $s_{k+1}\in S$, we have that $p=p^{\prime}\cdot s_{k+1}\in\Pi({\mathcal{M}},\mu)$ iff there exists $c_{k+1}\in A_{s_{k}}$ such that $\mu(p^{\prime})(c_{k+1})>0$ and $\mathrm{Pr}_{{\mathcal{M}}}(s_{k},c_{k+1},s_{k+1})>0$.
 Thus, a run in $\Pi({\mathcal{M}},\mu)$ can only be extended using actions that have a positive probability according to $\mu$.
@@
 Given a strategy $\mu$, we can obtain from $\mathcal{M}$} and $\mu$ the MC ${\mathcal{C}}({\mathcal{M}},\mu)=\left\langle\Pi({\mathcal{M}},\mu),s_{0},\mathrm{Pr}_{\mathcal{M}}^{\mu}\right\rangle$ in which the choice of actions is resolved according to $\mu$.
-Formally, for all $p\in\Pi({\mathcal{M}},\mu)$ and $s_{1},s_{2}\in S$ such that both $p_{1}=p\cdot s_{1}$ and $p_{2}=p_{1}\cdot s_{2}$ belong to $\Pi({\mathcal{M}},\mu)$, then $\mathrm{Pr}_{\mathcal{M}}^{\mu}(p_{1},p_{2})=\sum_{a\in A_{s_{1}}}\mu(p_{1})(a)\cdot\mathrm{Pr}_{\mathcal{M}}(s_{1},a,s_{2})$.
+Formally, for all $p\in\Pi({\mathcal{M}},\mu)$ and $s_{1},s_{2}\in S$ such that both $p_{1}=p\cdot s_{1}$ and $p_{2}=p_{1}\cdot s_{2}$ belong to $\Pi({\mathcal{M}},\mu)$, then $\mathrm{Pr}_{\mathcal{M}}^{\mu}(p_{1},p_{2})=\sum_{c\in A_{s_{1}}}\mu(p_{1})(c)\cdot\mathrm{Pr}_{\mathcal{M}}(s_{1},c,s_{2})$.
 It is clear that when $\mu$ is memoryless, that is, $\mu$ only depends on the last position in $p_{1}$, then ${\mathrm{Pr}}_{\mathcal{M}}^{\mu}$ only depends on $s_{1}$ and $s_{2}$, and hence the projection of the chain ${\mathcal{C}}({\mathcal{M}},\mu)$ onto the last position in a path is a finite MC.
@@
 Note that if no such $p\in\Pi({\mathcal{M}},\mu)$ with $\iota(p)=s$ exists, then we arbitrarily take $d_{s}$ to be some distribution in $A_{s}$.
-Let ${\mathrm{Pr}^{\prime}}_{{\mathcal{M}}}^{\mu}:S\times S\rightarrow[0,1]$ be the stochastic transition function that is defined by ${\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}(s,s^{\prime})=\sum_{a\in A_{s}}d_{s}(a)\cdot\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})$.
+Let ${\mathrm{Pr}^{\prime}}_{{\mathcal{M}}}^{\mu}:S\times S\rightarrow[0,1]$ be the stochastic transition function that is defined by ${\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}(s,s^{\prime})=\sum_{c\in A_{s}}d_{s}(c)\cdot\mathrm{Pr}_{\mathcal{M}}(s,c,s^{\prime})$.
 Then, if $p_{0},p_{1},p_{2},\ldots\in(\Pi({\mathcal{M}},\mu))^{\omega}$ is an infinite sequence of increasing paths that are sampled according to ${\mathcal{C}}({\mathcal{M}},\mu)$, then it is not hard to see that the sequence $\iota(p_{0}),\iota(p_{1}),\iota(p_{2}),\ldots\in S^{\omega}$ satisfies $\mathbb{P}(\iota(p_{i+1})=s^{\prime}|\iota(p_{i})=s)={\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}(s,s^{\prime})$ for all $s^{\prime},s\in S$.
@@
 The set of actions $A^{\prime}=A\left\langle s_{0},\bot\right\rangle\cup\bigcup_{\left\langle s,q\right\rangle\in S\times Q}(A_{\left\langle s,q\right\rangle})$ is defined as follows.
-For all $\left\langle s,q\right\rangle\in S\times Q$, the set of actions $A_{\left\langle s,q\right\rangle}$ available from $\left\langle s,q\right\rangle$ consists of all pairs $\left\langle a,q^{\prime}\right\rangle\in A_{s}\times\delta(q,\tau(s))$.
+For all $\left\langle s,q\right\rangle\in S\times Q$, the set of actions $A_{\left\langle s,q\right\rangle}$ available from $\left\langle s,q\right\rangle$ consists of all pairs $\left\langle c,q^{\prime}\right\rangle\in A_{s}\times\delta(q,\tau(s))$.
 In addition, the actions available from $\left\langle s_{0},\bot\right\rangle$ are $A_{\left\langle s_{0},\bot\right\rangle}=A_{s_{0}}\times\left\{q_{0}\right\}$.
-Then, for all $\left\langle s,\ast\right\rangle\in S^{\prime}$ and $\left\langle a,q^{\prime}\right\rangle\in A_{\left\langle s,\ast\right\rangle}$, we have $\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{N}}}(\left\langle s,\ast\right\rangle,\left\langle a,q^{\prime}\right\rangle,\left\langle s^{\prime},q^{\prime}\right\rangle)=\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$ and $\gamma(s,\ast)=\tau(s)$.
+Then, for all $\left\langle s,\ast\right\rangle\in S^{\prime}$ and $\left\langle c,q^{\prime}\right\rangle\in A_{\left\langle s,\ast\right\rangle}$, we have $\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{N}}}(\left\langle s,\ast\right\rangle,\left\langle c,q^{\prime}\right\rangle,\left\langle s^{\prime},q^{\prime}\right\rangle)=\mathrm{Pr}_{{\mathcal{M}}}(s,c,s^{\prime})$ and $\gamma(s,\ast)=\tau(s)$.
 Given a finite or infinite run $r\in\Pi({\mathcal{M}}\times{\mathcal{N}})\cup\Omega({\mathcal{M}}\times{\mathcal{N}})$, $r=\left\langle\left\langle s_{0},\bot\right\rangle,\left\langle s_{1},q_{0}\right\rangle,\left\langle s_{2},q_{1}\right\rangle,\ldots\right\rangle$, we define $\pi_{\mathcal{N}}(r)=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle$, and $\pi_{\mathcal{M}}(r)=\left\langle s_{0},s_{1},s_{2},\ldots\right\rangle$ to be the $\mathcal{N}$}-run and $\mathcal{M}$}-run that correspond to $r$ respectively.
@@
 Let $\mu:S\times M\rightarrow\mathrm{dist}(A)\times M$, be a finite memory strategy in ${\mathcal{F}}({\mathcal{M}})$, and let ${\mathcal{C}}={\mathcal{C}}({\mathcal{M}},\mu)$ be the Markov chain induced when taking actions in $\mathcal{M}$} according to $\mu$.
-Note that since $\mu$ uses $M$ as a memory structure, we may think of $\mathcal{C}$} as a Markov chain with positions $S\times M$, and a stochastic transition function $\mathrm{Pr}_{\mathcal{C}}:(S\times M)^{2}\rightarrow[0,1]$ that is defined by $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=\sum_{a\in A_{s}}d(a)\cdot\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$, when $\mu(s,m)=\left\langle d,m^{\prime}\right\rangle$, and $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=0$, otherwise.
+Note that since $\mu$ uses $M$ as a memory structure, we may think of $\mathcal{C}$} as a Markov chain with positions $S\times M$, and a stochastic transition function $\mathrm{Pr}_{\mathcal{C}}:(S\times M)^{2}\rightarrow[0,1]$ that is defined by $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=\sum_{c\in A_{s}}d(c)\cdot\mathrm{Pr}_{{\mathcal{M}}}(s,c,s^{\prime})$, when $\mu(s,m)=\left\langle d,m^{\prime}\right\rangle$, and $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=0$, otherwise.
 Let $r=\left\langle\left\langle s_{0},m_{0}\right\rangle,\left\langle s_{1},m_{1}\right\rangle,\left\langle s_{2},m_{2}\right\rangle,\ldots\right\rangle\in(S\times M)^{\omega}$ be a random process sampled according to $\mathcal{C}$}.
@@
 Consider an MDP $\mathcal{M}$}, and let $\mu:S\times M\to dist(A)\times M$ be a finite memory strategy such that $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)=\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}$.
-Consider the MC ${\mathcal{C}}={\mathcal{C}}({\mathcal{M}},\mu)$, and note that since $\mu$ uses $M$ as a memory structure, we may think of $\mathcal{C}$} as a finite MC with positions $S\times M$, and a stochastic transition function $\mathrm{Pr}_{\mathcal{C}}:(S\times M)^{2}\rightarrow[0,1]$ that is defined $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=\sum_{a\in A_{s}}d(a)\cdot\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$ when $\mu(s,m)=\left\langle d,m^{\prime}\right\rangle$, and otherwise $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=0$.
+Consider the MC ${\mathcal{C}}={\mathcal{C}}({\mathcal{M}},\mu)$, and note that since $\mu$ uses $M$ as a memory structure, we may think of $\mathcal{C}$} as a finite MC with positions $S\times M$, and a stochastic transition function $\mathrm{Pr}_{\mathcal{C}}:(S\times M)^{2}\rightarrow[0,1]$ that is defined $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=\sum_{c\in A_{s}}d(c)\cdot\mathrm{Pr}_{{\mathcal{M}}}(s,c,s^{\prime})$ when $\mu(s,m)=\left\langle d,m^{\prime}\right\rangle$, and otherwise $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=0$.
 To conclude the GFMness of $\mathcal{N}$}, we show that there is a strategy $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ such that $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})$.
 Recall that a strategy for ${\mathcal{M}}\times{\mathcal{N}}$ needs to take actions in $\mathcal{M}$} and in $\mathcal{N}$} simultaneously.
-In the first it needs to choose an action $a\in A_{s}$ where $s\in S$ is the current state in $\mathcal{M}$}, while in the latter it should choose a $\tau(s)$-successor of the current state $q$ in $\mathcal{N}$}.
+In the first it needs to choose an action $c\in A_{s}$ where $s\in S$ is the current state in $\mathcal{M}$}, while in the latter it should choose a $\tau(s)$-successor of the current state $q$ in $\mathcal{N}$}.
 The strategy $\mu^{\prime}$ resolves the first according to $\mu$, and hence ignores the $\mathcal{N}$} component.
@@
     $\left\{\begin{array}[]{ll}\gamma_{i}&\mbox{If $(q,\gamma_{i+1})\rightarrow(q^{\prime},\gamma^{\prime}_{i+1},R)$}\\ (q^{\prime},\gamma_{i})&\mbox{If $(q,\gamma_{i+1})\rightarrow(q^{\prime},\gamma^{\prime}_{i},L)$}\end{array}\right$.
-  • $next(\#,(q,\gamma_{i}),\gamma_{i+1})=(q^{\prime},\gamma^{\prime})$ where $(q,\gamma_{i})\rightarrow(q^{\prime},\gamma^{\prime}_{i},L)$ 111111 We assume that the reading head of $T$ stays in place when it is above the leftmost cell and the transition function directs it to move left..
+  • $next(\#,(q,\gamma_{i}),\gamma_{i+1})=(q^{\prime},\gamma^{\prime})$ where $(q,\gamma_{i})\rightarrow(q^{\prime},\gamma^{\prime}_{i},L)$  [^11] We assume that the reading head of $T$ stays in place when it is above the leftmost cell and the transition function directs it to move left..
   • $next(\gamma_{i-1},(q,\gamma_{i}),\gamma_{i+1})=next(\gamma_{i-1},(q,\gamma_{i}),\#)=\gamma^{\prime}_{i}$ where $(q,\gamma_{i})\rightarrow(q^{\prime},\gamma^{\prime}_{i},\Delta)$.